### v1

# Matrix Rosenthal and Concentration Inequalities for Markov Chains with Applications in Statistical Learning

Abstract
In this paper, we study moment and concentration inequalities of the spectral norm for sums of dependent random matrices. We establish novel Rosenthal-Burkholder inequalities for matrix martingales, as well as matrix Rosenthal, Hoeffding, and Bernstein inequalities for ergodic Markov chains. Compared with previous work on matrix concentration inequalities for Markov chains, our results do not require the assumptions of a non-zero absolute spectral gap and bounded matrix functions. Furthermore, our results have leading terms that match the Markov chain central limit theorem, rather than relying on variance proxies. We also give dimension-free versions of the inequalities, which are independent of the ambient dimension $d$ and relies on the effective rank of some matrix instead. This enables the generalization of our results to linear operators in infinite-dimensional Hilbert spaces. Our results have extensive applications in statistics and machine learning; in particular, we obtain improved bounds in covariance estimation and principal component analysis on Markovian data.
1 Introduction
Concentration inequalities for sums of random matrices (Wigderson and Xiao, 2005; Tropp, 2015) are of great significance for statistics, machine learning, theoretical computer science, signal processing, etc. Common examples include covariance matrix estimation (Chen et al., 2012), principal component analysis (PCA) (Kumar and Sarkar, 2023), and reinforcement learning (Duan and Wainwright, 2022; Wu et al., 2025; Mou, 2025). Although matrix concentration inequalities have been thoroughly studied in the literature (Tropp, 2011, 2012; Mackey et al., 2014; Tropp, 2015; Minsker, 2017; Garg et al., 2018; Klochkov and Zhivotovskiy, 2020; Qiu et al., 2020; Huang and Tropp, 2021; Van Werde and Sanders, 2023; Bandeira et al., 2023; Neeman et al., 2024; Kroshnin and Suvorikova, 2024; Jirak et al., 2024; Wang and Ramdas, 2024; Wu et al., 2025), to the best of our knowledge, the Rosenthal-Burkholder inequality for matrix martingales and matrix Rosenthal inequalities for Markov chains remain absent.
Furthermore, while some existing work (Garg et al., 2018; Qiu et al., 2020; Neeman et al., 2024; Wu et al., 2025) has derived matrix Hoeffding and Bernstein inequalities for Markov chains with a non-zero absolute spectral gap via the multi-matrix Golden-Thompson inequality (Sutter et al., 2017; Garg et al., 2018) and the León-Perron argument (León and Perron, 2004), these studies suffer from the following limitations. First, for non-reversible Markov chains, geometric ergodicity does not imply a non-zero absolute spectral gap (Kontoyiannis and Meyn, 2012, Theorem 1.4), making these results inapplicable to general non-reversible Markov chains, which occur in reinforcement learning (Duan and Wainwright, 2022) and PCA (Kumar and Sarkar, 2023) problems with Markovian data. Second, these works are confined to uniformly geometrically ergodic Markov chains and bounded matrix functions, thus cannot be extended to unbounded matrix functions. Third, the leading terms of these results do not match the Markov chain central limit theorem (CLT) (Jones, 2004, Corollary 5), relying instead on variance proxies. Finally, due to the use of the multi-matrix Golden-Thompson inequality (Garg et al., 2018, Theorem 1.5), this incurs an additional dependence on the dimension $d$, making it impossible to generalize these results to a dimension-free version. Accordingly, these results cannot be applied to linear operators in infinite-dimensional Hilbert spaces.
The goal of this paper is to bridge the gaps mentioned earlier. Specifically, we first give a Rosenthal-Burkholder inequality (Rosenthal, 1970; Burkholder, 1973; Hitczenko, 1990) for matrix martingales in Theorem 2.1. That is, for an $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$-adapted $d$-dimensional self-adjoint matrix martingale $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$, letting martingale differences ${\bm{X}}_{k}={\bm{M}}_{k}-{\bm{M}}_{k-1}$ and quadratic variation $\left\langle{{\bm{M}}}\right\rangle_{\infty}=\sum_{k=0}^{\infty}{\mathbb{E}}[{\bm{X}}_{k}^{2}|{\mathscr{F}}_{k-1}]$, one has that
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq C_{R,1}\ \sqrt{p\vee\log d}\Big{\|}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\Big{\|}_{p}+C_{R,2}\ (p\vee\log d)\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|}\right\|_{p}.$ (1)
Here, $\|\cdot\|$ is the spectral norm of the matrix, $\|\cdot\|_{p}$ is the $L^{p}$ norm of random variables for any $p\geq 2$, and $C_{R,1},C_{R,2}>0$ are universal constants. This result matches the classic Hilbert space-valued Rosenthal-Burkholder inequality (Pinelis, 1994, Theorem 4.1), with the only difference that an additional $\log d$ term appears. This is an unavoidable price to pay when extending Hilbert-space-valued probabilistic inequalities to matrix versions.
Based on the Rosenthal-Burkholder inequality in Eqn. (1), we further derive matrix Rosenthal inequalities for Markov chains by leveraging the approach of Poisson equations (see Eqn. (12)) (Durmus et al., 2023), which allows us to overcome the limitations mentioned above. In particular, we consider the stationary Markov chain $(Z_{k})_{k\in{\mathbb{N}}}$ in a measurable space $\mathcal{Z}$} with Markov kernel $\mathrm{Q}$} and stationary distribution $\pi$, and a self-adjoint matrix function ${\bm{F}}\colon{\mathcal{Z}}\to{\mathbb{H}}^{d}$ satisfying $\pi({\bm{F}}):=\int_{{\mathcal{Z}}}{\bm{F}}(z)\pi(dz)=\bm{0}_{d\times d}$. Our goal is to give an upper bound for the spectral norm of the summation ${\bm{S}}_{n}:=\sum_{k=0}^{n-1}{{\bm{F}}}(Z_{k})$. By the Markov chain CLT (Jones, 2004, Corollary 5), under some regularity conditions, we have $n^{-1/2}{\bm{S}}_{n}$ converges weakly to a Gaussian distribution and satisfies $\lim_{n\to\infty}n^{-1}{\mathbb{E}}[{\bm{S}}_{n}^{2}]={\bm{\Sigma}}_{\pi}({\bm{F}})$, where ${\bm{\Sigma}}_{\pi}({\bm{F}})$ is the long-run variance defined as
${\bm{\Sigma}}_{\pi}({\bm{F}}):=\pi\left({{{\bm{F}}}^{2}}\right)+\sum_{k=1}^{\infty}\pi\left({{{\bm{F}}}{\mathrm{Q}}^{k}{{\bm{F}}}+\left({{\mathrm{Q}}^{k}{{\bm{F}}}}\right){{\bm{F}}}}\right),$ (2)
with $({\mathrm{Q}}{\bm{F}})(\cdot)=\int_{{\mathcal{Z}}}{\bm{F}}(z){\mathrm{Q}}(\cdot,dz)$. Under the uniform geometric ergodicity assumption (Assumption 2.1) and the geometric $V$-ergodicity assumption with Lyapunov function $V$ (Assumption 2.2), in Theorem 2.2 and Theorem 2.4 we show that for any $p\geq 2$, the following Rosenthal-type inequality holds:
$\frac{1}{n}\left\|{\sup_{k\leq n}\|{\bm{S}}_{k}\|}\right\|_{p}\leq C_{R,1}\sqrt{\frac{(p\vee\log d)\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|}{n}}+O\left({\frac{1}{n^{3/4}}}\right).$ (3)
The first term of the r.h.s. of Eqn.(3) is the leading term, which matches the Markov chain CLT.
By Markov inequality, in Corollary 2.1 and Corollary 2.2 we also translate the Rosenthal inequality (Eqn. (3)) into a Hoeffding and Bernstein high-probability upper-bounds: for any $\delta\in(0,1)$, with probability at least $1-\delta$, it holds that
$\frac{1}{n}\sup_{k\leq n}\|{\bm{S}}_{k}\|\leq\frac{16}{3}C_{R,1}\sqrt{\frac{e^{2}\log(\delta^{-1}\vee d)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}^{2}}{n}}+O\left({\frac{1}{n}}\right),$ (4)
$\frac{1}{n}\sup_{k\leq n}\|{\bm{S}}_{k}\|\leq C_{R,1}\sqrt{\frac{e^{2}\log(\delta^{-1}\vee d)\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|}{n}}+O\left({\frac{1}{n^{3/4}}}\right),$ (5)
where $t_{\operatorname{mix}}$ is the mixing time and $t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}^{2}$ is a variance proxy, which is an upper bound of the long-run variance $\|{\bm{\Sigma}}_{\pi}({\bm{F}})\|$. Neeman et al. (2024, Theorem 2.6) gave a Bernstein inequality with a leading term relying on the variance proxy $\gamma^{-1}\|\pi({\bm{F}}^{2})\|$ which is also an upper bound of $\|{\bm{\Sigma}}_{\pi}({\bm{F}})\|$, where $\gamma$ is the absolute spectral gap. Hence our results are sharper in the leading term than the previous work on the matrix concentration inequalities for Markov chains (Garg et al., 2018; Qiu et al., 2020; Neeman et al., 2024; Wu et al., 2025) if we consider time-homogeneous matrix functions.
All of the above inequalities (1), (3), (4), (5) can be extended to their respective dimension-free versions, where the ambient dimension $d$ is replaced by the effective rank of some matrix that is an upper bound of the variance term. As a result, our results can be applied to linear operators in infinite-dimensional Hilbert spaces.
We apply our results in covariance estimation, PCA with Markovian data, and obtain improved bounds.
Notation.
Here and later, $a\wedge b:=\min\{a,b\}$, $a\vee b:=\max\{a,b\}$ for any $a,b\in{\mathbb{R}}$. We also denote ${\mathbb{R}}_{+}:=[0,\infty)$, ${\mathbb{N}}^{*}:=\{1,2,3,\ldots\}$. The symbol $\lesssim$” (resp. $\gtrsim$”) means no larger (resp. smaller) than up to a multiplicative universal constant, and $a\simeq b$ means $a\lesssim b$ and $a\gtrsim b$ hold simultaneously.
We denote by ${\bm{I}}_{d}\in{\mathbb{R}}^{d{\times}d}$ the $d\times d$ identity matrix, $\bm{0}_{d_{1}{\times}d_{2}}\in{\mathbb{R}}^{d_{1}{\times}d_{2}}$ the $d_{1}{\times}d_{2}$ zero matrix, $\left\|{{\bm{u}}}\right\|$ the Euclidean norm of vector $\bm{u}$}, $\left\|{{\bm{B}}}\right\|$ the spectral norm of matrix $\bm{B}$}, $\bm{B}$$^{*}$ the conjugate transpose of matrix $\bm{B}$}. For any $a\in{\mathbb{R}}$ and a self-adjoint matrix ${\bm{B}}\in{\mathbb{H}}^{d}$ (${\bm{B}}={\bm{B}}^{*}$) with spectral decomposition ${\bm{B}}=\sum_{i=1}^{d}\lambda_{i}{\bm{u}}_{i}{\bm{u}}_{i}^{*}$, we define ${\bm{B}}\wedge a:=\sum_{i=1}^{d}(\lambda_{i}\wedge a){\bm{u}}_{i}{\bm{u}}_{i}^{*}$. For any $\bm{B}_{1},\bm{B}_{2}\in{\mathbb{H}}^{d}$, $\bm{B}_{1}\preccurlyeq\bm{B}_{2}$ stands for $\bm{B}_{2}{-}\bm{B}_{1}$ is positive semi-definite (PSD), and we denote by ${\mathbb{H}}^{d}_{+}$ the set of all $d\times d$ PSD matrices.
We denote by $\mathds{1}$ the indicator function and $\left|{\xi}\right|$ the total variation measure of a signed measure $\xi$. For any $p\geq 1$ and a real-valued random variable $X$, we denote by $\|X\|_{p}=({\mathbb{E}}[|X|^{p}])^{1/p}$ the $L^{p}$ norm of $X$.
2 Main Results
In this section, we present our main results: the Rosenthal-Burkholder inequality for matrix martingales, as well as matrix Rosenthal, Hoeffding, and Bernstein inequalities for ergodic Markov chains. We will also give dimension-free versions of these results.
We find that the dimension-free results can be applied to linear operator-valued martingales in infinite-dimensional separable Hilbert spaces. For brevity, we will not change the notation, and our dimension-free results still hold if ${\mathbb{H}}^{d}$ is replaced by the space of all self-adjoint operators on the separable Hilbert space $\mathcal{X}$}, and ${\mathbb{C}}^{d_{1}{\times}d_{2}}$ is replaced by the space of all bounded linear operators from the separable Hilbert space ${\mathcal{X}}_{1}$ to the separable Hilbert space ${\mathcal{X}}_{2}$.
2.1 The Rosenthal-Burkholder Inequality for Matrix Martingales
Let $(\Omega,{\mathscr{F}},{\mathbb{P}})$ be a probability space and $\{\emptyset,\Omega\}={\mathscr{F}}_{-1}\subseteq{\mathscr{F}}_{0}\subseteq{\mathscr{F}}_{1}\subseteq\ldots\subseteq{\mathscr{F}}$ be the filtration. We denote by ${\mathbb{E}}_{k}$ the expectation conditioned on ${\mathscr{F}}_{k}$. We consider the $d$-dimensional self-adjoint matrix martingale $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ that is adapted to the filtration $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$. The matrix martingale $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ satisfies
${\mathbb{E}}_{k-1}{\bm{M}}_{k}={\bm{M}}_{k-1}\quad\text{ and }\quad{\mathbb{E}}\left\|{{\bm{M}}_{k}}\right\|<\infty\quad\text{ for }\ k\in{\mathbb{N}}.$
We denote by $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$ the martingale difference sequence, that is, ${\bm{X}}_{k}={\bm{M}}_{k}-{\bm{M}}_{k-1}\in{\mathbb{H}}^{d}$ (${\bm{M}}_{-1}:=\bm{0}_{d{\times}d}$). And we denote by $(\left\langle{{\bm{M}}}\right\rangle_{k})_{k\in{\mathbb{N}}}$ the quadratic variation sequence, that is, $\left\langle{{\bm{M}}}\right\rangle_{k}=\sum_{j=0}^{k}{\mathbb{E}}_{j-1}{\bm{X}}_{j}^{2}\in{\mathbb{H}}^{d}_{+}$. Throughout this section, we assume that $\left\langle{{\bm{M}}}\right\rangle_{\infty}:=\lim_{k\to\infty}\left\langle{{\bm{M}}}\right\rangle_{k}<\infty$ element-wisely almost surely. As a result, by the martingale convergence theorem (Durrett, 2019, Theorem 4.5.2), $\lim_{k\to\infty}{{\bm{M}}}_{k}={\bm{M}}_{\infty}\in{\mathbb{H}}^{d}$ almost surely.
Rectangular Case: For a $d_{1}{\times}d_{2}$ rectangular matrix martingale $({\bm{W}}_{k})_{k\in{\mathbb{N}}}$ adapted to $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$, we denote by ${\bm{D}}_{k}={\bm{W}}_{k}-{\bm{W}}_{k-1}$ the martingale difference and $\|\langle{\bm{W}}\rangle_{\infty}\|:=\|\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}{\bm{D}}_{k}^{*}]\|\vee\|\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}^{*}{\bm{D}}_{k}]\|$ the proxy of the quadratic variation. In this case, we also assume that
$\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}{\bm{D}}_{k}^{*}]<\infty\mbox{ and }\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}^{*}{\bm{D}}_{k}]<\infty\quad\mbox{element-wisely almost surely}.$
Dimension-Free: As mentioned in Minsker (2017), in the case of ultra-high dimensions or even infinite dimensions, it is desirable for the results not to depend on the dimension $d$ of the ambient space. To this end, we resort to the notion of effective rank (Koltchinskii and Lounici, 2016). For any ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$, its effective rank is defined as
$r({\bm{\Upsilon}})=\frac{\operatorname{tr}({\bm{\Upsilon}})}{\left\|{{\bm{\Upsilon}}}\right\|}.$
One can check that $r({\bm{\Upsilon}})\leq d$ and $r({\bm{\Upsilon}})$ can be much smaller than $d$ if the eigenvalues of $\bm{\Upsilon}$} rapidly decay to 0.
Now, we are ready to state one of the main results, the Rosenthal-Burkholder inequality for matrix martingales in the following theorem.
Theorem 2.1 (Rosenthal-Burkholder Inequality).
For any $p\geq 2$, suppose that
$\quad\Big{\|}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\Big{\|}_{p}<\infty\quad\text{ and }\quad\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|}\right\|_{p}<\infty.$
Then it holds that
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq C_{R,1}\sqrt{p\vee\log d}\Big{\|}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\Big{\|}_{p}+C_{R,2}(p\vee\log d)\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|}\right\|_{p},$ (6)
where $C_{R,1}=87$, $C_{R,2}=50$. When $p\geq 117$, the constants can be smaller: $C_{R,1}=64$, $C_{R,2}=28$.
Furthermore, the following extensions hold:
  1. (a)
    Rectangular Matrix: The above inequality also holds for the rectangular matrix martingale $({\bm{W}}_{k})_{k\in{\mathbb{N}}}$, in which we replace $\bm{M}$} by $\bm{W}$}, $\bm{X}$} by $\bm{D}$}, and $d$ by $d_{1}+d_{2}$.
  2. (b)
    Dimension-Free: In the self-adjoint case, if we further assume that there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that $\langle{\bm{M}}\rangle_{\infty}\preccurlyeq{\bm{\Upsilon}}$ almost surely, then the above inequality also holds when we replace $d$ by $e\cdot r({\bm{\Upsilon}}\wedge\|\|\left\langle{{\bm{M}}}\right\rangle_{\infty}\|\|_{\frac{p}{2}})$, and double the right-hand side.
    In the rectangular case (a), if we further assume there exist ${\bm{\Upsilon}}_{1}\in{\mathbb{H}}_{+}^{d_{1}}$, ${\bm{\Upsilon}}_{2}\in{\mathbb{H}}_{+}^{d_{2}}$ such that $\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}{\bm{D}}_{k}^{*}]\preccurlyeq{\bm{\Upsilon}}_{1}$ and $\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}^{*}{\bm{D}}_{k}]\preccurlyeq{\bm{\Upsilon}}_{2}$ almost surely, then the above inequality also holds, while we replace $\bm{M}$} by $\bm{W}$}, $\bm{X}$} by $\bm{D}$}, $d$ by $2[r({\bm{\Upsilon}}_{1}\wedge\|\|\left\langle{{\bm{W}}}\right\rangle_{\infty}\|\|_{\frac{p}{2}})\vee r({\bm{\Upsilon}}_{2}\wedge\|\|\left\langle{{\bm{W}}}\right\rangle_{\infty}\|\|_{\frac{p}{2}})]$, and double the right-hand side.
For self-adjoint matrix martingales $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$, the proof strategy is to combine the extrapolation method (good $\lambda$ inequality) (Hitczenko, 1990, 1994; Pinelis, 1994) to derive a scalar-version Rosenthal-Burkholder inequality, with techniques in matrix concentration inequalities (Tropp, 2011, 2012, 2015). A detailed outline of the proof can be found in Section 4.1.
For the rectangular matrix martingale $({\bm{W}}_{k})_{k\in{\mathbb{N}}}$, we can directly apply the previous results to the self-adjoint matrix martingale $({\mathscr{H}}({\bm{W}}_{k}))_{k\in{\mathbb{N}}}$ to obtain the desired inequality, where $\mathscr{H}$} is the Hermitian dilation operator (Paulsen, 2002). For a rectangular matrix $\bm{B}$}, its Hermitian dilation ${\mathscr{H}}\left({{\bm{B}}}\right)$ is a self-adjoint matrix, which satisfies
${\mathscr{H}}\left({{\bm{B}}}\right):=\begin{bmatrix}\bm{0}&{\bm{B}}\\ {\bm{B}}^{*}&\bm{0}\end{bmatrix},\quad{\mathscr{H}}\left({{\bm{B}}}\right)^{2}=\begin{bmatrix}{\bm{B}}{\bm{B}}^{*}&\bm{0}\\ \bm{0}&{\bm{B}}^{*}{\bm{B}}\end{bmatrix},\quad\left\|{{\mathscr{H}}\left({{\bm{B}}}\right)}\right\|=\left\|{{\bm{B}}}\right\|.$
It is worth noting that the assumption in Theorem (b), the quadratic variation $\langle{\bm{M}}\rangle_{\infty}$ has an almost sure upper bound $\bm{\Upsilon}$}, seems quite restrictive. However, such an assumption is exactly satisfied in deriving the dimension-free matrix Rosenthal inequality for uniformly geometrically ergodic Markov chains. We also note that doubling the r.h.s. in Theorem 2.1 (b) is not essential; we could reduce this multiplier from $2$ to a constant close to $1$ by increasing the effective rank term that only appears as the logarithmic term. However, for brevity, we use this presentation and do not strive for the optimal constants.
Discussion.
In the i.i.d. case, Chen et al. (2012); Mackey et al. (2014) provided a matrix Rosenthal inequality, and Jirak et al. (2024) further gave a dimension-free version. We extend their results from the i.i.d. setting to the martingale setting. See also Junge and Xu (2003, 2008); Junge and Zeng (2013); Jiao et al. (2025) for the Rosenthal-Burkholder inequality and other probability inequalities for noncommutative martingales.
We also present another version of the Rosenthal-Burkholder inequality: $\|\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|\|_{p}\lesssim[(p/\log p)\vee\log d](\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\|_{p}+\|\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\|_{p})$, which can be found in Section 4.1.
2.2 Matrix Rosenthal and Concentration Inequalities for Markov Chains
Consider the Markov kernel ${\mathrm{Q}}\colon{\mathcal{Z}}\times{\mathscr{Z}}\to[0,1]$ on the measurable space $({\mathcal{Z}},{\mathscr{Z}})$ with a unique stationary distribution $\pi\in{\mathcal{P}}$, where $\mathcal{P}$} is the set of all probability distributions on $({\mathcal{Z}},{\mathscr{Z}})$. For any $\xi\in{\mathcal{P}}$, we denote by ${\mathbb{P}}_{\xi}$ (resp. ${\mathbb{E}}_{\xi}$) the probability (resp. expectation) of the canonical process $(Z_{k})_{k\in{\mathbb{N}}}$ with initial distribution $\xi$, and $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$ the nature filtration where ${\mathscr{F}}_{k}=\sigma(Z_{0},\ldots,Z_{k})$. That is, $(Z_{k})_{k\in{\mathbb{N}}}$ is a stochastic process in $({\mathcal{Z}}^{\mathbb{N}},{\mathscr{Z}}^{\mathbb{N}},{\mathbb{P}}_{\xi})$ adapted to $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$, which satisfies ${\mathbb{P}}_{\xi}(Z_{0}\in A)=\xi(A)$, ${\mathbb{P}}_{\xi}(Z_{k+1}\in A\mid{\mathscr{F}}_{k})={\mathbb{P}}_{\xi}(Z_{k+1}\in A\mid Z_{k})={\mathrm{Q}}(Z_{k},A)$ ${\mathbb{P}}_{\xi}$-almost surely for any $A\in{\mathscr{Z}}$ and $k\in{\mathbb{N}}$. Thus, $(Z_{k})_{k\in{\mathbb{N}}}$ is an Markov chain.
For any $p\geq 1$, $\xi\in{\mathcal{P}}$ and random variable $X$ defined on $({\mathcal{Z}}^{\mathbb{N}},{\mathscr{Z}}^{\mathbb{N}},{\mathbb{P}}_{\xi})$, we denote by $\|X\|_{\xi,p}=({\mathbb{E}}_{\xi}[|X|^{p}])^{1/p}$ the $L^{p}$ norm of $X$ under the initial distribution $\xi$.
For any measurable function $f$ in $({\mathcal{Z}},{\mathscr{Z}})$ and distribution $\xi\in{\mathcal{P}}$, we denote $\left\|{f}\right\|_{\infty}=\sup_{z\in{\mathcal{Z}}}\left|{f(z)}\right|$ ($f$ is real-valued), $\xi(f)=\int_{\mathcal{Z}}f(z)\xi(dz)$, $(\xi{\mathrm{Q}})(\cdot)=\int_{{\mathcal{Z}}}{\mathrm{Q}}(z,\cdot)\xi(dz)$, and $({\mathrm{Q}}f)(\cdot)=\int_{{\mathcal{Z}}}f(z){\mathrm{Q}}(\cdot,dz)$.
Now, we introduce the ergodicity assumptions we consider in this paper, we will only consider the uniformly geometrically ergodic Markov chains in this section, and the more general geometrically $V$-ergodic Markov chains in Section 2.2.1. Under other ergodicity condition, such as geometrically ergodic Markov chains with respect to the Wasserstein semimetric (Durmus et al., 2023, Section 2.3), similar results can be derived by using the same techniques, so we omit them in this paper.
Assumption 2.1 (Uniformly Geometrically Ergodic Markov Chains).
There exists a $t_{\operatorname{mix}}\in{\mathbb{N}}$ such that for any $k\in{\mathbb{N}}$,
$\sup_{z\in{\mathcal{Z}}}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\pi}\right\|_{\textup{TV}}\leq 2\left({1/4}\right)^{\lfloor k/t_{\operatorname{mix}}\rfloor},$
where $\left\|{\xi}\right\|_{\textup{TV}}:=\left|{\xi}\right|({\mathcal{Z}})$ is the total variation norm of the signed measure $\xi$ on $({\mathcal{Z}},{\mathscr{Z}})$.
We consider a measurable self-adjoint matrix function ${\bm{F}}\colon{\mathcal{Z}}\to{\mathbb{H}}^{d}$. We assume $\pi({\bm{F}})=\int_{{\mathcal{Z}}}{\bm{F}}(z)\pi(dz)=\bm{0}_{d{\times}d}$ and $\|\|{\bm{F}}\|\|_{\infty}=\sup_{z\in{\mathcal{Z}}}\|{\bm{F}}(z)\|<\infty$. Our goal is to give counterparts to matrix Rosenthal, Hoeffding, Bernstein inequalities for the spectral norm of the summation ${\bm{S}}_{n}=\sum_{k=0}^{n-1}{{\bm{F}}}(Z_{k})$. We observe that the boundedness condition is indispensable. As argued by Fan et al. (2021, Theorem 4), even for scalar unbounded functions, it is impossible to obtain Hoeffding inequalities for ${\bm{S}}_{n}$ under Assumption 2.1, whereas our Rosenthal inequality implies the Hoeffding-typ inequality for ${\bm{S}}_{n}$.
Recall that the Markov chain CLT (Jones, 2004, Corollary 5) says that $n^{-1/2}{\bm{S}}_{n}$ converges weakly to a Gaussian distribution and satisfies $\lim_{n\to\infty}n^{-1}{\mathbb{E}}_{\xi}[{\bm{S}}_{n}^{2}]={\bm{\Sigma}}_{\pi}({\bm{F}})$ for any $\xi\in{\mathcal{P}}$, where ${\bm{\Sigma}}_{\pi}({\bm{F}})$ is the long-run variance defined as
${\bm{\Sigma}}_{\pi}({\bm{F}}):=\pi\left({{{\bm{F}}}^{2}}\right)+\sum_{k=1}^{\infty}\pi\left({{{\bm{F}}}{\mathrm{Q}}^{k}{{\bm{F}}}+\left({{\mathrm{Q}}^{k}{{\bm{F}}}}\right){{\bm{F}}}}\right).$
We aim to obtain a Rosenthal inequality whose leading term is given by the long-run variance ${\bm{\Sigma}}_{\pi}({\bm{F}})$, matching the Markov chain CLT.
Dimension-Free: When we address dimension-free results, we need to make an additional assumption, that is, there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$. Note that the effective rank $r({\bm{\Upsilon}})$ could be much smaller than the ambient dimension $d$.
Rectangular Case: In this case, we consider a rectangular matrix function ${\bm{R}}\colon{\mathcal{Z}}\to{\mathbb{C}}^{d_{1}{\times}d_{2}}$ that satisfies $\pi({\bm{R}})=\int_{{\mathcal{Z}}}{\bm{R}}(z)\pi(dz)=\bm{0}_{d_{1}{\times}d_{2}}$. We denote by ${\bm{V}}_{n}=\sum_{k=0}^{n-1}{\bm{R}}(Z_{k})$ the summation and
$\left\|{{\bm{\Sigma}}_{\pi}({\bm{R}})}\right\|:=\left\|{\pi\left({{\bm{R}}{\bm{R}}^{*}}\right){+}\sum_{k=1}^{\infty}\pi\left({{{\bm{R}}}{\mathrm{Q}}^{k}{{\bm{R}}^{*}}{+}\left({{\mathrm{Q}}^{k}{{\bm{R}}}}\right){{\bm{R}}^{*}}}\right)}\right\|\vee\left\|{\pi\left({{\bm{R}}^{*}{\bm{R}}}\right){+}\sum_{k=1}^{\infty}\pi\left({{{\bm{R}}^{*}}{\mathrm{Q}}^{k}{{\bm{R}}}{+}\left({{\mathrm{Q}}^{k}{{\bm{R}}^{*}}}\right){{\bm{R}}}}\right)}\right\|,$ (7)
called the proxy of the long-run variance.
Now we state one of our main results, the matrix Rosenthal inequality for uniformly geometrically ergodic Markov chains.
Theorem 2.2 (Matrix Rosenthal Inequality for Uniformly Geometrically Ergodic Markov Chains).
Under Assumption 2.1 and $\|\|{\bm{F}}\|\|_{\infty}<\infty$, we have that for any $p\geq 2$, $n\geq(p\vee\log d)t_{\operatorname{mix}}$, and $\xi\in{\mathcal{P}}$,
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $C_{R,1}\sqrt{\frac{(p\vee\log d)\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|}{n}}+D_{1}t_{\operatorname{mix}}^{1/2}\left({\frac{(p\vee\log d)t_{\operatorname{mix}}}{n}}\right)^{3/4}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty},$
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}+D_{2}\frac{pt_{\operatorname{mix}}}{n}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty},$ (8)
where $C_{R,1}$ is defined in Theorem 2.1 and $D_{1},D_{2}>0$ are universal constants, which can be traced in the proof.
Furthermore, the following extensions hold:
  1. (a)
    Rectangular Matrix: the above inequality also holds for the rectangular case if we further assume $\|\|{\bm{R}}\|\|_{\infty}<\infty$ and replace $\bm{S}$} by $\bm{V}$}, $\bm{F}$} by $\bm{R}$}, and $d$ by $d_{1}+d_{2}$.
  2. (b)
    Dimension-Free: in the self-adjoint case, if we further assume that there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$, then the above inequality also holds when we replace the $d$ in the first term by $e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\pi,{p}/{2}}])$, and replace the $d$ in the second term by the dimension factor defined after Eqn. (17), and double the right-hand side.
    In the rectangular case (a), under similar assumptions, analogous dimension-free results hold, and we omit the details.
The proof strategy of Theorem 2.2 is to leverage the Poisson equation ${\bm{G}}-\left({{\mathrm{Q}}{\bm{G}}}\right)={\bm{F}}$ to construct a matrix martingale ${\bm{M}}_{k}=\sum_{j=1}^{k}{\bm{G}}(Z_{j})-({\mathrm{Q}}{\bm{G}})(Z_{j-1})$, and then we apply our Rosenthal-Burkholder inequality (Theorem 2.1). This method is first introduced by Durmus et al. (2023), who derived a Hilbert space-valued Rosenthal inequality for Markov chains. A detailed outline of the proof can be found in Section 4.2.
For the dimension-free results, as $n\to\infty$, we can show that $d_{n,p}\to e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge\|{\bm{\Sigma}}_{\pi}({\bm{F}})\|)$, which is a natural dimension factor in this problem.
We note that the dependence of the second term in Theorem 2.2 on $t_{\operatorname{mix}}$ is suboptimal. If we want the first term to dominate the second term, we need to take $n\gtrsim(p\vee\log d)t_{\operatorname{mix}}^{3}$, which is not linear in $t_{\operatorname{mix}}$. In Durmus et al. (2023), they used dyadic descent argument to derive a Hilbert space-valued Rosenthal inequality with residual terms that have a sharp dependence on $t_{\operatorname{mix}}$. However, this argument does not work for the matrix case, because under the Löwner order, the square function is not operator monotone, and the $p$-th power function ($p>2$) is not operator convex, making the core inequality (23) in the dyadic descent argument of Durmus et al. (2023) invalid in our matrix case.
Fortunately, if we do not insist on obtaining the result with a leading term matching the Markov chain CLT, we can still derive results with the sharp dependence on $t_{\operatorname{mix}}$ by using the blocking argument. In the following theorem, we present another version of the Rosenthal inequality, whose leading term is given by the variance proxy $t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{\infty}^{2}$, which is an upper bound for the long-run variance $\|{\bm{\Sigma}}_{\pi}({\bm{F}})\|$.
Theorem 2.3.
Under Assumption 2.1 and $\|\|{\bm{F}}\|\|_{\infty}<\infty$, we have that for all $p\geq 2$ and $n\geq 2$,
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $\frac{16e}{3}C_{R,1}\sqrt{\frac{(p\vee\log d)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}^{2}}{n}}+\left({\frac{16e}{3}C_{R,2}(p\vee\log d)+\frac{19}{3}}\right)\frac{t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}}{n}.$
For brevity, we omit the inequality for the rectangular case, and defer the dimension-free results as well as the proofs to Lemma 4.3 (a).
By our Rosenthal inequalities and applying Markov inequality to $\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|^{\log(1/\delta)}$, we immediately obtain the following Hoeffding and Bernstein inequalities.
Corollary 2.1 (Matrix Hoeffding Inequalities for Uniformly Geometrically Ergodic Markov Chains).
Under Assumption 2.1 and $\|\|{\bm{F}}\|\|_{\infty}<\infty$, for any $\delta\in(0,1)$ and $n\geq 2$, with ${\mathbb{P}}_{\xi}$-probability at least $1-\delta$, it holds that
$\frac{1}{n}\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|\leq$ $\frac{16e}{3}C_{R,1}\sqrt{\frac{\log(\delta^{-1}{\vee}d)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}^{2}}{n}}+D_{3}\log(\delta^{-1}{\vee}d)\frac{t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}}{n},$
where $D_{3}>0$ is a universal constant, which can be taken as $6eC_{R,2}+eD_{2}$.
Corollary 2.2 (Matrix Bernstein Inequalities for Uniformly Geometrically Ergodic Markov Chains).
Under Assumption 2.1 and $\|\|{\bm{F}}\|\|_{\infty}<\infty$, for any $\delta\in(0,1)$ and $n\geq 2$, with ${\mathbb{P}}_{\xi}$-probability at least $1-\delta$, it holds that
$\frac{1}{n}\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|\leq$ $eC_{R,1}\sqrt{\frac{\log(\delta^{-1}{\vee}d)\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|}{n}}+e(D_{1}+D_{2})t_{\operatorname{mix}}^{1/2}\left({\frac{\log(\delta^{-1}{\vee}d)t_{\operatorname{mix}}}{n}}\right)^{3/4}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
For the above concentration inequalities, the inequalities for both the rectangular case and the dimension-free version (where one only needs to replace $p$ by $\log(1/\delta)$ in the dimension factor) still hold. We omit them here.
Discussion.
Garg et al. (2018); Qiu et al. (2020); Neeman et al. (2024); Wu et al. (2025) derived matrix concentration inequalities for Markov chains with a non-zero absolute spectral gap. Instead of the multi-matrix Golden-Thompson inequality (Sutter et al., 2017; Garg et al., 2018) and the León-Perron argument (León and Perron, 2004) used in these works, we leverage the Poisson equation approach (Durmus et al., 2023), which allows us to overcome several limitations of the previous results: First, our results can be applied to general non-reversible Markov chains without assuming a non-zero absolute spectral gap. Second, our results can be extended to more general ergodicity assumptions, under which we do not need to assume the boundedness of functions, as detailed in Section 2.2.1. Third, the leading term of our results matches the Markov chain CLT, rather than relying on a variance proxy, e.g. $t_{\operatorname{mix}}\|{\mathbb{E}}_{\pi}[{\bm{F}}^{2}(Z_{0})]\|$ or $t_{\operatorname{mix}}\|\|{\bm{F}}\|\|^{2}$. Finally, under further assumptions, we obtain dimension-free results, which the previous works, which used the multi-matrix Golden-Thompson inequality, fail to achieve them. Applying the dimension-free results, we can also obtain the concentration inequalities for linear operators in infinite-dimensional Hilbert spaces.
Our results have the following limitations: As mentioned before, the residual term in the results where the leading term matches the Markov chain CLT (Theorem 2.2, Theorem 2.2, Theorem 2.4) has a suboptimal dependence on $t_{\operatorname{mix}}$, failing to achieve the sharp dependence on $t_{\operatorname{mix}}$ in the residual term as in the Hilbert space-valued version of the inequality (Durmus et al., 2023). We leave the improvement to the sharp dependence on $t_{\operatorname{mix}}$ as future work. Another limitation is that the Poisson equation approach is only applicable to time-homogeneous functions (additive functionals of Markov chains), whereas the results in Neeman et al. (2024) apply to time-inhomogeneous functions. This makes our results inapplicable to problems such as stochastic approximation with Markovian data (Li et al., 2023; Mou et al., 2024; Huo et al., 2024; Samsonov et al., 2025). We also leave the extension to time-inhomogeneous functions as future work.
2.2.1 Extension: Results for Geometrically $V$-Ergodic Markov Chains
We now introduce the more general geometrically $V$-ergodic Markov chains (Meyn and Tweedie, 2012; Douc et al., 2018; Durmus et al., 2023) and present matrix Rosenthal inequalities under the ergodicity assumption. Let $V\colon{\mathcal{Z}}\to[e,\infty)$ be a Lyapunov function satisfying $\pi(V)<\infty$. We denote by ${\mathcal{P}}_{V}$ the set of all probability distributions whose elements $\xi$ satisfy $\xi(V)<\infty$. For any real-valued measurable function $f$, we also define its $V$-norm as $\|f\|_{V}=\sup_{z\in{\mathcal{Z}}}{\left|{f(z)}\right|}/{V(z)}$. We will see that, under the ergodicity assumption, for matrix functions $\bm{F}$} whose growth rate does not exceed $V^{1/p}$, the convergence of $n^{-1}{\bm{S}}_{n}$ in the $L^{p}$ norm can be obtained, which allows us to extend the result to unbounded functions. In this section we omit the results in the rectangular case.
Assumption 2.2 (Geometrically $V$-Ergodic Markov Chains).
There exist $t_{\operatorname{mix}}\in{\mathbb{N}}$ and $\varkappa\geq 1$ such that for any $k\in{\mathbb{N}}$,
$\sup_{z,z^{\prime}\in{\mathcal{Z}}}\frac{\left\|{{\mathrm{Q}}^{k}(z,\cdot)-{\mathrm{Q}}^{k}(z^{\prime},\cdot)}\right\|_{V}}{V(z)+V(z^{\prime})}\leq\varkappa\left({\frac{1}{4}}\right)^{\lfloor k/t_{\operatorname{mix}}\rfloor},$
where $\left\|{\xi}\right\|_{V}:=\left|{\xi}\right|(V)$ is the the $V$-total variation norm of the signed measure $\xi$ on $({\mathcal{Z}},{\mathscr{Z}})$.
Assumption 2.2 is equivalent to the statement that for any $\xi\in{\mathcal{P}}_{V}$ and $k\in{\mathbb{N}}$,
$\left\|{\xi{\mathrm{Q}}^{k}-\pi}\right\|_{V}\leq\varkappa\left({\xi(V)+\pi(V)}\right)\left({1/4}\right)^{\lfloor k/t_{\operatorname{mix}}\rfloor}.$
Hence, Assumption 2.1 is a special case of Assumption 2.2 with $V\equiv e$ and $\varkappa=1$. For more details on the total variation norm and the $V$-total variation norm, one can refer to Douc et al. (2018, Appendix D).
Now we are ready to present matrix Rosenthal inequalities for geometrically $V$-ergodic Markov chains. We first present the result parallel to Theorem 2.2 whose leading term matches the Markov chain CLT, with the proof in Section 4.2.
Theorem 2.4 (Matrix Rosenthal Inequality for Geometrically $V$-Ergodic Markov Chains).
Under Assumption 2.2 and $\|\|{\bm{F}}\|\|_{V^{1/p}}<\infty$ for $p\geq 2$, we have that for any $n\geq 2$, and $\xi\in{\mathcal{P}}_{V}$,
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $C_{R,1}\sqrt{\frac{(p\vee\log d)\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|}{n}}+D_{4}p^{3/2}\left({p\vee\log d}\right)^{3/4}t_{\operatorname{mix}}^{5/4}n^{-3/4}\left[{\varkappa\pi(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
$+D_{5}p^{3/2}\left({p\vee\log d}\right)t_{\operatorname{mix}}^{3/2}n^{-(1-1/p)}\left[{\varkappa\pi(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}},$
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}+\frac{32}{15n}\varkappa^{1/p}\left({\xi(V)+\pi(V)}\right)^{1/p}pt_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}},$
where $C_{R,1}$ is defined in Theorem 2.1 and $D_{4},D_{5}>0$ are universal constants, which can be traced in the proof.
The proof strategy is identical to that under uniform geometric ergodicity assumption. And the result also has a suboptimal dependence on $t_{\operatorname{mix}}$.
Next, we present a parallel result to Theorem 2.3, whose leading term is given by the variance proxy $t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{1/p}}^{2}$.
Theorem 2.5.
Under Assumption 2.2 and $\|\|{\bm{F}}\|\|_{V^{1/p}}<\infty$ for $p\geq 2$, we have that,
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $\frac{8}{3}C_{R,1}p\sqrt{\frac{(p\vee\log d)t_{\operatorname{mix}}}{n}}\left[{\varkappa\pi(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
$+\left({\frac{16}{3}C_{R,2}(p\vee\log d)+\frac{11}{3}}\right)\frac{pt_{\operatorname{mix}}}{n^{1-1/p}}\left[{\varkappa\pi(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$
The proof can be found in Lemma 4.3 (b).
We can also use Markov inequality to obtain high-probability upper bounds similar to Corollary 2.1 and Corollary $2.$ For brevity, we omit them here.
3 Applications
In this section, we apply our results to covariance estimation and principal component analysis (PCA) problems on Markovian data, to showcase their effectiveness.
We use the same notation as in Section 2.2. Suppose $(Z_{k})_{k\in{\mathbb{N}}}$ is a uniformly geometrically ergodic stationary Markov chain (Assumption 2.1) and $\bm{f}\colon{\mathcal{Z}}\to{\mathbb{R}}^{d}$ is a bounded vector-valued function. We first consider the problem of covariance estimation on Markovian data, which is also considered in Fan et al. (2021). That is, we are interested in analyzing the spectral norm error in estimating the covariance matrix ${\bm{\Sigma}}=\pi(\bm{f}\bm{f}^{\top})={\mathbb{E}}_{\pi}[\bm{f}(Z_{0})\bm{f}(Z_{0})^{\top}]$ using the sample covariance matrix $\hat{{\bm{\Sigma}}}_{n}=n^{-1}\sum_{k=0}^{n-1}\bm{f}(Z_{k})\bm{f}(Z_{k})^{\top}$. Let ${\bm{F}}(z):=\bm{f}(z)\bm{f}(z)^{\top}-{\bm{\Sigma}}$. Then $\bm{F}$} satisfies $\pi({\bm{F}})=\bm{0}_{d{\times}d}$ and $\|\|{\bm{F}}\|\|_{\infty}\leq\|\|\bm{f}\|\|_{\infty}^{2}$. Using the matrix Bernstein inequality for Markov chains (Corollary 2.2), we obtain that for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\frac{1}{n}\sup_{k\leq n}\left\|{\sum_{j=0}^{k-1}\left({\bm{f}(Z_{j})\bm{f}(Z_{j})^{\top}-{\bm{\Sigma}}}\right)}\right\|\leq C_{R,1}\sqrt{\frac{e^{2}\log(\delta^{-1}\vee d)\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|}{n}}+O\left({\frac{1}{n^{3/4}}}\right),$ (9)
where ${\bm{\Sigma}}_{\pi}({\bm{F}})$ is the long-run variance of $\bm{F}$} defined in Eqn. (2).
Next, we consider the problem of principal component analysis on Markovian data, which is also considered in Neeman et al. (2024); Kumar and Sarkar (2023). We denote by $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{d}\geq 0$ the eigenvalues of $\bm{\Sigma}$}, and ${\bm{v}}_{1},{\bm{v}}_{2},\ldots,{\bm{v}}_{n}$ the corresponding eigenvectors. We assume $\lambda_{1}-\lambda_{2}>0$. In the PCA problem, we are interested in analyzing the $\sin^{2}$ error in estimating the leading eigenvector ${\bm{v}}_{1}$ of $\bm{\Sigma}$} using the leading eigenvector $\hat{{\bm{v}}}_{1,n}$ of $\hat{{\bm{\Sigma}}}_{n}$. By Eqn. (9) and Wedin’s Theorem (Wedin, 1972), we immediately have
$1-\left({{\bm{v}}_{1}^{\top}\hat{{\bm{v}}}_{1,n}}\right)^{2}\leq C\frac{\log(\delta^{-1}\vee d)\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|}{(\lambda_{1}-\lambda_{2})^{2}n}+O\left({\frac{1}{n^{3/2}}}\right),$
where $C>0$ is a universal constant.
Under further assumptions on $\bm{f}$, we can obtain dimension-free results. Let ${\bm{\Upsilon}}\in{\mathbb{H}}_{d}^{+}$ be a real PSD matrix with spectral decomposition ${\bm{\Upsilon}}=\sum_{i=1}^{d}\mu_{i}{\bm{u}}_{i}{\bm{u}}_{i}^{\top}$ such that $\mu_{1}\geq\mu_{2}\geq\ldots\geq\mu_{d}\geq 0$, and $\bm{f}(z)=\sum_{i=1}^{d}\alpha_{i}(z){\bm{u}}_{i}$—the representation of $\bm{f}$ on the orthonormal basis $\{{\bm{u}}_{1},\ldots,{\bm{u}}_{d}\}$. Suppose that $\sup_{z\in{\mathcal{Z}}}\sum_{i=1}^{d}\alpha_{i}^{2}(z)/\mu_{i}\leq 1$ ($0/0:=0$ and $1/0:=\infty$). Then by the Schur complement lemma (Zhang, 2006), one has that $\bm{f}(z)\bm{f}(z)^{\top}\preccurlyeq{\bm{\Upsilon}}$ for any $z\in{\mathcal{Z}}$. Furthermore, ${\bm{F}}^{2}(z)=(\bm{f}(z)\bm{f}(z)^{\top}-{\bm{\Sigma}})^{2}\preccurlyeq 2[\bm{f}(z)\bm{f}(z)^{\top}]^{2}+2{\bm{\Sigma}}^{2}\preccurlyeq 2\|\|\bm{f}\|\|_{\infty}^{2}[\bm{f}(z)\bm{f}(z)^{\top}+{\bm{\Sigma}}]\preccurlyeq 4\|\|\bm{f}\|\|_{\infty}^{2}{\bm{\Upsilon}}$. Using the dimension-free results, we can replace the ambient $d$ in the bounds by the effective rank $e\cdot r([(64/9)t_{\operatorname{mix}}^{2}\|\|\bm{f}\|\|_{\infty}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\pi,(1/2)\log(1/\delta)}])$. Similar results hold for the more general geometrically $V$-ergodic Markov chains and unbounded function $\bm{f}$, and we omit them for brevity.
Compared with the results in Kumar and Sarkar (2023); Neeman et al. (2024), our results do not require the Markov chain to have an absolute spectral gap, while obtaining a leading term that matches the Markov chain CLT and gaining dimension-free results. Moreover, our results can be extended to more general ergodicity assumptions and unbounded functions. These aspects demonstrate the advantages of our results over the previous ones as we discussed earlier.
4 Proofs
In this section, we outline the proofs of our main results.
4.1 Proof of the Rosenthal-Burkholder Inequality for Matrix Martingales
Based on the Bennett inequality for matrix martingales (Lemma A.1), we can obtain the following good $\lambda$ inequality for conditionally symmetric matrix martingales, which is a matrix version of Pinelis (1994, Lemma 4.2). Before that, we need to introduce the concept of conditionally symmetric martingales.
Definition 4.1 (Conditionally Symmetric Martingales).
The martingale $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ is said to be conditionally symmetric if the conditional distributions of ${\bm{X}}_{k}$ and ${\bm{X}}_{k}$ given ${\mathscr{F}}_{k-1}$ coincide almost surely for any $k\in{\mathbb{N}}$.
With the assumption that the martingale $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ is conditionally symmetric, the analysis can be greatly simplified, because the truncated martingale remains a martingale. Later on, we will introduce how to perform symmetrization to extend the conclusion to general martingales.
Theorem 4.1 (Good $\lambda$ Inequality).
Suppose that $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ is conditionally symmetric. For any $\lambda,\delta_{1},\delta_{2}>0$ and $\beta>1+\delta_{2}$, letting $N=[\beta-(1+\delta_{2})]/{\delta_{2}}$, we have that
${\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|>\beta\lambda,\ \left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\leq\delta_{1}\lambda,\ \sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda}\right)\leq 2d\left({\frac{e}{N}\frac{\delta_{1}^{2}}{\delta_{2}^{2}}}\right)^{N}{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|>\lambda}\right).$
Furthermore, if there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that $\langle{\bm{M}}\rangle_{\infty}\preccurlyeq{\bm{\Upsilon}}$ almost surely, then the above inequality also holds in which $d$ is replaced by $e\cdot r({\bm{\Upsilon}}\wedge(\delta_{1}^{2}\lambda^{2}))$.
The proof is similar to Pinelis (1994, Lemma 4.2) and can be found in Appendix A.1. The key ingredient is the Bennett inequality for matrix martingales (Lemma A.1). Roughly speaking, this good $\lambda$ inequality indicates that if the supremum of a conditionally symmetric martingale is sufficiently large, then with high probability, both the quadratic variation and the supremum of the martingale difference will also be large.
We observe that the dimension-free factor $r\left({{\bm{\Upsilon}}\wedge(\delta_{1}^{2}\lambda^{2})}\right)$ in Theorem 4.1 increases to $d$ as $\lambda$ decreases to 0. Therefore, to obtain a valid dimension-free result, it is necessary to truncate the error level $\lambda$ before applying the good $\lambda$ inequality.
The following lemma, the truncated $\Phi$ inequality, translates a truncated good $\lambda$ inequality into a moment inequality, which is an extension of Burkholder (1973, Lemma 7.1).
Lemma 4.1 (Truncated $\Phi$ Inequality).
Suppose that $\Phi\colon{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$ is non-decreasing continuous satisfying $\Phi(\beta\lambda)\leq\gamma\Phi(\lambda)$ for all $\lambda\geq 0$, where $\beta>1$ and $\gamma>0$ are some constants; and $Y,Z$ are nonnegative random variables satisfying the following good $\lambda$ inequality
${\mathbb{P}}\left({Y>\beta\lambda,Z\leq\lambda}\right)\leq\epsilon{\mathbb{P}}\left({Y>\lambda}\right),\quad\forall\lambda\geq\frac{a}{\beta}$
for some constants $a\geq 0$ and $\epsilon\in(0,1/\gamma)$. Then we have that
${\mathbb{E}}\left[{\Phi(Y)}\right]\leq\frac{1}{1-\gamma\epsilon}\Phi(a)+\frac{\gamma}{1-\gamma\epsilon}{\mathbb{E}}\left[{\Phi(Z)}\right].$
The proof can be found in Appendix A.2.
Now, we are ready to give a full proof of the Rosenthal-Burkholder inequality (Theorem 2.1). As in the proof of the scalar Rosenthal-Burkholder inequality (Hitczenko, 1990; Pinelis, 1994), we first deal with conditionally symmetric martingales, and then generalize the results to general martingales.
Step 1: Conditionally Symmetric Martingales.
Assume that $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ is conditionally symmetric. We start with the dimension-dependent case. Applying the good $\lambda$ inequality (Theorem 4.1) and the truncated $\Phi$ inequality (Lemma 4.1) with $Y=\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|$, $Z=\delta_{1}^{-1}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}+\delta_{2}^{-1}\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|$, $\Phi(x)=x^{p}$, $\beta>1$ to be determined, $\gamma=\beta^{p}$, $a=0$, and
$\epsilon=2d\left({\frac{e}{N}\frac{\delta_{1}^{2}}{\delta_{2}^{2}}}\right)^{N},$
we have the following inequality
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq\left({\frac{1}{1-\gamma\epsilon}}\right)^{1/p}\beta\left({\delta_{1}^{-1}\left\|{\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}}\right\|_{p}+\delta_{2}^{-1}\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|}\right\|_{p}}\right),$
provided that
$\gamma\epsilon=2d\left({\frac{e}{N}\frac{\delta_{1}^{2}}{\delta_{2}^{2}}}\right)^{N}\beta^{p}<1,\quad\text{ where }\ N=\frac{\beta-(1+\delta_{2})}{\delta_{2}}.$ (10)
We can choose the following parameters such that $\beta>1+\delta_{2}$ and the inequality (10) holds:
$\beta=1+e^{-p/c}+\frac{1}{c\vee\log d},\quad\delta_{1}=\frac{1}{5\sqrt{c\vee\log d}\ e^{p/c}},\quad\delta_{2}=\frac{1}{5(c\vee\log d)},$
where $c\in[1,p]$ is a parameter to be determined. Now, we check that inequality (10) indeed holds. First note that $e^{-p/c}\geq e^{-1}$ and then
$N=\frac{\beta-1-\delta_{2}}{\delta_{2}}=4+5(c\vee\log d)e^{-p/c}\geq 4+5e^{-1}\log d,$
$\beta^{p}=\left[{1+(N+1)\delta_{2}}\right]^{p}\leq e^{2pN\delta_{2}}=e^{2pN/[5(c\vee\log d)]}\leq e^{2pN/(5c)},$
$\frac{e}{N}\frac{\delta_{1}^{2}}{\delta_{2}^{2}}=\frac{e}{4e^{2p/c}(c\vee\log d)^{-1}+5e^{p/c}}\leq\frac{e}{5}e^{-p/c},$
hence
$\gamma\epsilon=2d\left({\frac{e}{N}\frac{\delta_{1}^{2}}{\delta_{2}^{2}}}\right)^{N}\beta^{p}$ $\leq 2d\left({\frac{e}{5}}\right)^{4+5e^{-1}\log d}e^{-21p/(5c)}<2(e/5)^{4}e^{-21/5}<1,$
which is the desired conclusion. Here we used the fact $(e/5)^{5/e}<e^{-1}$ and $2(e/5)^{4}<1/2$. Then we have
$\left({\frac{1}{1-\gamma\epsilon}}\right)^{1/p}<\frac{1}{\sqrt{1-2(e/5)^{4}e^{-21/5}}}<1.002.$
Up to this point, we have shown that for the conditionally symmetric martingale $\{{\bm{M}}_{k}\}_{k\in{\mathbb{N}}}$, the following inequality holds:
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq 5.01\!\left({1{+}e^{-p/c}{+}\frac{1}{c{\vee}\log d}}\right)\left({\sqrt{c{\vee}\log d}\ e^{p{/}c}\left\|{\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1{/}2}}\right\|_{p}{+}(c{\vee}\log d)\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|}\right\|_{p}}\right).$ (11)
Setting $c=p$, we obtain the inequality in the form given in Theorem 2.1. Alternatively, we can choose $c$ as the unique solution to equation $\sqrt{c}=e^{p/c}$, which yields another common form of the Rosenthal-Burkholder inequality $\|\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|\|_{p}\lesssim[(p/\log p)\vee\log d](\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\|_{p}+\|\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\|_{p})$. We shall not delve into this form in further detail in this paper.
Dimension-Free: To extend the result to a dimension-free version, we only need to replace $a=0$ with
$\tilde{a}=\beta\left({\delta_{1}^{-1}\left\|{\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}}\right\|_{p}+\delta_{2}^{-1}\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|}\right\|_{p}}\right),$
and $d$ with
$\tilde{d}=e\cdot r\left({{\bm{\Upsilon}}\wedge\left[{\left\|{\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|}\right\|_{\frac{p}{2}}}\right]}\right).$
And we need to double the right-hand side of the inequality (11) due to the additional appearance of $\Phi(a)$, that is, if $\langle{\bm{M}}\rangle_{\infty}\preccurlyeq{\bm{\Upsilon}}$ almost surely, then it holds that
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq 10.02\!\left({1{+}e^{-p/c}{+}\frac{1}{c{\vee}\log\tilde{d}}}\right)\left({\sqrt{c{\vee}\log\tilde{d}}\ e^{p{/}c}\left\|{\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1{/}2}}\right\|_{p}{+}(c{\vee}\log\tilde{d})\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|}\right\|_{p}}\right).$
Step 2: Generalize to Arbitrary Martingales
In order to generalize the result under the conditionally symmetry assumption to arbitrary martingales, in the special case of independence increment, it suffices to introduce i.i.d. Rademacher random variables; while in the case of martingales, a more sophisticated technique, named decoupling tangent sequence (Kwapien and Woyczynski, 1989; Hitczenko, 1988; De la Pena and Giné, 2012), is required. The proof in this part is similar to those of the scalar Rosenthal-Burkholder inequality (Hitczenko, 1990), and we defer the presentation to Appendix A.3.
4.2 Proof of the Matrix Rosenthal Inequality for Markov Chains
We consider the Poisson equation with the matrix forcing function $\bm{F}$}:
${\bm{G}}-\left({{\mathrm{Q}}{\bm{G}}}\right)={\bm{F}}.$ (12)
In the following lemma, we show that ${\bm{G}}=\sum_{k=0}^{\infty}{\mathrm{Q}}^{k}{\bm{F}}$ is a solution to the Poisson equation under different assumptions, and provide several upper bounds for $\bm{G}$}. The proof can be found in Appendix B.1.
Lemma 4.2 (Solution to the Poisson Equation).
Consider the matrix Poisson equation (12). Then
  1. (a)
    Under Assumption 2.1 and $\|\|{\bm{F}}\|\|_{\infty}<\infty$, we have that ${\bm{G}}=\sum_{k=0}^{\infty}{\mathrm{Q}}^{k}{\bm{F}}$ is a solution to the Poisson equation and satisfies $\|\|{\bm{G}}\|\|_{\infty}\leq(8/3)t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{\infty}$. Furthermore, if there exists an ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$, then ${\bm{G}}^{2}(z)\preccurlyeq(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$.
  2. (b)
    Under Assumption 2.2 and $\|\|{\bm{F}}\|\|_{V^{\alpha}}<\infty$ for some $\alpha\in(0,1]$, we have that ${\bm{G}}=\sum_{k=0}^{\infty}{\mathrm{Q}}^{k}{\bm{F}}$ is a solution to the Poisson equation and satisfies $\|{\bm{G}}(z)\|\leq[\varkappa^{\alpha}2^{3-\alpha}/(3\alpha)](\pi(V)+V(z))^{\alpha}t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{\alpha}}$ for any $z\in{\mathcal{Z}}$.
Using the Poisson equation Eqn. (12), we decompose the summation into a martingale, which is the dominant term, plus a remainder term. That is,
${\bm{S}}_{n}=\sum_{k=0}^{n-1}{{\bm{F}}}(Z_{k})=\sum_{k=0}^{n-1}\left[{{\bm{G}}(Z_{k})-({\mathrm{Q}}{\bm{G}})(Z_{k})}\right]=\sum_{k=1}^{n}{\bm{X}}_{k}+{\bm{G}}(Z_{0})-{\bm{G}}(Z_{n})={\bm{M}}_{k}+{\bm{G}}(Z_{0})-{\bm{G}}(Z_{n}),$
where ${\bm{X}}_{k}={\bm{G}}(Z_{k})-({\mathrm{Q}}{\bm{G}})(Z_{k-1})$ is a $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$-adapted martingale difference sequence, and we denote ${\bm{M}}_{k}=\sum_{j=1}^{k}{\bm{X}}_{j}$ as the related martingale and $\langle{\bm{M}}\rangle_{k}=\sum_{j=1}^{k}{\mathbb{E}}_{j-1}[({\bm{X}}_{j})^{2}]$ as the quadratic variation. We define ${\bm{M}}_{0}={\bm{X}}_{0}=\bm{0}_{d\times d}$.
As in the proof of the Hilbert space-valued Rosenthal inequality for ergodic Markov chains (Durmus et al., 2023), we first deal with stationary Markov chains, and then generalize the results to non-stationary chains.
Step 1: Stationary Markov Chains.
We assume $(Z_{k})_{k\in{\mathbb{N}}}$ is a stationary Markov chain, namely $Z_{0}\sim\pi$. By Minkowski inequality and Lemma 4.2, for all $p\geq 2$, we have
$\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq\left\|{\sup_{k\leq n}\left\|{{\bm{M}}_{k}}\right\|}\right\|_{\pi,p}+2\left\|{\left\|{{\bm{G}}(Z_{0})}\right\|}\right\|_{\pi,p}.$ (13)
We start with the dimension-dependent results. For the martingale term, by the Rosenthal-Burkholder inequality (Theorem 2.1), we have
$\left\|{\sup_{k\leq n}\left\|{{\bm{M}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $C_{R,1}\sqrt{p\vee\log d}\Big{\|}\left\|{\left\langle{{\bm{M}}}\right\rangle_{n}}\right\|^{1/2}\Big{\|}_{\pi,p}+C_{R,2}(p\vee\log d)\left\|{\sup_{k\leq n}\|{\bm{X}}_{k}\|}\right\|_{\pi,p},$ (14)
where the quadratic variation term is dominant, because the second term is independent of $n$.
Now, we aim to give an upper bound for $\|\|\langle{\bm{M}}\rangle_{n}\|^{1/2}\|_{\pi,p}$. Note that $\langle{{\bm{M}}}\rangle_{n}=\sum_{k=1}^{n}{\mathbb{E}}_{k-1}[({\bm{X}}_{k})^{2}]$ and
${\mathbb{E}}_{k-1}\left[{\left({{\bm{X}}_{k}}\right)^{2}}\right]=$ ${\mathbb{E}}_{k-1}\left[{\left({{\bm{G}}(Z_{k})-({\mathrm{Q}}{\bm{G}})(Z_{k-1})}\right)^{2}}\right]$
$=$ $\left({{\mathrm{Q}}{\bm{G}}^{2}}\right)(Z_{k-1})-\left[{\left({{\mathrm{Q}}{\bm{G}}}\right)(Z_{k-1})}\right]^{2}$
$=$ ${\bm{H}}(Z_{k-1}),$
where we defined ${\bm{H}}(z)=({\mathrm{Q}}{\bm{G}}^{2})(z)-[({\mathrm{Q}}{\bm{G}})(z)]^{2}$ for any $z\in{\mathcal{Z}}$. According to the Poisson equation Eqn. (12) and the definition of ${\bm{\Sigma}}_{\pi}({\bm{F}})$ (Eqn. (2)), we have
$\pi\left({{\bm{H}}}\right)=$ $\pi\left({{\mathrm{Q}}{\bm{G}}^{2}-({\mathrm{Q}}{\bm{G}})^{2}}\right)$
$=$ $\pi\left({{\bm{G}}^{2}-\left({{\bm{G}}-{\bm{F}}}\right)^{2}}\right)$
$=$ $\pi\left({{{\bm{F}}}{\bm{G}}+{\bm{G}}{\bm{F}}-{{\bm{F}}}^{2}}\right)$
$=$ $\pi\left({{{\bm{F}}}^{2}}\right)+\sum_{k=1}^{\infty}\pi\left({{{\bm{F}}}{\mathrm{Q}}^{k}{{\bm{F}}}+\left({{\mathrm{Q}}^{k}{{\bm{F}}}}\right){{\bm{F}}}}\right)$
$=$ ${\bm{\Sigma}}_{\pi}({\bm{F}}).$
By Minkowski’s inequality and the inequality $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$, we have that
$\Big{\|}\left\|{\left\langle{{\bm{M}}}\right\rangle_{n}}\right\|^{1/2}\Big{\|}_{\pi,p}=$ $\sqrt{\left\|{\left\|{\sum_{k=1}^{n}{\bm{H}}(Z_{k-1})}\right\|}\right\|_{\pi,p/2}}$ (15)
$\leq$ $\sqrt{\left\|{n\left\|{\pi({\bm{H}})}\right\|+\left\|{\sum_{k=1}^{n}\left({{\bm{H}}(Z_{k-1})-\pi({\bm{H}})}\right)}\right\|}\right\|_{\pi,p/2}}$
$\leq$ $\sqrt{n\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|+\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\pi,p/2}}$
$\leq$ $\sqrt{n}\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|^{1/2}+\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\pi,p/2}^{1/2},$
where we denoted $\bar{{\bm{H}}}={{\bm{H}}}-\pi({\bm{H}})$.
Combining the upper bounds provided in Eqn. (13), Eqn. (14), Eqn. (15), we obtain
$\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $C_{R,1}\sqrt{n(p\vee\log d)}\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|^{1/2}+C_{R,1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\pi,p/2}^{1/2}$ (16)
$+C_{R,2}(p\vee\log d)\left\|{\sup_{k\leq n}\|{\bm{X}}_{k}\|}\right\|_{\pi,p}+2\left\|{\left\|{{\bm{G}}(Z_{0})}\right\|}\right\|_{\pi,p},$
where the first term is the leading term of order $\sqrt{n}$, and the second term is of order $n^{1/4}$ because it involves the square root of the sum of $n$ zero-mean functions of the Markov chain. The remaining two terms are independent of $n$. Thus, the sum of these three residual terms is of order $O(n^{1/4})$.
Dimension-Free: To get a dimension-free counterpart of the inequality (16) under Assumption 2.1, the only difference is that we need to apply the dimension-free Rosenthal-Burkholder inequality to get a dimension-free counterpart of the inequality (14). To this end, we need an upper bound for the quadratic variation:
$\langle{{\bm{M}}}\rangle_{n}=\sum_{k=1}^{n}{\bm{H}}(Z_{k-1})=\sum_{k=1}^{n}\left\{{\left({{\mathrm{Q}}{\bm{G}}^{2}}\right)(Z_{k-1})-\left[{\left({{\mathrm{Q}}{\bm{G}}}\right)(Z_{k-1})}\right]^{2}}\right\}\preccurlyeq\sum_{k=1}^{n}\left({{\mathrm{Q}}{\bm{G}}^{2}}\right)(Z_{k-1}).$
If we assume Assumption 2.1 and ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$, then by Lemma 4.2 (a), it holds that $\left({{\mathrm{Q}}{\bm{G}}^{2}}\right)(Z_{k-1})\preccurlyeq(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}$, thus $\langle{{\bm{M}}}\rangle_{n}\preccurlyeq(64/9)nt_{\operatorname{mix}}^{2}{\bm{\Upsilon}}$. By Theorem 2.1 (b), the inequality (14) still holds if we replace $d$ by $d_{n,p}=e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\pi,\frac{p}{2}}])$, and double the right-hand side. As $n\to\infty$, by ergodicity and the fact that $\pi({\bm{H}})={\bm{\Sigma}}_{\pi}({\bm{F}})$, we find that $d_{n,p}\to e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge\|{\bm{\Sigma}}_{\pi}({\bm{F}})\|)$, which is a natural dimension factor.
Residual Terms: Next, we provide the specific forms of the upper bounds for these three residual terms under different assumptions. For the term of order $n^{1/4}$, we need the following lemma, which is a crude version of the Rosenthal inequality for uniformly geometrically ergodic Markov chains.
Lemma 4.3.
  1. (a)
    Under Assumption 2.1 and $\|\|{\bm{F}}\|\|_{\infty}<\infty$, we have that for any $p\geq 2$,
    $\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $\frac{16}{3}C_{R,1}\sqrt{(p\vee\log d)nt_{\operatorname{mix}}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}+\left({\frac{16}{3}C_{R,2}(p\vee\log d)+\frac{19}{3}}\right)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty},$
    Furthermore, if there exists an ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$, then the above inequality also holds when we replace $d$ by $e\cdot r([(64/9){\bm{\Upsilon}}]\wedge[(1/\lfloor{n}/{t_{\operatorname{mix}}}\rfloor)\|\|\langle{{\bm{M}}^{(t_{\operatorname{mix}})}}\rangle_{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\|\|_{\pi,\frac{p}{2}}])$, and double the right-hand side, where ${\bm{M}}^{(t_{\operatorname{mix}})}$ is a matrix martingale defined in Appendix B.2.
  2. (b)
    Under Assumption 2.2 and $\|\|{\bm{F}}\|\|_{V^{1/p}}<\infty$ for $p\geq 2$, we have that,
    $\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $\frac{8}{3}C_{R,1}p\sqrt{(p\vee\log d)nt_{\operatorname{mix}}}\left[{\varkappa\pi(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
    $+\left({\frac{16}{3}C_{R,2}(p\vee\log d)+\frac{11}{3}}\right)pt_{\operatorname{mix}}n^{1/p}\left[{\varkappa\pi(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$
The proof can be found in Appendix B.2.
Under Assumption 2.1 and $\|\|{\bm{F}}\|\|_{\infty}<\infty$, by Lemma 4.2 (a) and Lemma 4.3 (a), $\|\|\bar{{\bm{H}}}\|\|_{\infty}\leq(64/9)t_{\operatorname{mix}}^{2}\|\|{{\bm{F}}}\|\|_{\infty}$ and the term of order $n^{1/4}$ has the following upper bound:
$C_{R,1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\pi,p/2}^{1/2}\lesssim\left({p\vee\log d}\right)^{3/4}t_{\operatorname{mix}}^{5/4}n^{1/4}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}+\left({p\vee\log d}\right)t_{\operatorname{mix}}^{3/2}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$ (17)
And it is easy to see that the other two residual terms are dominated by the term. If we further assume ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$, then by Lemma 4.2 (a), the above inequality still holds if we replace $d$ introduced in the l.h.s. by $e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\pi,\frac{p}{2}}])$, and $d$ introduced in the r.h.s. (see Lemma 4.3) by $e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/\lfloor{n}/{t_{\operatorname{mix}}}\rfloor)\|\|\langle{{\bm{M}}^{(t_{\operatorname{mix}})}_{{\bm{H}}}}\rangle_{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\|\|_{\pi,\frac{p}{4}}])$, and double the r.h.s., where the matrix martingale ${\bm{M}}^{(t_{\operatorname{mix}})}_{\bar{\bm{H}}}$ is defined in Appendix B.2 with $\bm{F}$} replaced by $\bar{\bm{H}}$.
Under Assumption 2.2 and $\|\|{\bm{F}}\|\|_{V^{1/p}}<\infty$, by Lemma 4.2 (b), $\|\|\bar{{\bm{H}}}\|\|_{V^{2/p}}^{1/2}$ has the following upper bound:
$\left\|{\left\|{\bar{{\bm{H}}}}\right\|}\right\|_{V^{2/p}}^{1/2}\leq$ $\left\|{\left\|{{{\bm{G}}}}\right\|^{2}}\right\|_{V^{2/p}}^{1/2}$
$=$ $\left\|{\left\|{{{\bm{G}}}}\right\|}\right\|_{V^{1/p}}$
$\leq$ $\sup_{z\in{\mathcal{Z}}}\frac{[p\varkappa^{1/p}2^{3-1/p}/3](\pi(V)+V(z))^{1/p}t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{1/p}}}{V^{1/p}(z)}$
$\lesssim$ $p\left[{\varkappa\pi(V)}\right]^{1/p}t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{1/p}},$
and by Lemma 4.3 (b) the term of order $n^{1/4}$ has the following upper bound:
$C_{R,1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\pi,p/2}^{1/2}\lesssim$ $p^{3/2}\left({p\vee\log d}\right)^{3/4}t_{\operatorname{mix}}^{5/4}n^{1/4}\left[{\varkappa\pi(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
$+p^{3/2}\left({p\vee\log d}\right)t_{\operatorname{mix}}^{3/2}n^{1/p}\left[{\varkappa\pi(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$
And it is easy to see that the other two residual terms are dominated by the term.
Step 2: Generalize to Non-Stationary Markov Chains
In order to generalize the result to non-stationary Markov chains, we need to leverage the maximal exact coupling argument (Douc et al., 2018, Lemma 19.3.6 and Lemma 19.3.9). Following (Durmus et al., 2023, Section 4), under Assumption 2.1 and $\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}<\infty$, we obtain
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}+\frac{2}{n}\left({1+128(t_{\operatorname{mix}}/\log 4)^{p}(p+1)^{p+1/2}e^{-p}}\right)^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
Under Assumption 2.2 and $\|\|{\bm{F}}\|\|_{V^{1/p}}<\infty$, we obtain
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}+\frac{32}{15n}\varkappa^{1/p}\left({\xi(V)+\pi(V)}\right)^{1/p}pt_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$
5 Conclusions
In this work, we have studied probability bounds for the spectral norm of sums of dependent random matrices. First, we have proved a Rosenthal-Burkholder inequality for matrix martingales using the extrapolation method (good $\lambda$ inequality) (Hitczenko, 1990, 1994; Pinelis, 1994) and matrix Bennett inequality (Tropp, 2011; Kroshnin and Suvorikova, 2024). Furthermore, combining this martingale inequality with the Poisson equation approach (Durmus et al., 2023), we have derived matrix Rosenthal inequalities for Markov chains under both the uniform geometric ergodicity assumption and the geometric $V$-ergodicity assumption. By Markov inequality, these moment inequalities can be further translated into concentration inequalities including Hoeffding and Bernstein inequalities. Through more careful analysis, we have also established dimension-free versions of the above inequalities.
Compared with previous results based on the León-Perron argument (León and Perron, 2004), our results yield a leading term that matches the Markov chain CLT, rather than relying on a variance proxy ($t_{\operatorname{mix}}\|{\mathbb{E}}_{\pi}[{\bm{F}}^{2}(Z_{0})]\|$ or $t_{\operatorname{mix}}\|\|{\bm{F}}\|\|^{2}$). And our results do not require the assumption of a non-zero absolute spectral gap. This expands the applicability to a broader range of problems, such as policy evaluation in reinforcement learning (Duan and Wainwright, 2022) and PCA (Kumar and Sarkar, 2023) with Markovian data. In contrast to prior works relying on the multi-matrix Golden-Thompson inequality (Sutter et al., 2017; Garg et al., 2018), our results can be extended to dimension-free versions, enabling applications to infinite-dimensional settings. Moreover, our approach is not restricted to the uniform geometric ergodicity assumption, so we can derive the results under more general ergodicity assumptions, and in some cases we do not require the functions to be bounded. The advantages of our results over previous works are demonstrated in the covariance estimation and PCA problems on Markovian data.
One direction for future work is to explore whether we can obtain residual terms with sharp dependence on $t_{\operatorname{mix}}$. This might require more sophisticated tools from matrix analysis, beyond the properties of self-adjoint matrices related to the Löwner order. Another promising direction is to extend our results to time-inhomogeneous functions while retaining the advantages mentioned above. For time-inhomogeneous functions, the standard Poisson equation approach no longer works, and more advanced tools from Markov chain theory would be necessary.
Appendix A Omitted Proof for the Rosenthal-Burkholder Inequality for Matrix Martingales
In this section, we complete the proofs omitted in Section 4.1.
A.1 Proof of Theorem 4.1
Proof.
We first recall the Bennett inequality for matrix martingales with bounded martingale difference, which is provided in Tropp [2011, Theorem 3.1] Kroshnin and Suvorikova [2024, Theorem 2.3].
Lemma A.1 (Bennett Inequality for Matrix Martingale).
Suppose that there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that $\langle{\bm{M}}\rangle_{\infty}\preccurlyeq{\bm{\Upsilon}}$ almost surely, and $\|{\bm{X}}_{k}\|\leq U$ for any $k\in{\mathbb{N}}$ almost surely for some $U>0$. Then for any $\varepsilon>0$, it holds that
${\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\left\|{{\bm{M}}_{k}}\right\|\geq\varepsilon}\right)\leq$ $2(d\wedge(e\cdot r({\bm{\Upsilon}})))\exp\left\{{-\frac{\left\|{{\bm{\Upsilon}}}\right\|}{U^{2}}h\left({\frac{U\varepsilon}{\left\|{{\bm{\Upsilon}}}\right\|}}\right)}\right\}\leq 2(d\wedge(e\cdot r({\bm{\Upsilon}})))\left({\frac{e\left\|{{\bm{\Upsilon}}}\right\|}{\varepsilon U}}\right)^{{\varepsilon}/{U}},$
where $h(x)=(1+x)\log(1+x)-x$ for any $x>0$.
For any $k\in{\mathbb{N}}$, let $\bar{{\bm{X}}}_{k}={\bm{X}}_{k}\mathds{1}\{\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda\}$ be the truncated version of ${\bm{X}}_{k}$, $\bar{{\bm{M}}}_{k}=\sum_{j=0}^{k}\bar{{\bm{X}}}_{j}$ be the truncated martingale, and $\langle\bar{{\bm{M}}}\rangle_{k}=\sum_{j=0}^{k}{\mathbb{E}}_{j-1}\bar{{\bm{X}}}_{j}^{2}\preccurlyeq\langle{\bm{M}}\rangle_{k}$ be the corresponding quadratic variation sequence. $\bar{{\bm{M}}}_{k}$ is indeed a martingale because
${\mathbb{E}}_{k-1}\left[{\bar{{\bm{X}}}_{k}}\right]=$ ${\mathbb{E}}_{k-1}\left[{{\bm{X}}_{k}\mathds{1}\{\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda\}}\right]$
$=$ ${\mathbb{E}}_{k-1}\left[{{{\bm{X}}}_{k}}\right]-{\mathbb{E}}_{k-1}\left[{{\bm{X}}_{k}\mathds{1}\{\|{\bm{X}}_{k}\|>\delta_{2}\lambda\}}\right]$
$=$ ${\mathbb{E}}_{k-1}\left[{{\bm{X}}_{k}\mathds{1}\{\|{\bm{X}}_{k}\|>\delta_{2}\lambda\}}\right]$
$=$ ${\mathbb{E}}_{k-1}\left[{{{\bm{X}}}_{k}}\right]-{\mathbb{E}}_{k-1}\left[{{\bm{X}}_{k}\mathds{1}\{\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda\}}\right]$
$=$ $-{\mathbb{E}}_{k-1}\left[{\bar{{\bm{X}}}_{k}}\right]$
$=$ $\bm{0}_{d\times d},$
where we used the fact that $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ is a martingale and it is conditionally symmetric. By definition of the truncated martingale, on the event $\{\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda\}$, we have ${\bm{M}}_{k}=\bar{{\bm{M}}}_{k}$ for any $k\in{\mathbb{N}}$.
Next, we define three stopping times: $\tau_{1}=\inf\{k\in{\mathbb{N}}\colon\|{{\bm{M}}_{k}}\|>\lambda\}$, $\tau_{2}=\inf\{k\in{\mathbb{N}}\colon\|{{\bm{M}}_{k}}\|>\beta\lambda\}$ ($\tau_{1}\leq\tau_{2}$), $\tau_{3}=\inf\{k\in{\mathbb{N}}\colon\|\langle\bar{{\bm{M}}}\rangle_{k+1}\|^{1/2}>\delta_{1}\lambda\}$. By definition of the stopping times, we have
$\left\{{\sup_{k\in{\mathbb{N}}}\|{{\bm{M}}}_{k}\|>\lambda}\right\}=\left\{{\tau_{1}<\infty}\right\},\quad\left\{{\sup_{k\in{\mathbb{N}}}\|{{\bm{M}}}_{k}\|>\beta\lambda}\right\}=\left\{{\tau_{2}<\infty}\right\},\quad\left\{{\left\|{\left\langle{\bar{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\leq\delta_{1}\lambda}\right\}=\left\{{\tau_{3}=\infty}\right\}.$ (18)
We also consider the stopped martingale $\hat{{\bm{M}}}_{k}=\bar{{\bm{M}}}_{(k\wedge\tau_{2}\wedge\tau_{3})\vee\tau_{1}}-\bar{{\bm{M}}}_{\tau_{1}}$ conditionally on the sigma-algebra of the $\tau_{1}$-past
${\mathscr{F}}_{\tau_{1}}=\left\{{A\in{\mathscr{F}}\colon A\cap\{\tau_{1}\leq k\}\in{\mathscr{F}}_{k},\quad\forall k\in{\mathbb{N}}}\right\}.$
That is $(\hat{{\bm{M}}}_{k})_{k\in{\mathbb{N}}}$ is a $({\mathscr{F}}_{k\vee\tau_{1}})_{k\in{\mathbb{N}}}$-adapted martingale with the martingale difference $\hat{{\bm{X}}}_{k}=\hat{{\bm{M}}}_{k}-\hat{{\bm{M}}}_{k-1}=\bar{{\bm{X}}}_{k}\mathds{1}\{\tau_{1}<k<\tau_{2}\wedge\tau_{3}\}$ upper-bounded by $\delta_{2}\lambda$ and square root of the quadratic variation upper-bounded by $\delta_{1}\lambda$.
Now, we are ready to derive the good $\lambda$ inequality.
${\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|>\beta\lambda,\ \left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\leq\delta_{1}\lambda,\ \sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda}\right)$ (19)
$\qquad\leq{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\|{{\bm{M}}}_{k}\|>\beta\lambda,\ \left\|{\left\langle{\bar{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\leq\delta_{1}\lambda,\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda}\right)$
$\qquad\leq{\mathbb{P}}\left({\tau_{2}<\infty,\tau_{3}=\infty,\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda}\right)$
$\qquad\leq{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\|\hat{{\bm{M}}}_{k}\|>(\beta-1-\delta_{2})\lambda}\right)$
$\qquad={\mathbb{E}}\left[{{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\|\hat{{\bm{M}}}_{k}\|>(\beta-1-\delta_{2})\lambda\Big{|}{\mathscr{F}}_{\tau_{1}}}\right)\mathds{1}\left\{{\tau_{1}<\infty}\right\}}\right]$
$\qquad\leq 2d\left({\frac{e}{N}\frac{\delta_{1}^{2}}{\delta_{2}^{2}}}\right)^{N}{\mathbb{E}}\left[{\mathds{1}\left\{{\tau_{1}<\infty}\right\}}\right]$
$\qquad=2d\left({\frac{e}{N}\frac{\delta_{1}^{2}}{\delta_{2}^{2}}}\right)^{N}{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|>\lambda}\right),$
which is the desired conclusion. Here, we explain some details in the derivation of Eqn (19). In the first inequality, we used the fact that $\langle\bar{{\bm{M}}}\rangle_{\infty}\preccurlyeq\langle{\bm{M}}\rangle_{\infty}$ almost surely. In the second inequality, we used Eqn. (18). In the third inequality, we used the fact that on the event $\{\tau_{2}<\infty,\tau_{3}=\infty,\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda\}$, it holds that $\sup_{k<\tau_{1}}\left\|{{{\bm{M}}}_{k}}\right\|=\sup_{k<\tau_{1}}\left\|{\bar{{\bm{M}}}_{k}}\right\|<\lambda$ and $\left\|{\bar{{\bm{M}}}_{(\tau_{2}\wedge\tau_{3})-1}}\right\|>\left\|{{{\bm{M}}}_{\tau_{2}}}\right\|-\left\|{{\bm{X}}_{\tau_{2}}}\right\|>(\beta-\delta_{2})\lambda$. And in the last inequality, we used the matrix Bennett inequality (Lemma A.1) with error term $(\beta-1-\delta_{2})\lambda$ parameter $U=\delta_{2}\lambda$ and $\sigma=\delta_{1}\lambda$.
If we further assume that there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that $\langle{\bm{M}}\rangle_{\infty}\preccurlyeq{\bm{\Upsilon}}$ almost surely, then the third inequality also holds if we replace $d$ with $e\cdot r({\bm{\Upsilon}}\wedge(\delta_{1}^{2}\lambda^{2}))$ by the dimension-free matrix Bennett inequality (Lemma A.1). ∎
A.2 Proof of Lemma 4.1
Proof.
Consider the Lebesgue-Stieltjes integral, we have
${\mathbb{E}}\left[{\Phi(Y)}\right]=$ ${\mathbb{E}}\left[{\Phi(Y)\mathds{1}\left\{{Y\leq a}\right\}}\right]+{\mathbb{E}}\left[{\Phi(Y)\mathds{1}\left\{{Y>a}\right\}}\right]$
$\leq$ $\Phi(a)+{\mathbb{E}}\left[{\Phi(Y)\mathds{1}\left\{{Y>a}\right\}}\right],$
and by the good $\lambda$ inequality, we further have
${\mathbb{E}}\left[{\Phi(Y)\mathds{1}\left\{{Y>a}\right\}}\right]=$ ${\mathbb{E}}\left[{\Phi(\beta\beta^{-1}Y)\mathds{1}\left\{{Y>a}\right\}}\right]$
$\leq$ $\gamma{\mathbb{E}}\left[{\Phi(\beta^{-1}Y)\mathds{1}\left\{{Y>a}\right\}}\right]$
$=$ $\gamma\int_{{a}/{\beta}}^{\infty}{\mathbb{P}}\left({Y>\beta\lambda}\right)\Phi(d\lambda)$
$=$ $\gamma\int_{{a}/{\beta}}^{\infty}\left[{{\mathbb{P}}\left({Y>\beta\lambda,Z\leq\lambda}\right)+{\mathbb{P}}\left({Y>\beta\lambda,Z>\lambda}\right)}\right]\Phi(d\lambda)$
$\leq$ $\gamma\int_{{a}/{\beta}}^{\infty}\left[{\epsilon{\mathbb{P}}\left({Y>\lambda}\right)+{\mathbb{P}}\left({Z>\lambda}\right)}\right]\Phi(d\lambda)$
$\leq$ $\gamma\epsilon{\mathbb{E}}\left[{\Phi(Y)\mathds{1}\left\{{Y>{a}/{\beta}}\right\}}\right]+\gamma{\mathbb{E}}\left[{\Phi(Z)}\right]$
$\leq$ $\gamma\epsilon{\mathbb{E}}\left[{\Phi(Y)\mathds{1}\left\{{Y>a}\right\}}\right]+\gamma\epsilon\Phi(a)+\gamma{\mathbb{E}}\left[{\Phi(Z)}\right]$
$\leq$ $\frac{\gamma\epsilon}{1-\gamma\epsilon}\Phi(a)+\frac{\gamma}{1-\gamma\epsilon}{\mathbb{E}}\left[{\Phi(Z)}\right].$
hence
${\mathbb{E}}\left[{\Phi(Y)}\right]\leq\frac{1}{1-\gamma\epsilon}\Phi(a)+\frac{\gamma}{1-\gamma\epsilon}{\mathbb{E}}\left[{\Phi(Z)}\right],$
which is the desired conclusion. ∎
A.3 Remaining Proof of Theorem 2.1
We continue from the Step 2 in Section 4.1. In this section, we only prove the dimension-dependent version of the inequality. For the dimension-free version, it can be obtained by repeating the proof line by line as presented in this section, and we omit it for brevity.
First, let us recall some important concepts. The first one is the tangent sequence, which is used to construct conditionally symmetric martingales. The second is the conditional independence (CI) condition; the additional random components introduced in constructing conditionally symmetric martingales will satisfy this condition as shown in Theorem A.2, and sequences satisfying the CI condition can be treated as independent sequences.
Definition A.1.
[Hitczenko, 1990, Tangent Sequences, Definition 2.1] Two $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$-adapted sequences $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$ and $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ are said to be tangent if the conditional distributions of ${\bm{X}}_{k}$ and ${\bm{Y}}_{k}$ given ${\mathscr{F}}_{k-1}$ coincide almost surely for any $k\in{\mathbb{N}}$.
From this definition, it follows that $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ is conditionally symmetric if and only if $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$ and $(-{\bm{X}}_{k})_{k\in{\mathbb{N}}}$ are tangent. Furthermore, suppose that there exists a sequence $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ tangent to $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$, and ${\bm{X}}_{k},{\bm{Y}}_{k}$ are ${\mathscr{F}}_{k-1}$-conditionally independent for any $k\in{\mathbb{N}}$, it holds that $(\sum_{j=0}^{k}({\bm{X}}_{j}-{\bm{Y}}_{j}))_{k\in{\mathbb{N}}}$ is a conditionally symmetric martingale.
Definition A.2.
[Hitczenko, 1990, Conditional Independence (CI) Condition, Definition 2.2] An $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$-adapted sequences $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ satisfies (CI) condition if there exists a $\sigma$-algebra $\mathscr{G}$$\subset{\mathscr{F}}$ such that the conditional distributions of ${\bm{Y}}_{k}$ given ${\mathscr{F}}_{k-1}$ and $\mathscr{G}$} coincide almost surely for any $k\in{\mathbb{N}}$, and the sequence $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ is $\mathscr{G}$}-conditionally independent.
The following lemma guarantees the existence of the desired sequence $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$.
Lemma A.2.
[Hitczenko, 1990, Lemma 2.3] In some enlarged probability space of $(\Omega,{\mathscr{F}},{\mathbb{P}})$, there exists a sequence $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ which is tangent to $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$, satisfies the (CI) condition with $\sigma$-algebra $\mathscr{G}$}, and satisfies ${\bm{X}}_{k},{\bm{Y}}_{k}$ are ${\mathscr{F}}_{k-1}$-conditionally independent for any $k\in{\mathbb{N}}$.
The statement and proof of this lemma require rigorous measure-theoretic language (such as regular conditional probabilities). To avoid introducing excessive notations, we omit the details and merely explain the idea of constructing $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ in words. Interested readers may refer to Hitczenko [1990, Lemma 2.3]. For simplicity, from now on, we shall not change the notations even though we will work in the enlarged probability space. At time 0, we generate an independent copy ${\bm{Y}}_{0}$ of ${\bm{X}}_{0}$ (in an enlarged probability space to accommodate the sequence $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$). At time $k\geq 1$, given ${\mathscr{F}}_{k-1}$, we generate a copy ${\bm{Y}}_{k}$ that is conditionally independent of ${\bm{X}}_{k}$. It is easy to see that $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ is tangent to $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$ and satisfies the (CI) condition with $\mathscr{G}$$\mathscr{F}$}.
Based on the sequence $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ given in Theorem A.2, we introduce some notations. For any $k\in{\mathbb{N}}$, we denote ${\bm{N}}_{k}=\sum_{j=0}^{k}{\bm{Y}}_{j}$, $\tilde{{\bm{X}}}_{k}={\bm{X}}_{k}-{\bm{Y}}_{k}$, $\tilde{{\bm{M}}}_{k}=\sum_{j=0}^{k}\tilde{{\bm{X}}}_{k}={\bm{M}}_{k}-{\bm{N}}_{k}$. Then $({\bm{N}}_{k})_{k\in{\mathbb{N}}}$ is a martingale with (CI) increments, and we denote by $(\langle{\bm{N}}\rangle_{k})_{k\in{\mathbb{N}}}$ the related quadratic variation process. In addition, $(\tilde{{\bm{M}}}_{k})_{k\in{\mathbb{N}}}$ is a conditionally symmetric martingale, and we denote by $(\langle\tilde{\bm{M}}\rangle_{k})_{k\in{\mathbb{N}}}$ the related quadratic variation process.
Because ${\bm{X}}_{k},{\bm{Y}}_{k}$ are ${\mathscr{F}}_{k-1}$-conditionally independent, we have $\langle\tilde{\bm{M}}\rangle_{k}=\langle{\bm{M}}\rangle_{k}+\langle{\bm{N}}\rangle_{k}$ almost surely. In addition, because $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ is tangent to $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$, we have $\langle{\bm{M}}\rangle_{k}=\langle{\bm{N}}\rangle_{k}$ almost surely, thus $\langle\tilde{\bm{M}}\rangle_{k}=2\langle{\bm{M}}\rangle_{k}$ almost surely.
By Minkowski’s inequality, we have
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq\left\|{\sup_{k\in{\mathbb{N}}}\|\tilde{\bm{M}}_{k}\|}\right\|_{p}+\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{N}}_{k}\|}\right\|_{p}.$
According to Rosenthal-Burkholder inequality for conditionally symmetric martingale (Eqn. (11)), we have
[…]
where
$\left\|{\left\|{\left\langle{\tilde{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}}\right\|_{p}=\sqrt{2}\left\|{\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}}\right\|_{p},$
and
$\left\|{\sup_{k\in{\mathbb{N}}}\left\|{\tilde{\bm{X}}_{k}}\right\|}\right\|_{p}\leq\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|}\right\|_{p}+\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{Y}}_{k}}\right\|}\right\|_{p}.$
To give an upper bound for $\|\sup_{k\in{\mathbb{N}}}\|{\bm{Y}}_{k}\|\|_{p}$, we need the following lemma.
Lemma A.3.
[Hitczenko, 1988, Lemma 1] Suppose that $A_{k},B_{k}\in{\mathscr{F}}_{k}$ satisfy ${\mathbb{P}}(A_{k}\mid{\mathscr{F}}_{k-1})={\mathbb{P}}(B_{k}\mid{\mathscr{F}}_{k-1})$ for any $k\in{\mathbb{N}}$. Then it holds that
${\mathbb{P}}\left({\bigcup_{k\in{\mathbb{N}}}A_{k}}\right)\leq 2{\mathbb{P}}\left({\bigcup_{k\in{\mathbb{N}}}B_{k}}\right).$
For any $t>0$, we define the events $A_{k}=\{\|{\bm{Y}}_{k}\|\geq t\}$, $B_{k}=\{\|{\bm{X}}_{k}\|\geq t\}$ which satisfy ${\mathbb{P}}(A_{k}\mid{\mathscr{F}}_{k-1})={\mathbb{P}}(B_{k}\mid{\mathscr{F}}_{k-1})$ because $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ is tangent to $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$. By Lemma A.3, we obtain
${\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\left\|{{\bm{Y}}_{k}}\right\|\geq t}\right)=$ ${\mathbb{P}}\left({\bigcup_{k\in{\mathbb{N}}}A_{k}}\right)\leq 2{\mathbb{P}}\left({\bigcup_{k\in{\mathbb{N}}}B_{k}}\right)=2{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|\geq t}\right),$
hence
$\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{Y}}_{k}}\right\|}\right\|_{p}^{p}=$ ${\mathbb{E}}\left[{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{Y}}_{k}}\right\|^{p}}\right]$
$=$ $\int_{0}^{\infty}{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\left\|{{\bm{Y}}_{k}}\right\|^{p}\geq u}\right)du$
$=$ $p\int_{0}^{\infty}{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\left\|{{\bm{Y}}_{k}}\right\|\geq t}\right)t^{p-1}dt$
$\leq$ $2p\int_{0}^{\infty}{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|\geq t}\right)t^{p-1}dt$
$=$ $2\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|}\right\|_{p}^{p},$
which implies $\|\sup_{k\in{\mathbb{N}}}\|{\bm{Y}}_{k}\|\|_{p}\leq 2^{1/p}\|\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\|_{p}$.
Therefore, we have
[…]
We only need to give an upper bound for $\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{N}}_{k}\|}\right\|_{p}$. Since $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ satisfies (CI) condition, the proof strategy is almost identical to that of an independent sum. The only difference is that we need to replace the expectation with the mathematical expectation conditional on $\mathscr{G}$}. The key tool is the following maximal version of the symmetrization inequality [Ledoux and Talagrand, 2013, Lemma 6.3].
Lemma A.4 (Maximal Symmetrization Inequality).
Let $F\colon{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$ be a non-decreasing convex function, and $({\mathcal{X}},\|\cdot\|)$ be a separable Banach space. Suppose that $(X_{k})_{k\in{\mathbb{N}}}$ is a sequence of independent $\mathcal{X}$}-valued mean zero random elements such that ${\mathbb{E}}[F(\sup_{k\in{\mathbb{N}}}\|\sum_{j=0}^{k}X_{j}\|)]<\infty$ for any $k\in{\mathbb{N}}$, and $(\epsilon_{k})_{k\in{\mathbb{N}}}$ is a sequence of i.i.d. Rademacher random variables (${\mathbb{P}}(\epsilon_{1}=1)={\mathbb{P}}(\epsilon_{1}=-1)=1/2$) independent of $(X_{k})_{k\in{\mathbb{N}}}$. Then it holds that
${\mathbb{E}}\left[{F\left({\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}X_{j}}\right\|}\right)}\right]\leq{\mathbb{E}}\left[{F\left({2\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}X_{j}}\right\|}\right)}\right].$
The proof can be found in Appendix A.4.
By applying Lemma A.4 with $F(x)=x^{p}$, we have
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{N}}_{k}\|}\right\|_{p}^{p}=$ ${\mathbb{E}}\left[{\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}{\bm{Y}}_{j}}\right\|^{p}}\right]$
$=$ ${\mathbb{E}}\left[{{\mathbb{E}}\left[{\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}{\bm{Y}}_{j}}\right\|^{p}\Bigg{|}{\mathscr{G}}}\right]}\right]$
$\leq$ $2^{p}{\mathbb{E}}\left[{{\mathbb{E}}\left[{\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}{\bm{Y}}_{j}}\right\|^{p}\Bigg{|}{\mathscr{G}}}\right]}\right]$
$=$ $2^{p}{\mathbb{E}}\left[{\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}{\bm{Y}}_{j}}\right\|^{p}}\right]$
$=$ $2^{p}\left\|{\sup_{k\in{\mathbb{N}}}\|\tilde{\bm{N}}_{k}\|}\right\|_{p}^{p},$
where we used the fact that $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ satisfies (CI) condition and we denoted $\tilde{\bm{N}}_{k}=\sum_{j=0}^{k}\epsilon_{j}{\bm{Y}}_{j}$ which is also a conditionally symmetric martingale with $\langle\tilde{\bm{N}}\rangle_{k}=\langle{\bm{N}}\rangle_{k}=\langle{\bm{M}}\rangle_{k}$ and $\sup_{k\in{\mathbb{N}}}\left\|{\epsilon_{k}{\bm{Y}}_{k}}\right\|=\sup_{k\in{\mathbb{N}}}\left\|{{\bm{Y}}_{k}}\right\|$ which implies $\|\sup_{k\in{\mathbb{N}}}\|\epsilon_{k}{\bm{Y}}_{k}\|\|_{p}=\|\sup_{k\in{\mathbb{N}}}\|{\bm{Y}}_{k}\|\|_{p}\leq 2^{1/p}\|\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\|_{p}$ for any $k\in{\mathbb{N}}$.
Applying the Rosenthal-Burkholder inequality for conditionally symmetric martingale (Eqn. (11)) again, we have
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{N}}_{k}\|}\right\|_{p}\leq$ $2\left\|{\sup_{k\in{\mathbb{N}}}\|\tilde{\bm{N}}_{k}\|}\right\|_{p}$ (21)
[…]
Combining Eqn. (20) with Eqn. (21), we arrive at the conclusion.
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq$ $\left\|{\sup_{k\in{\mathbb{N}}}\|\tilde{\bm{M}}_{k}\|}\right\|_{p}+\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{N}}_{k}\|}\right\|_{p}$
[…]
When we take $c=p$, we determine the constants in the inequality: The coefficient of the quadratic variation term is
$5.01(1+e^{-1}+\frac{1}{p\vee\log d})(2+\sqrt{2})e\leq 87,$
when $p\vee\log d\geq 117$, the constant can be taken as $64$. The coefficient of the martingale difference term is
$5.01(1+e^{-1}+\frac{1}{p\vee\log d})(1+3\cdot 2^{1/p})\leq 50,$
when $p\geq 117$, the constant can be taken as $28$.
A.4 Proof of Lemma A.4
Proof.
We denote by $(X_{k}^{\prime})_{k\in{\mathbb{N}}}$ an independent copy of $(X_{k})_{k\in{\mathbb{N}}}$, which is also independent of $(\epsilon_{k})_{k\in{\mathbb{N}}}$. And we denote $\tilde{X}_{k}=X_{k}-X_{k}^{\prime}$. Then $\epsilon_{k}\tilde{X}_{k}$ and $\tilde{X}_{k}$ have the same distribution. Hence
${\mathbb{E}}\left[{F\left({\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}X_{j}}\right\|}\right)}\right]=$ ${\mathbb{E}}\left[{F\left({\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\left({X_{j}-{\mathbb{E}}[X_{j}^{\prime}]}\right)}\right\|}\right)}\right]$
$\leq$ ${\mathbb{E}}\left[{F\left({{\mathbb{E}}\left[{{\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\tilde{X}_{j}}\right\|\Bigg{|}(X_{k}^{\prime})_{k\in{\mathbb{N}}}}}\right]}\right)}\right]$
$\leq$ ${\mathbb{E}}\left[{F\left({\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\tilde{X}_{j}}\right\|}\right)}\right]$
$=$ ${\mathbb{E}}\left[{F\left({\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}\tilde{X}_{j}}\right\|}\right)}\right]$
$\leq$ ${\mathbb{E}}\left[{F\left({\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}X_{j}}\right\|+\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}X_{j}^{\prime}}\right\|}\right)}\right]$
$\leq$ $\frac{1}{2}\left\{{{\mathbb{E}}\left[{F\left({2\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}X_{j}}\right\|}\right)}\right]+{\mathbb{E}}\left[{F\left({2\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}X_{j}^{\prime}}\right\|}\right)}\right]}\right\}$
$=$ ${\mathbb{E}}\left[{F\left({2\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}X_{j}}\right\|}\right)}\right],$
where we used the fact that the function $(x_{k})_{k\in{\mathbb{N}}}\mapsto\sup_{k\in{\mathbb{N}}}\left\|{x_{k}}\right\|$ from $\mathcal{X}$$^{\mathbb{N}}$ to $\mathbb{R}$$_{+}$ is a convex function and Jensen’s inequality. ∎
Appendix B Omitted Proof for the Matrix Rosenthal Inequality for Markov Chains
In this section, we complete the proofs omitted in Section 4.2.
B.1 Proof of Lemma 4.2
Proof.
To derive the desired conclusion, we need the following useful lemma, whose proof is in Appendix B.3.
Lemma B.1.
For any $\xi_{1},\xi_{2}\in{\mathcal{P}}$, vector-valued function ${\bm{h}}\colon{\mathcal{Z}}\to{\mathbb{C}}^{d}$, and matrix-valued function ${\bm{H}}\colon{\mathcal{Z}}\to{\mathbb{C}}^{d_{1}{\times}d_{2}}$, it holds that
  1. 1.
    Assume $\left\|{\left\|{{\bm{h}}}\right\|}\right\|_{\infty},\left\|{\left\|{{\bm{H}}}\right\|}\right\|_{\infty}<\infty$. Then it holds that
    $\left\|{\xi_{1}({\bm{h}})-\xi_{2}({\bm{h}})}\right\|\leq\left\|{\left\|{{\bm{h}}}\right\|}\right\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}},$
    $\left\|{\xi_{1}({\bm{H}})-\xi_{2}({\bm{H}})}\right\|\leq\left\|{\left\|{{\bm{H}}}\right\|}\right\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}}.$
  2. 2.
    Assume $\left\|{\left\|{{\bm{h}}}\right\|}\right\|_{V^{\alpha}},\left\|{\left\|{{\bm{H}}}\right\|}\right\|_{V^{\alpha}}<\infty$ for some $\alpha\in(0,1]$. Then it holds that
    $\left\|{\xi_{1}({\bm{h}})-\xi_{2}({\bm{h}})}\right\|\leq\left\|{\left\|{{\bm{h}}}\right\|}\right\|_{V^{\alpha}}\left\|{\xi_{1}-\xi_{2}}\right\|_{V^{\alpha}},$
    $\left\|{\xi_{1}({\bm{H}})-\xi_{2}({\bm{H}})}\right\|\leq\left\|{\left\|{{\bm{H}}}\right\|}\right\|_{V^{\alpha}}\left\|{\xi_{1}-\xi_{2}}\right\|_{V^{\alpha}}.$
Equipped with Lemma B.1, we are ready to prove this lemma.
  1. (a)
    By Douc et al. [2018, Proposition 21.2.3], we only need to show that $\|{\bm{G}}(z)\|<\infty$ for any $z\in{\mathcal{Z}}$.
    $\left\|{\left\|{{\bm{G}}}\right\|}\right\|_{\infty}=$ $\sup_{z\in{\mathcal{Z}}}\left\|{\sum_{k=0}^{\infty}\left({{\mathrm{Q}}^{k}{\bm{F}}}\right)(z)}\right\|$
    $\leq$ $\sup_{z\in{\mathcal{Z}}}\sum_{k=0}^{\infty}\left\|{\left({{\mathrm{Q}}^{k}{\bm{F}}}\right)(z)}\right\|$
    $=$ $\sup_{z\in{\mathcal{Z}}}\sum_{k=0}^{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)({\bm{F}})-\pi({\bm{F}})}\right\|$
    $\leq$ $\sup_{z\in{\mathcal{Z}}}\sum_{k=0}^{\infty}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\pi}\right\|_{\textup{TV}}$
    $\leq$ $\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}\sum_{k=0}^{\infty}\sup_{z\in{\mathcal{Z}}}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\pi}\right\|_{\textup{TV}}$
    $\leq$ $2\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}\sum_{k=0}^{\infty}\left({\frac{1}{4}}\right)^{\lfloor k/t_{\operatorname{mix}}\rfloor}$
    $\leq$ $2t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}\sum_{k=0}^{\infty}\left({\frac{1}{4}}\right)^{k}$
    $=$ $\frac{8}{3}t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
    where we used Lemma B.1 and Assumption 2.1.
    If we further assume ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$, for some ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$, then we have the following result: for any ${\bm{u}}\in{\mathbb{C}}^{d}$ and $z\in{\mathcal{Z}}$, it holds that
    $\left\|{{\bm{G}}(z){\bm{u}}}\right\|=$ $\left\|{\sum_{k=0}^{\infty}\left({{\mathrm{Q}}^{k}{\bm{F}}}\right)(z){\bm{u}}}\right\|$
    $\leq$ $\sum_{k=0}^{\infty}\left\|{\left({{\mathrm{Q}}^{k}{\bm{F}}}\right)(z){\bm{u}}}\right\|$
    $=$ $\sum_{k=0}^{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)({\bm{F}}{\bm{u}})-\pi({\bm{F}}{\bm{u}})}\right\|$
    $\leq$ $\sum_{k=0}^{\infty}\left\|{\left\|{{\bm{F}}{\bm{u}}}\right\|}\right\|_{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\pi}\right\|_{\textup{TV}}$
    $\leq$ $\frac{8}{3}t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}{\bm{u}}}\right\|}\right\|_{\infty},$
    where we used Lemma B.1. Hence
    ${\bm{u}}^{*}{\bm{G}}^{2}(z){\bm{u}}=$ $\left\|{{\bm{G}}(z){\bm{u}}}\right\|^{2}$
    $\leq$ $\frac{64}{9}t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}{\bm{u}}}\right\|}\right\|_{\infty}^{2}$
    $=$ $\frac{64}{9}t_{\operatorname{mix}}\sup_{z^{\prime}\in{\mathcal{Z}}}{\bm{u}}^{*}{\bm{F}}^{2}(z^{\prime}){\bm{u}}$
    $\leq$ $\frac{64}{9}t_{\operatorname{mix}}{\bm{u}}^{*}{\bm{\Upsilon}}{\bm{u}},$
    which is the desired conclusion.
  2. (b)
    We first state a lemma, which is a immediate corollary of Jensen’s inequality.
    Lemma B.2.
    [Durmus et al., 2023, Lemma 12] Assume Assumption $2.$ Then for any $\alpha\in(0,1]$, $k\in{\mathbb{N}}$, and $\xi\in{\mathcal{P}}_{V}$, it holds that
    $\left\|{\xi{\mathrm{Q}}^{k}-\pi}\right\|_{V^{\alpha}}\leq 2^{1-\alpha}\varkappa^{\alpha}\left({\pi(V)+\xi(V)}\right)^{\alpha}(1/4)^{{\lfloor k/t_{\operatorname{mix}}\rfloor}\alpha}.$
    The proof is almost identical:
    $\left\|{{\bm{G}}(z)}\right\|=$ $\left\|{\sum_{k=0}^{\infty}\left({{\mathrm{Q}}^{k}{\bm{F}}}\right)(z)}\right\|$
    $\leq$ $\sum_{k=0}^{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)({\bm{F}})-\pi({\bm{F}})}\right\|$
    $\leq$ $\sum_{k=0}^{\infty}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{\alpha}}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\pi}\right\|_{V^{\alpha}}$
    $\leq$ $2^{1-\alpha}\varkappa^{\alpha}\left({\pi(V)+V(z)}\right)^{\alpha}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{\alpha}}\sum_{k=0}^{\infty}(1/4)^{{\lfloor k/t_{\operatorname{mix}}\rfloor}\alpha}$
    $\leq$ $\left[{\varkappa^{\alpha}2^{3-\alpha}/(3\alpha)}\right](\pi(V)+V(z))^{\alpha}t_{\operatorname{mix}}\left\|{\|{\bm{F}}\|}\right\|_{V^{\alpha}},$
    where we used Lemma B.1, Lemma B.2, and the estimate in Durmus et al. [2023, Eqn (146)] in the last inequality.
B.2 Proof of Lemma 4.3
Proof.
  1. (a)
    By Minkowski inequality, we have
    $\left\|{\left\|{{\bm{S}}_{n}}\right\|}\right\|_{\pi,p}\leq\left\|{\left\|{\sum_{k=0}^{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor t_{\operatorname{mix}}-1}{{\bm{F}}}(Z_{k})}\right\|}\right\|_{\pi,p}+t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
    We define the function ${\bm{G}}_{t_{\operatorname{mix}}}\colon{\mathcal{Z}}\to{\mathbb{H}}^{d}$ as ${\bm{G}}_{t_{\operatorname{mix}}}=\sum_{k=0}^{\infty}{\mathrm{Q}}^{kt_{\operatorname{mix}}}{\bm{F}}$, which satisfies the Poisson equation ${\bm{G}}_{t_{\operatorname{mix}}}-{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}={\bm{F}}$ . By Lemma 4.2 (a), ${\bm{G}}_{t_{\operatorname{mix}}}$ satisfies the Poisson equation ${\bm{G}}_{t_{\operatorname{mix}}}-{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}={\bm{F}}$ and $\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}\leq({8}/{3})\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}$. By Minkowski inequality, we have
    $\left\|{\left\|{\sum_{k=0}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor t_{\operatorname{mix}}-1}{{\bm{F}}}(Z_{k})}\right\|}\right\|_{\pi,p}$ $=\left\|{\left\|{\sum_{r=0}^{t_{\operatorname{mix}}-1}\sum_{k=0}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor-1}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{kt_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\pi,p}$
    $\leq\sum_{r=0}^{t_{\operatorname{mix}}-1}\left\|{\left\|{\sum_{k=0}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor-1}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{kt_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\pi,p}$
    $\leq\sum_{r=0}^{t_{\operatorname{mix}}-1}\left\|{\left\|{\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\pi,p}+2t_{\operatorname{mix}}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}.$
    The summand of the first term is the spectral norm of a matrix martingale ${\bm{M}}^{(t_{\operatorname{mix}})}_{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}$, by Rosenthal-Burkholder inequality for matrix martingales (Theorem 2.1), we have
    $\left\|{\left\|{\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\pi,p}$
    $\leq$ $C_{R,1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}{\mathbb{E}}\left[{\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})}\right)^{2}\big{|}{{\mathscr{F}}_{r,k-1}}}\right]}\right\|^{1/2}}\right\|_{\pi,p}$
    $+C_{R,2}(p\vee\log d)\left\|{\sup_{k\leq\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\|{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})\|}\right\|_{\pi,p}$
    $\leq$ $C_{R,1}\sqrt{p\vee\log d}\Big{\|}\left({\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left({2\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}}\right)^{2}}\right)^{1/2}\Big{\|}_{\pi,p}+C_{R,2}(p\vee\log d)\left\|{2\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}}\right\|_{\pi,p}$
    $=$ $2C_{R,1}\sqrt{(p\vee\log d)\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}+2C_{R,2}(p\vee\log d)\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}.$
    where ${\mathscr{F}}_{r,k-1}=\sigma(Z_{jt_{\operatorname{mix}}+r},0\leq j\leq k-1)$. Combining the upper bounds above, we have
    $\left\|{\left\|{{\bm{S}}_{n}}\right\|}\right\|_{\pi,p}$ $\leq t_{\operatorname{mix}}(2C_{R,1}\sqrt{(p{\vee}\log d)\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}{+}2C_{R,2}(p{\vee}\log d)\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}){+}2t_{\operatorname{mix}}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}{+}t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}$
    $\leq\frac{16}{3}C_{R,1}\sqrt{(p\vee\log d)nt_{\operatorname{mix}}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}+\left({\frac{16}{3}C_{R,2}(p\vee\log d)+\frac{19}{3}}\right)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
    Furthermore, if there exists an ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$, then by Lemma 4.2 (a), it holds that ${\bm{G}}^{2}_{t_{\operatorname{mix}}}(z)\preccurlyeq(64/9){\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$. By Theorem 2.1 (b), the above inequality still holds if we replace $d$ with $e\cdot r([(64/9){\bm{\Upsilon}}]\wedge[(1/\lfloor{n}/{t_{\operatorname{mix}}}\rfloor)\|\|\langle{{\bm{M}}^{(t_{\operatorname{mix}})}}\rangle_{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\|\|_{\pi,\frac{p}{2}}])$, and double the right-hand side.
  2. (b)
    The proof is identical to that of part (a), except that we need to use Lemma 4.2 (b) instead.
B.3 Proof of Lemma B.1
Proof.
  1. 1.
    We first show the inequality for the vector-valued function. Recall the dual representation of the total variation [Douc et al., 2018, Proposition D.2.4]:
    $\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}}=\sup_{f\colon\|f\|_{\infty}\leq 1}(\xi_{1}(f)-\xi_{2}(f)),$
    then by the dual representation of the Euclidean norm, we have
    $\left\|{\xi_{1}({\bm{h}})-\xi_{2}({\bm{h}})}\right\|=$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}{\bm{u}}^{*}\left({\xi_{1}({\bm{h}})-\xi_{2}({\bm{h}})}\right)$
    $=$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}\left({\xi_{1}({\bm{u}}^{*}{\bm{h}})-\xi_{2}({\bm{u}}^{*}{\bm{h}})}\right)$
    $\leq$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}\|{\bm{u}}^{*}{\bm{h}}\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}}$
    $\leq$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}\|{\bm{u}}\|\left\|{\|{\bm{h}}\|}\right\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}}$
    $=$ $\left\|{\|{\bm{h}}\|}\right\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}},$
    which is desired.
    Now, we deal with the matrix case:
    $\left\|{\xi_{1}({\bm{H}})-\xi_{2}({\bm{H}})}\right\|=$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}\left\|{\left({\xi_{1}({\bm{H}})-\xi_{2}({\bm{H}})}\right){\bm{u}}}\right\|$
    $=$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}\left\|{\xi_{1}({\bm{H}}{\bm{u}})-\xi_{2}({\bm{H}}{\bm{u}})}\right\|$
    $\leq$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}\left\|{\|{\bm{H}}{\bm{u}}\|}\right\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}}$
    $\leq$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}\left\|{{\bm{u}}}\right\|\left\|{\|{\bm{H}}\|}\right\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}}$
    $=$ $\left\|{\|{\bm{H}}\|}\right\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}},$
    where we used the inequality for the vector-valued function.
  2. 2.
    The proof is identical to (a) if we note that
    $\xi_{1}(f)-\xi_{2}(f)=$ $\int_{\mathcal{Z}}f(z)(\xi_{1}-\xi_{2})(dz)$
    $=$ $\int_{\mathcal{Z}}\frac{f(z)}{V^{\alpha}(z)}V^{\alpha}(z)(\xi_{1}-\xi_{2})(dz)$
    $\leq$ $\int_{\mathcal{Z}}\frac{\left|{f(z)}\right|}{V^{\alpha}(z)}V^{\alpha}(z)\left|{\xi_{1}-\xi_{2}}\right|(dz)$
    $\leq$ $\sup_{z\in{\mathcal{Z}}}\frac{\left|{f(z)}\right|}{V^{\alpha}(z)}\int_{\mathcal{Z}}V^{\alpha}(z)\left|{\xi_{1}-\xi_{2}}\right|(dz)$
    $=$ $\left\|{f}\right\|_{V^{\alpha}}\left\|{\xi_{1}-\xi_{2}}\right\|_{V^{\alpha}}.$
    Then we only need to replace $\|\cdot\|_{\infty}$ by $\left\|{\cdot}\right\|_{V^{\alpha}}$, and $\|\cdot\|_{\textup{TV}}$ by $\left\|{\cdot}\right\|_{V^{\alpha}}$ in (a).