### v2

# Matrix Moment and Concentration Inequalities for Martingales and Ergodic Markov Chains with Applications in Statistical Learning

Abstract
In this paper, we study moment and concentration inequalities for the spectral norm of sums of dependent random matrices. We establish novel Rosenthal-Burkholder inequalities for discrete-time matrix local martingales, Burkholder-Davis-Gundy inequality for continuous matrix local martingales, as well as matrix Rosenthal, Hoeffding, and Bernstein inequalities for ergodic Markov chains. Compared with previous work on matrix concentration inequalities for Markov chains, which assume a non-zero absolute $L^{2}$-spectral gap or the stronger $\psi$-mixing condition, our results assume geometric ergodicity, a condition commonly used in statistical applications. Furthermore, our results have leading terms that match the Markov chain central limit theorem, rather than relying on suboptimal variance proxies. We also give dimension-free versions of the inequalities, which are independent of the ambient dimension $d$ and relies on the effective rank instead. This enables the generalization of our results to linear operators in infinite-dimensional Hilbert spaces. Our results have extensive applications in statistics and machine learning; in particular, we obtain improved bounds in covariance estimation and principal component analysis on Markovian data.
1 Introduction
Moment and concentration inequalities for the spectral norm of sums of random matrices (Wigderson and Xiao, 2005; Tropp, 2015) are of great significance for statistics, machine learning, theoretical computer science, signal processing, etc. Common examples include covariance matrix estimation (Chen et al., 2012), principal component analysis (PCA) (Kumar and Sarkar, 2023), and reinforcement learning (Duan and Wainwright, 2022; Wu et al., 2025; Mou, 2025). Although matrix probability inequalities have been studied extensively in the literature (Tropp, 2011, 2012; Mackey et al., 2014; Tropp, 2015; Paulin et al., 2016; Banna et al., 2016; Minsker, 2017; Tropp, 2018; Garg et al., 2018; Bacry et al., 2018; Klochkov and Zhivotovskiy, 2020; Qiu et al., 2020; Han and Li, 2020; Huang and Tropp, 2021; Cai et al., 2022; Bandeira et al., 2023; Zhivotovskiy, 2024; Neeman et al., 2024; Kroshnin and Suvorikova, 2024; Brailovskaya and van Handel, 2024; Jirak et al., 2025; Wang and Ramdas, 2024; Wu et al., 2025; Maitre, 2025; Van Werde and Sanders, 2025), to the best of our knowledge, the Rosenthal-Burkholder type inequalities for matrix martingales and matrix functions of Markov chains remain absent.
We would bridge the gaps in the current paper. Specifically, we first give a Rosenthal-Burkholder inequality (Rosenthal, 1970; Burkholder, 1973; Hitczenko, 1990) for discrete-time matrix martingales in Theorem 2.1. That is, for an $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$-adapted $d$-dimensional self-adjoint matrix martingale $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$, letting martingale differences ${\bm{X}}_{k}={\bm{M}}_{k}-{\bm{M}}_{k-1}$ and predictable quadratic variation $\left\langle{{\bm{M}}}\right\rangle_{\infty}=\sum_{k=0}^{\infty}{\mathbb{E}}[{\bm{X}}_{k}^{2}|{\mathscr{F}}_{k-1}]$, one has that
[…]
for any $p\geq 2$. Here, $\|\cdot\|$ is the spectral norm of the matrix, $\|\cdot\|_{p}$ is the $L^{p}$ norm of the random variable, and $C_{1},C_{2}>0$ are universal constants. This result matches the classic Hilbert space-valued Rosenthal-Burkholder inequality (Pinelis, 1994, Theorem 4.1), with the only difference that an additional $\log d$ term appears. And the proof is based on the extrapolation method (Theorem 4.1) (Hitczenko, 1990, 1994; Pinelis, 1994) and the techniques in the classic matrix concentration inequalities (Tropp, 2011, 2012, 2015; Kroshnin and Suvorikova, 2024).
As a direct application of our Rosenthal-Burkholder inequality, we further establish a Burkholder-Davis-Gundy inequality for continuous matrix local martingales $({\bm{M}}_{t})_{t\geq 0}$ in Theorem 2.2 by discretization, namely,
$\left\|{\sup_{0\leq t\leq\tau}\|{\bm{M}}_{t}\|}\right\|_{p}\leq C_{1}\sqrt{p\vee\log d}\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|^{1/2}\Big\|_{p},$ (2)
for any $p\geq 2$ and stopping time $\tau$ such that $\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|^{1/2}\|_{p}<\infty$.
Regarding Markov chains, based on the Rosenthal-Burkholder inequality in Eqn. (1), we further derive Rosenthal inequalities for sum of time-homogeneous matrix functions of ergodic Markov chains by using the approach of Poisson equations (see Eqn. (14)) (Durmus et al., 2023). In particular, we consider a stationary Markov chain $(Z_{k})_{k\in{\mathbb{N}}}$ in a measurable space $\mathcal{Z}$} with a Markov kernel $\mathrm{Q}$} and a stationary distribution $\mu$, and a square-integrable self-adjoint matrix function of ${\bm{F}}\in L^{2}_{\mu,0}({\mathbb{H}}^{d})$ with $\mu({\bm{F}}^{2}):=\int_{{\mathcal{Z}}}{\bm{F}}^{2}(z)\mu(dz)<\infty$ and $\mu({\bm{F}})=\bm{0}_{d\times d}$. Our goal is to give an upper bound for the spectral norm of the summation ${\bm{S}}_{n}:=\sum_{k=0}^{n-1}{{\bm{F}}}(Z_{k})$. By the Markov chain central limit theorem (CLT) (Jones, 2004), under some regularity conditions, it holds that $n^{-1/2}{\bm{S}}_{n}$ converges weakly to a random matrix with Gaussian entries and satisfies $\lim_{n\to\infty}n^{-1}{\mathbb{E}}[{\bm{S}}_{n}^{2}]={\bm{\Sigma}}_{\mu}({\bm{F}})$, where ${\bm{\Sigma}}_{\mu}({\bm{F}})$ is the long-run variance defined as
[…]
where $({\mathrm{Q}}{\bm{F}})(\cdot):=\int_{{\mathcal{Z}}}{\bm{F}}(z){\mathrm{Q}}(\cdot,dz)$. Under the uniform geometric ergodicity assumption (Assumption 2.1) and the geometric ergodicity with Lyapunov function $V$ assumption (Assumption 2.2), in Theorem 2.3 and Theorem 2.5 we show that for any $p\geq 2$, the following Rosenthal-type inequality holds:
[…]
where the leading term of the r.h.s. relies on the long-run variance ${\bm{\Sigma}}_{\mu}({\bm{F}})$, matching the Markov chain CLT. Throughout this paper, the $O$-notation captures only the order of $n$ as $n\to\infty$, with other constants omitted.
By the Markov inequality, in Corollary 2.1 and Corollary 2.2 we also translate the Rosenthal inequality (Eqn. (4)) into Hoeffding-type and Bernstein-type high-probability upper-bounds under the uniform geometric ergodicity assumption: for any $\delta\in(0,1)$, with probability at least $1-\delta$, it holds that
[…]
where $t_{\operatorname{mix}}$ is the mixing time, and $t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}^{2}$ is a variance proxy which is an upper bound of the long-run variance $\|{\bm{\Sigma}}_{\mu}({\bm{F}})\|$. Neeman et al. (2024, Theorem 2.6) gave a Bernstein inequality with a leading term relying on the variance proxy $\gamma^{-1}\|\mu({\bm{F}}^{2})\|$ which is also an upper bound of $\|{\bm{\Sigma}}_{\mu}({\bm{F}})\|$, where $\gamma$ is the absolute $L^{2}$-spectral gap. Hence, our results are sharper in the leading term. We apply our results in covariance estimation, PCA with Markovian data, and obtain improved bounds in these applications.
All of the above inequalities (1), (4), (5), (6) can be extended to their respective dimension-free versions, where the ambient dimension $d$ is replaced by the effective rank (see Eqn. (7)) of some matrix that is an upper bound of the variance term. As a result, our results can be applied to linear operators in infinite-dimensional Hilbert spaces.
Finally, we discuss other extensions of these results, which will not be elaborated on in this paper due to space constraints. First, with some efforts, we can prove using the same techniques that the inequality (4) also holds for geometrically ergodic Markov chains with respect to the Wasserstein semimetric (instead of total variation) (Hairer et al., 2011) as in (Durmus et al., 2023). This ergodicity assumption has applications in Markov chain Monte Carlo algorithms (Eberle, 2014; Hairer et al., 2014). Second, based on the Burkholder-Davis-Gundy inequality (2), we can extend the results for Markov chains to continuous Markov processes.
Related Work
Until now, probabilistic inequalities for matrix martingales have been extensively studied. For example, Tropp (2011) proposed a Freedman inequality for matrix martingales; Minsker (2017); Kroshnin and Suvorikova (2024) derived dimension-free matrix Freedman inequalities in different forms, respectively; and Bacry et al. (2018) established a Freedman inequality for continuous-time matrix martingales. Regarding Rosenthal-type moment inequalities, Chen et al. (2012) developed a sharp Rosenthal inequality in the independent case; Jirak et al. (2025) further extend the inequality to its dimension-free version; and Maitre (2025) established a version of Burkholder-Davis-Gundy inequality for continuous matrix martingales. Additionally, in the field of noncommutative probability (Junge and Xu, 2003, 2008; Junge and Zeng, 2013, 2015; Jiao et al., 2025), the probabilistic inequalities for noncommutative martingales proposed therein can also be translated into probabilistic inequalities for the Schatten $p$-norms of matrix martingales. However, the existing results do not imply ours; see the Discussion in Section 2.1 for details.
Regarding probability inequalities for Markov chains, some existing work (Garg et al., 2018; Qiu et al., 2020; Neeman et al., 2024; Wu et al., 2025; Van Werde and Sanders, 2025) has derived matrix concentration inequalities for Markov chains under various ergodicity assumptions. Here we elaborate on the differences between these existing results and ours in detail.
Garg et al. (2018); Qiu et al. (2020); Neeman et al. (2024); Wu et al. (2025) considered Markov chains with a non-zero absolute $L^{2}$-spectral gap (the gap between $1$ and the operator norm of $\mathrm{Q}$} over the subspace $L^{2}_{\mu,0}$), and they used the multi-matrix Golden-Thompson inequality (Sutter et al., 2017; Garg et al., 2018) and the León-Perron argument (León and Perron, 2004) in their proof. However, these studies suffer from the following limitations. First, they assumed a non-zero absolute $L^{2}$-spectral gap $\gamma\in(0,1]$, which is a strong ergodicity condition and implies geometric ergodicity with some Lyapunov function (Assumption 2.2) if we assume the Markov chain is $\psi$-irreducible and aperiodic (Kontoyiannis and Meyn, 2012, Theorem 1.3). This assumption does not hold in many examples, especially for non-reversible Markov chains. For example, Huang and Li (2025, Example A.4) provided a finite-state, non-reversible, irreducible, and aperiodic Markov chain; due to the finiteness of its state space, this Markov chain is uniformly geometrically ergodic (Assumption 2.1), but its absolute $L^{2}$-spectral gap is zero. This strong assumption renders these results inapplicable to general non-reversible Markov chains, which arise in reinforcement learning problems with Markovian data (Duan and Wainwright, 2022). Second, the leading terms of these results rely on suboptimal variance proxies $\gamma^{-1}\|\mu({\bm{F}}^{2})\|$ instead of long-run variance $\|{\bm{\Sigma}}_{\mu}({\bm{F}})\|$. Finally, the multi-matrix Golden-Thompson inequality (Garg et al., 2018, Theorem 1.5) used in their proof incurs an additional dependence on the dimension $d$, making it impossible to generalize these results to a dimension-free version. Accordingly, these results cannot be easily extended to linear operators in infinite-dimensional Hilbert spaces.
In one recent work, Van Werde and Sanders (2025) considered $\psi$-mixing Markov chains and employed the universality-based approach (Bandeira et al., 2023; Brailovskaya and van Handel, 2024) in their proof. Compared with the classic matrix concentration inequalities derived in our paper, the universality-based concentration inequalities they obtained are sharper in cases involving matrices with specific structures, especially when matrices in the summation are far from being commutative. In their upper bounds, the coefficient of the long-run variance $\|{\bm{\Sigma}}_{\mu}({\bm{F}})\|$ no longer includes $\log d$. Instead, $\log d$ appears in a new variance term, which is a lower bound of $\|{\bm{\Sigma}}_{\mu}({\bm{F}})\|$. For a detailed introduction of universality-based concentration inequalities, one may refer to the seminal work (Bandeira et al., 2023; Brailovskaya and van Handel, 2024). However, the results in (Van Werde and Sanders, 2025) have the following limitations. On the one hand, $\psi$-mixing is a stronger ergodicity assumption than a non-zero absolute $L^{2}$-spectral gap and uniform geometric ergodicity (Assumption 2.1). Markov chains with infinite states generally fail to satisfy this condition, which limits the scope of application of the results. On the other hand, the approach they employed remains difficult to extend to infinite-dimensional Hilbert spaces.
Notation.
Here and later, $a\wedge b:=\min\{a,b\}$, $a\vee b:=\max\{a,b\}$ for any $a,b\in{\mathbb{R}}$. We also denote ${\mathbb{R}}_{+}:=[0,\infty)$, ${\mathbb{N}}^{*}:=\{1,2,3,\ldots\}$. The symbol $\lesssim$” (resp. $\gtrsim$”) means no larger (resp. smaller) than up to a multiplicative universal constant, and $a\simeq b$ means $a\lesssim b$ and $a\gtrsim b$ hold simultaneously. We denote by $\mathds{1}$ the indicator function and $\left|{\xi}\right|$ the total variation measure of a signed measure $\xi$. For any $p\geq 1$ and a real-valued random variable $X$, we denote by $\|X\|_{p}={\mathbb{E}}^{1/p}[|X|^{p}]$ the $L^{p}$ norm of $X$.
We denote by ${\bm{I}}_{d}\in{\mathbb{R}}^{d{\times}d}$ the $d\times d$ identity matrix, $\bm{0}_{d_{1}{\times}d_{2}}\in{\mathbb{R}}^{d_{1}{\times}d_{2}}$ the $d_{1}{\times}d_{2}$ zero matrix, $\left\|{{\bm{u}}}\right\|$ the Euclidean norm of vector $\bm{u}$}, $\left\|{{\bm{B}}}\right\|=\sup_{{\bm{u}}\colon\|{\bm{u}}\|\leq 1}\|{\bm{B}}{\bm{u}}\|$ the spectral norm of matrix $\bm{B}$}, $\bm{B}$$^{*}$ the conjugate transpose of matrix $\bm{B}$}, and ${\bm{B}}]_{ij}$ the $(i,j)$-entry of matrix $\bm{B}$}. For any self-adjoint matrix ${\bm{B}}\in{\mathbb{H}}^{d}$ (${\bm{B}}={\bm{B}}^{*}$) with spectral decomposition ${\bm{B}}=\sum_{i=1}^{d}\lambda_{i}{\bm{u}}_{i}{\bm{u}}_{i}^{*}$, we denote by $\operatorname{Tr}{{\bm{B}}}=\sum_{i=1}^{d}\lambda_{i}$ the trace of matrix $\bm{B}$}, and we define $|{\bm{B}}|:=\sum_{i=1}^{d}|\lambda_{i}|{\bm{u}}_{i}{\bm{u}}_{i}^{*}$ , ${\bm{B}}\wedge a:=\sum_{i=1}^{d}(\lambda_{i}\wedge a){\bm{u}}_{i}{\bm{u}}_{i}^{*}$ for any $a\in{\mathbb{R}}$. For any $\bm{B}_{1},\bm{B}_{2}\in{\mathbb{H}}^{d}$, $\bm{B}_{1}\preccurlyeq\bm{B}_{2}$ stands for $\bm{B}_{2}{-}\bm{B}_{1}$ is positive semi-definite (PSD), and we denote by ${\mathbb{H}}^{d}_{+}$ the set of all $d\times d$ PSD matrices.
Organization.
The remainder of this paper is organized as follows. In Section 2, we present our main results. In Section 3, we apply our results in covariance estimation, PCA problems with Markovian data, and obtain improved error bounds. The proof is outlined in Section 4. In Section 5, we conclude our work. Details of the proof are given in the appendices.
2 Main Results
In this section, we present our main results: the Rosenthal-Burkholder inequality for discrete-time matrix martingales and Burkholder-Davis-Gundy inequality for continuous matrix local martingales in Section 2.1, as well as matrix Rosenthal, Hoeffding, and Bernstein inequalities for ergodic Markov chains in Section 2.2. We will also give dimension-free versions of these results.
We find that the dimension-free results can be applied to linear operator-valued martingales in infinite-dimensional separable Hilbert spaces. For brevity, we will not change the notation, and our dimension-free results still hold if ${\mathbb{H}}^{d}$ is replaced by the space of all self-adjoint operators in the separable Hilbert space $\mathcal{X}$}, and ${\mathbb{C}}^{d_{1}{\times}d_{2}}$ is replaced by the space of all bounded linear operators from the separable Hilbert space ${\mathcal{X}}_{1}$ to the separable Hilbert space ${\mathcal{X}}_{2}$.
2.1 The Rosenthal-Burkholder Inequality for Matrix Martingales
Let $(\Omega,{\mathscr{F}},{\mathbb{P}})$ be a probability space and $\{\emptyset,\Omega\}={\mathscr{F}}_{-1}\subseteq{\mathscr{F}}_{0}\subseteq{\mathscr{F}}_{1}\subseteq\ldots\subseteq{\mathscr{F}}$ be the filtration. We denote by ${\mathbb{E}}_{k}$ the expectation conditioned on ${\mathscr{F}}_{k}$. We consider the $d$-dimensional self-adjoint matrix martingale $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ that is adapted to the filtration $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$. The matrix martingale $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ satisfies
${\mathbb{E}}_{k-1}{\bm{M}}_{k}={\bm{M}}_{k-1}\quad\text{ and }\quad{\mathbb{E}}\left\|{{\bm{M}}_{k}}\right\|<\infty\quad\text{ for }\ k\in{\mathbb{N}}.$
We denote by $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$ the martingale difference sequence, that is, ${\bm{X}}_{k}={\bm{M}}_{k}-{\bm{M}}_{k-1}\in{\mathbb{H}}^{d}$ (${\bm{M}}_{-1}:=\bm{0}_{d{\times}d}$). And we denote by $(\left\langle{{\bm{M}}}\right\rangle_{k})_{k\in{\mathbb{N}}}$ the predictable quadratic variation sequence, that is, $\left\langle{{\bm{M}}}\right\rangle_{k}=\sum_{j=0}^{k}{\mathbb{E}}_{j-1}{\bm{X}}_{j}^{2}\in{\mathbb{H}}^{d}_{+}$. Throughout this section, we assume that $\left\langle{{\bm{M}}}\right\rangle_{\infty}:=\lim_{k\to\infty}\left\langle{{\bm{M}}}\right\rangle_{k}<\infty$ element-wisely almost surely. As a result, by the martingale convergence theorem (Durrett, 2019, Theorem 4.5.2), $\lim_{k\to\infty}{{\bm{M}}}_{k}={\bm{M}}_{\infty}\in{\mathbb{H}}^{d}$ almost surely.
Rectangular Case: For a $d_{1}{\times}d_{2}$ rectangular matrix martingale $({\bm{W}}_{k})_{k\in{\mathbb{N}}}$ adapted to $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$, we denote by ${\bm{D}}_{k}={\bm{W}}_{k}-{\bm{W}}_{k-1}$ the martingale difference and $\|\langle{\bm{W}}\rangle_{\infty}\|:=\|\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}{\bm{D}}_{k}^{*}]\|\vee\|\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}^{*}{\bm{D}}_{k}]\|$ the proxy of the quadratic variation. In this case, we also assume that
$\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}{\bm{D}}_{k}^{*}]<\infty\mbox{ and }\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}^{*}{\bm{D}}_{k}]<\infty\quad\mbox{element-wisely almost surely}.$
Dimension-Free: As mentioned in Minsker (2017), in the case of ultra-high dimensions or even infinite dimensions, it is desirable that the results do not depend on the dimension $d$ of the ambient space. To this end, we resort to the notion of effective rank (Koltchinskii and Lounici, 2016). For any ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$, its effective rank is defined as
[…]
It is easily seen that $r({\bm{\Upsilon}})\leq d$ and $r({\bm{\Upsilon}})$ can be much smaller than $d$ if the eigenvalues of $\bm{\Upsilon}$} rapidly decay to 0.
Now, we are ready to state one of the main results, the Rosenthal-Burkholder inequality for matrix martingales in the following theorem.
Theorem 2.1 (Rosenthal-Burkholder Inequality).
For any $p\geq 2$, suppose that
$\quad\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\Big\|_{p}<\infty\quad\text{ and }\quad\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|}\right\|_{p}<\infty.$
Then it holds that
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq C_{1}\sqrt{p\vee\log d}\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\Big\|_{p}+C_{2}(p\vee\log d)\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|}\right\|_{p},$ (8)
where $C_{1}=87$, $C_{2}=50$. When $p\geq 117$, the constants can be smaller: $C_{1}=64$, $C_{2}=28$.
Furthermore, the following extensions hold:
  1. (a)
    Rectangular Matrix: The above inequality also holds for the rectangular matrix martingale $({\bm{W}}_{k})_{k\in{\mathbb{N}}}$, in which we replace $\bm{M}$} by $\bm{W}$}, $\bm{X}$} by $\bm{D}$}, and $d$ by $d_{1}+d_{2}$.
  2. (b)
    Dimension-Free: In the self-adjoint case, if we further assume that there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that $\langle{\bm{M}}\rangle_{\infty}\preccurlyeq{\bm{\Upsilon}}$ almost surely, then the above inequality also holds when we replace $d$ by $e\cdot r({\bm{\Upsilon}}\wedge\|\|\left\langle{{\bm{M}}}\right\rangle_{\infty}\|\|_{\frac{p}{2}})$, and double the right-hand side.
    In the rectangular case (a), if we further assume that there exist ${\bm{\Upsilon}}_{1}\in{\mathbb{H}}_{+}^{d_{1}}$, ${\bm{\Upsilon}}_{2}\in{\mathbb{H}}_{+}^{d_{2}}$ such that $\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}{\bm{D}}_{k}^{*}]\preccurlyeq{\bm{\Upsilon}}_{1}$ and $\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}^{*}{\bm{D}}_{k}]\preccurlyeq{\bm{\Upsilon}}_{2}$ almost surely, then the above inequality also holds, while we replace $\bm{M}$} by $\bm{W}$}, $\bm{X}$} by $\bm{D}$}, $d$ by $2[r({\bm{\Upsilon}}_{1}\wedge\|\|\left\langle{{\bm{W}}}\right\rangle_{\infty}\|\|_{\frac{p}{2}})\vee r({\bm{\Upsilon}}_{2}\wedge\|\|\left\langle{{\bm{W}}}\right\rangle_{\infty}\|\|_{\frac{p}{2}})]$, and double the right-hand side.
By localization, our Rosenthal-Burkholder inequalities also hold if we consider matrix local martingales and replace $k\in{\mathbb{N}}$ by $0\leq k\leq\tau$ for any stopping time $\tau$; we omit the details for brevity. We also present another version of the Rosenthal-Burkholder inequality: $\|\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|\|_{p}\lesssim[(p/\log p)\vee\log d](\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\|_{p}+\|\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\|_{p})$, which can be found in Section 4.1.
For self-adjoint matrix martingales $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$, the proof strategy is to combine the extrapolation method (good $\lambda$ inequality) (Hitczenko, 1990, 1994; Pinelis, 1994) with techniques in matrix concentration inequalities (especially matrix Bennett inequalities) (Tropp, 2011, 2012, 2015; Kroshnin and Suvorikova, 2024). A detailed outline of the proof can be found in Section 4.1.
For the rectangular matrix martingale $({\bm{W}}_{k})_{k\in{\mathbb{N}}}$, we can directly apply the previous results to the self-adjoint matrix martingale $({\mathscr{H}}({\bm{W}}_{k}))_{k\in{\mathbb{N}}}$ to obtain the desired inequality, where $\mathscr{H}$} is the Hermitian dilation operator (Paulsen, 2002). For a rectangular matrix $\bm{B}$}, its Hermitian dilation ${\mathscr{H}}\left({{\bm{B}}}\right)$ is a self-adjoint matrix, which satisfies
${\mathscr{H}}\left({{\bm{B}}}\right):=\begin{bmatrix}\bm{0}&{\bm{B}}\\ {\bm{B}}^{*}&\bm{0}\end{bmatrix},\quad{\mathscr{H}}\left({{\bm{B}}}\right)^{2}=\begin{bmatrix}{\bm{B}}{\bm{B}}^{*}&\bm{0}\\ \bm{0}&{\bm{B}}^{*}{\bm{B}}\end{bmatrix},\quad\left\|{{\mathscr{H}}\left({{\bm{B}}}\right)}\right\|=\left\|{{\bm{B}}}\right\|.$
It is worth noting that the assumption in Theorem 2.1 (b), the quadratic variation $\langle{\bm{M}}\rangle_{\infty}$ has an almost sure upper bound $\bm{\Upsilon}$}, seems quite restrictive. However, such an assumption is exactly satisfied in deriving the dimension-free matrix Rosenthal inequality for uniformly geometrically ergodic Markov chains.
We also note that doubling the r.h.s. in Theorem 2.1 (b) is not essential; we could reduce this multiplier from $2$ to a constant close to $1$ by increasing the effective rank term that only appears as the logarithmic term. However, for brevity, we use this presentation and do not strive for the optimal constants.
A direct application of the Rosenthal-Burkholder inequality (Theorem 2.1) is to show the Burkholder-Davis-Gundy inequality for continuous matrix local martingales by discretization in the following theorem.
Theorem 2.2 (Burkholder-Davis-Gundy Inequality).
Let $({\bm{M}}_{t})_{t\geq 0}$ be a continuous local martingale taking values in ${\mathbb{H}}^{d}$. We denote by $(\left\langle{{\bm{M}}}\right\rangle_{t})_{t\geq 0}$ its quadratic variation process, that is, $\left\langle{{\bm{M}}}\right\rangle_{t}]_{ij}=\sum_{k=1}^{d}\left\langle{[{\bm{M}}]_{ik},[{\bm{M}}]_{kj}}\right\rangle_{t}$. For any $p\geq 2$ and stopping time $\tau$, suppose that $\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|^{1/2}\|_{p}<\infty$. Then it holds that
$\left\|{\sup_{0\leq t\leq\tau}\|{\bm{M}}_{t}\|}\right\|_{p}\leq C_{1}\sqrt{p\vee\log d}\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|^{1/2}\Big\|_{p},$
where $C_{1}$ is defined in Theorem 2.1.
The proof can be found in Appendix A.2. Maitre (2025, Theorem 1.1) also provided a version of Burkholder-Davis-Gundy inequality for continuous matrix martingales in the form: $\|\sup_{0\leq t\leq\tau}\|{\bm{M}}_{t}\|\|_{p}\lesssim\sqrt{p}\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|\|_{p}^{1/2}+\sqrt{\log d}\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|^{1/2}\|_{p}$, our version (Theorem 2.2) is sharper because $\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|^{1/2}\|_{p}\leq\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|\|_{p}^{1/2}$ by Jensen’s inequality.
Discussion.
In the independent case, Chen et al. (2012, Theorem A.1) provided a matrix Rosenthal inequality, and Jirak et al. (2025, Theorem 3) further gave a dimension-free version. We extend their results from the independent case to the martingale case. We remark that, in the independent case, it is possible to translate the universality-based concentration inequalities (Brailovskaya and van Handel, 2024, Theorem 2.8 and Theorem 2.16) into Rosenthal-type inequalities by integrating the tail bounds. In some cases, this yields tighter upper bounds than the classic results Chen et al. (2012, Theorem A.1). This observation gives rise to an interesting question: in the martingale case, does certain form of universality-based moment inequality hold? We leave this as a direction for future research.
In the field of noncommutative probability, Junge and Xu (2003, 2008); Junge and Zeng (2013, 2015); Jiao et al. (2025) investigated probability inequalities for noncommutative martingales. In our random matrix setting, their results can be translated into probability inequalities concerning the Schatten $p$-norm, rather than the spectral norm that is the focus of this paper. In particular, in the independent case, Junge and Zeng (2013, Theorem 0.4) gives the following Rosenthal inequality
${\mathbb{E}}^{{1}/{p}}\left[{\operatorname{Tr}\left({\left|{{\bm{M}}_{\infty}}\right|^{p}}\right)}\right]\lesssim\sqrt{p}{\mathbb{E}}^{{1}/{p}}\left[{\operatorname{Tr}\left({\left\langle{{\bm{M}}}\right\rangle_{\infty}^{p/2}}\right)}\right]+p{\mathbb{E}}^{{1}/{p}}\left[{\sup_{k\in{\mathbb{N}}}\operatorname{Tr}\left({\left|{{\bm{X}}_{k}}\right|^{p}}\right)}\right],$
for $p\geq 2.5$. However, in the martingale case, whether a Rosenthal-Burkholder inequality of the following form holds remains an open problem (Junge and Zeng, 2015, Problem 2.5)
${\mathbb{E}}^{{1}/{p}}\left[{\operatorname{Tr}\left({\left|{{\bm{M}}_{\infty}}\right|^{p}}\right)}\right]\lesssim\sqrt{p}{\mathbb{E}}^{{1}/{p}}\left[{\operatorname{Tr}\left({\left\langle{{\bm{M}}}\right\rangle_{\infty}^{p/2}}\right)}\right]+f(p){\mathbb{E}}^{{1}/{p}}\left[{\sum_{k\in{\mathbb{N}}}\operatorname{Tr}\left({\left|{{\bm{X}}_{k}}\right|^{p}}\right)}\right],$
where the coefficient of the dominant term (quadratic variation) has an optimal $\sqrt{p}$ dependence on $p$, and $f(p)$ is some function for which they conjectured $f(p)=p$ as in the commutative case.
2.2 Matrix Rosenthal and Concentration Inequalities for Markov Chains
Consider the Markov kernel ${\mathrm{Q}}\colon{\mathcal{Z}}\times{\mathscr{Z}}\to[0,1]$ on the measurable space $({\mathcal{Z}},{\mathscr{Z}})$ with a unique stationary distribution $\mu\in{\mathcal{P}}$, where $\mathcal{P}$} is the set of all probability distributions on $({\mathcal{Z}},{\mathscr{Z}})$. For any $\xi\in{\mathcal{P}}$, we denote by ${\mathbb{P}}_{\xi}$ (resp. ${\mathbb{E}}_{\xi}$) the probability (resp. expectation) of the canonical process $(Z_{k})_{k\in{\mathbb{N}}}$ with initial distribution $\xi$, and $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$ the nature filtration where ${\mathscr{F}}_{k}=\sigma(Z_{0},\ldots,Z_{k})$. That is, $(Z_{k})_{k\in{\mathbb{N}}}$ is a stochastic process in $({\mathcal{Z}}^{\mathbb{N}},{\mathscr{Z}}^{\mathbb{N}},{\mathbb{P}}_{\xi})$ adapted to $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$, which satisfies ${\mathbb{P}}_{\xi}(Z_{0}\in A)=\xi(A)$, ${\mathbb{P}}_{\xi}(Z_{k+1}\in A\mid{\mathscr{F}}_{k})={\mathbb{P}}_{\xi}(Z_{k+1}\in A\mid Z_{k})={\mathrm{Q}}(Z_{k},A)$ ${\mathbb{P}}_{\xi}$-almost surely for any $A\in{\mathscr{Z}}$ and $k\in{\mathbb{N}}$. Thus, $(Z_{k})_{k\in{\mathbb{N}}}$ is an Markov chain.
For any $p\geq 1$, $\xi\in{\mathcal{P}}$ and random variable $X$ defined on $({\mathcal{Z}}^{\mathbb{N}},{\mathscr{Z}}^{\mathbb{N}},{\mathbb{P}}_{\xi})$, we denote by $\|X\|_{\xi,p}=({\mathbb{E}}_{\xi}[|X|^{p}])^{1/p}$ the $L^{p}$ norm of $X$ under the initial distribution $\xi$.
For any measurable function $f$ in $({\mathcal{Z}},{\mathscr{Z}})$ and distribution $\xi\in{\mathcal{P}}$, we denote $\left\|{f}\right\|_{\infty}=\sup_{z\in{\mathcal{Z}}}\left|{f(z)}\right|$ ($f$ is real-valued), $\xi(f)=\int_{\mathcal{Z}}f(z)\xi(dz)$, $(\xi{\mathrm{Q}})(\cdot)=\int_{{\mathcal{Z}}}{\mathrm{Q}}(z,\cdot)\xi(dz)$, and $({\mathrm{Q}}f)(\cdot)=\int_{{\mathcal{Z}}}f(z){\mathrm{Q}}(\cdot,dz)$.
Now, we introduce the ergodicity assumptions we consider in this paper, we will only consider the uniformly geometrically ergodic Markov chains in this section, and the more general geometrically $V$-ergodic Markov chains in Section 2.2.1. We recall that an irreducible and aperiodic Markov chain with a finite state space is always uniformly geometrically ergodic.
Assumption 2.1 (Uniformly Geometrically Ergodic Markov Chains).
There exists a $t_{\operatorname{mix}}\in{\mathbb{N}}$ such that for any $k\in{\mathbb{N}}$,
$\sup_{z\in{\mathcal{Z}}}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\mu}\right\|_{\textup{TV}}\leq 2\left({1/4}\right)^{\lfloor k/t_{\operatorname{mix}}\rfloor},$
where $\left\|{\xi}\right\|_{\textup{TV}}:=\left|{\xi}\right|({\mathcal{Z}})$ is the total variation norm of the signed measure $\xi$ on $({\mathcal{Z}},{\mathscr{Z}})$.
We consider a measurable self-adjoint matrix function ${\bm{F}}\colon{\mathcal{Z}}\to{\mathbb{H}}^{d}$. We assume $\mu({\bm{F}})=\int_{{\mathcal{Z}}}{\bm{F}}(z)\mu(dz)=\bm{0}_{d{\times}d}$ and $\|\|{\bm{F}}\|\|_{\infty}=\sup_{z\in{\mathcal{Z}}}\|{\bm{F}}(z)\|<\infty$. Our goal is to give counterparts to matrix Rosenthal, Hoeffding, Bernstein inequalities for the spectral norm of the summation ${\bm{S}}_{n}=\sum_{k=0}^{n-1}{{\bm{F}}}(Z_{k})$. We observe that the boundedness condition is indispensable. As argued by Fan et al. (2021, Theorem 4), even for scalar unbounded functions, it is impossible to obtain Hoeffding-type inequalities for ${\bm{S}}_{n}$ under Assumption 2.1 or a non-zero absolute $L^{2}$-spectral gap, whereas our Rosenthal inequality implies the Hoeffding-type inequality for ${\bm{S}}_{n}$.
Recall that the Markov chain CLT (Jones, 2004, Corollary 5) says that $n^{-1/2}{\bm{S}}_{n}$ converges weakly to a Gaussian distribution and satisfies $\lim_{n\to\infty}n^{-1}{\mathbb{E}}_{\xi}[{\bm{S}}_{n}^{2}]={\bm{\Sigma}}_{\mu}({\bm{F}})$ for any $\xi\in{\mathcal{P}}$, where ${\bm{\Sigma}}_{\mu}({\bm{F}})$ is the long-run variance defined as
${\bm{\Sigma}}_{\mu}({\bm{F}}):=\mu\left({{{\bm{F}}}^{2}}\right)+\sum_{k=1}^{\infty}\mu\left({{{\bm{F}}}{\mathrm{Q}}^{k}{{\bm{F}}}+\left({{\mathrm{Q}}^{k}{{\bm{F}}}}\right){{\bm{F}}}}\right).$
We aim to obtain a Rosenthal inequality whose leading term is given by the long-run variance ${\bm{\Sigma}}_{\mu}({\bm{F}})$, matching the Markov chain CLT.
Dimension-Free: When we address dimension-free results, we need to make an additional assumption, that is, there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$. Note that the effective rank $r({\bm{\Upsilon}})$ could be much smaller than the ambient dimension $d$.
Rectangular Case: In this case, we consider a rectangular matrix function ${\bm{R}}\colon{\mathcal{Z}}\to{\mathbb{C}}^{d_{1}{\times}d_{2}}$ that satisfies $\mu({\bm{R}})=\int_{{\mathcal{Z}}}{\bm{R}}(z)\mu(dz)=\bm{0}_{d_{1}{\times}d_{2}}$. We denote by ${\bm{V}}_{n}=\sum_{k=0}^{n-1}{\bm{R}}(Z_{k})$ the summation and
[…]
called the proxy of the long-run variance.
Now we state one of our main results, the matrix Rosenthal inequality for uniformly geometrically ergodic Markov chains.
Theorem 2.3 (Matrix Rosenthal Inequality for Uniformly Geometrically Ergodic Markov Chains).
Under Assumption 2.1 and $\|\|{\bm{F}}\|\|_{\infty}<\infty$, we have that for any $p\geq 2$, $n\geq(p\vee\log d)t_{\operatorname{mix}}$, and $\xi\in{\mathcal{P}}$,
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $C_{1}\sqrt{\frac{(p\vee\log d)\left\|{{\bm{\Sigma}}_{\mu}({\bm{F}})}\right\|}{n}}+D_{1}t_{\operatorname{mix}}^{1/2}\left({\frac{(p\vee\log d)t_{\operatorname{mix}}}{n}}\right)^{3/4}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty},$
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}+D_{2}\frac{pt_{\operatorname{mix}}}{n}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty},$ (10)
where $C_{1}$ is defined in Theorem 2.1 and $D_{1},D_{2}>0$ are universal constants, which can be traced in the proof.
Furthermore, the following extensions hold:
  1. (a)
    Rectangular Matrix: the above inequality also holds for the rectangular case if we further assume $\|\|{\bm{R}}\|\|_{\infty}<\infty$ and replace $\bm{S}$} by $\bm{V}$}, $\bm{F}$} by $\bm{R}$}, and $d$ by $d_{1}+d_{2}$.
  2. (b)
    Dimension-Free: in the self-adjoint case, if we further assume that there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$, then the above inequality also holds when we replace the $d$ in the first term by $d_{n,p}=e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\mu,{p}/{2}}])$, and replace the $d$ in the second term by the dimension factor defined after Eqn. (19), and double the right-hand side.
    In the rectangular case (a), under similar assumptions, analogous dimension-free results hold, and we omit the details.
The proof strategy of Theorem 2.3 is to leverage the Poisson equation ${\bm{G}}-\left({{\mathrm{Q}}{\bm{G}}}\right)={\bm{F}}$ to construct a matrix martingale ${\bm{M}}_{k}=\sum_{j=1}^{k}{\bm{G}}(Z_{j})-({\mathrm{Q}}{\bm{G}})(Z_{j-1})$, and then we apply our Rosenthal-Burkholder inequality (Theorem 2.1). This method was introduced by Durmus et al. (2023), who derived a Hilbert space-valued Rosenthal inequality for Markov chains. A detailed outline of the proof can be found in Section 4.2.
We note that the dependence of the second term in Theorem 2.3 on $t_{\operatorname{mix}}$ is suboptimal. If we want the first term to dominate the second term, we need to take $n\gtrsim(p\vee\log d)t_{\operatorname{mix}}^{3}$, which is not linear in $t_{\operatorname{mix}}$. In Durmus et al. (2023), they used dyadic descent argument to derive a Hilbert space-valued Rosenthal inequality with residual terms that have a sharp dependence on $t_{\operatorname{mix}}$. However, this argument does not work for the matrix case, because under the Löwner order, the square function is not operator monotone, and the $p$-th power function ($p>2$) is not operator convex, making the core inequality (23) in the dyadic descent argument of Durmus et al. (2023) invalid in our matrix case.
Fortunately, if we do not insist on obtaining the result with a leading term matching the Markov chain CLT, we can still derive results with the sharp dependence on $t_{\operatorname{mix}}$ by using the blocking argument. In the following theorem, we present another version of the Rosenthal inequality, whose leading term is given by the variance proxy $t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{\infty}^{2}$, which is an upper bound for the long-run variance $\|{\bm{\Sigma}}_{\mu}({\bm{F}})\|$.
Theorem 2.4.
Under Assumption 2.1 and $\|\|{\bm{F}}\|\|_{\infty}<\infty$, we have that for all $p\geq 2$ and $n\geq 2$,
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $\frac{16e}{3}C_{1}\sqrt{\frac{(p\vee\log d)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}^{2}}{n}}+\left({\frac{16e}{3}C_{2}(p\vee\log d)+\frac{19}{3}}\right)\frac{t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}}{n}.$
For brevity, we omit the inequality for the rectangular case, and defer the dimension-free results as well as the proofs to Lemma 4.3 (a).
By our Rosenthal inequalities and applying Markov inequality to $\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|^{\log(1/\delta)}$, we immediately obtain the following Hoeffding and Bernstein inequalities.
Corollary 2.1 (Matrix Hoeffding Inequalities for Uniformly Geometrically Ergodic Markov Chains).
Under Assumption 2.1 and $\|\|{\bm{F}}\|\|_{\infty}<\infty$, for any $\delta\in(0,1)$ and $n\geq 2$, with ${\mathbb{P}}_{\xi}$-probability at least $1-\delta$, it holds that
$\frac{1}{n}\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|\leq$ $\frac{16e}{3}C_{1}\sqrt{\frac{\log(d/\delta)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}^{2}}{n}}+D_{3}\frac{\log(d/\delta)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}}{n},$
where $D_{3}>0$ is a universal constant, which can be taken as $6eC_{2}+eD_{2}$.
Corollary 2.2 (Matrix Bernstein Inequalities for Uniformly Geometrically Ergodic Markov Chains).
Under Assumption 2.1 and $\|\|{\bm{F}}\|\|_{\infty}<\infty$, for any $\delta\in(0,1)$ and $n\geq 2$, with ${\mathbb{P}}_{\xi}$-probability at least $1-\delta$, it holds that
$\frac{1}{n}\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|\leq$ $eC_{1}\sqrt{\frac{\log(d/\delta)\left\|{{\bm{\Sigma}}_{\mu}({\bm{F}})}\right\|}{n}}+e(D_{1}+D_{2})t_{\operatorname{mix}}^{1/2}\left({\frac{\log(d/\delta)t_{\operatorname{mix}}}{n}}\right)^{3/4}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
For the above concentration inequalities, the inequalities for both the rectangular case and the dimension-free version (where one only needs to replace $p$ by $\log(1/\delta)$ in the dimension factor) still hold. We omit them here.
2.2.1 Results for Geometrically $V$-Ergodic Markov Chains
The assumptions of uniform geometric ergodicity (Assumption 2.1) and bounded functions are excessively strong and fail to hold in many applications. We now introduce the more general geometrically ergodic Markov chains with Lyapunov function $V$ and present matrix Rosenthal inequalities under the ergodicity assumption. Let $V\colon{\mathcal{Z}}\to[e,\infty)$ be a Lyapunov function satisfying $\mu(V)<\infty$. We denote by ${\mathcal{P}}_{V}$ the set of all probability distributions whose elements $\xi$ satisfy $\xi(V)<\infty$. For any real-valued measurable function $f$, we also define its $V$-norm as $\|f\|_{V}=\sup_{z\in{\mathcal{Z}}}{\left|{f(z)}\right|}/{V(z)}$. We will see that, under the ergodicity assumption, for matrix functions $\bm{F}$} whose growth rate does not exceed $V^{1/p}$, the convergence of $n^{-1}{\bm{S}}_{n}$ in the $L^{p}$ norm can be obtained, which allows us to extend the result to unbounded functions. In this section we omit the results in the rectangular case for brevity.
Assumption 2.2 (Geometrically $V$-Ergodic Markov Chains).
There exist $t_{\operatorname{mix}}\in{\mathbb{N}}$ and $\varkappa\geq 1$ such that for any $k\in{\mathbb{N}}$,
$\sup_{z,z^{\prime}\in{\mathcal{Z}}}\frac{\left\|{{\mathrm{Q}}^{k}(z,\cdot)-{\mathrm{Q}}^{k}(z^{\prime},\cdot)}\right\|_{V}}{V(z)+V(z^{\prime})}\leq\varkappa\left({\frac{1}{4}}\right)^{\lfloor k/t_{\operatorname{mix}}\rfloor},$
where $\left\|{\xi}\right\|_{V}:=\left|{\xi}\right|(V)$ is the the $V$-total variation norm of the signed measure $\xi$ on $({\mathcal{Z}},{\mathscr{Z}})$.
This assumption can be verified via the Foster-Lyapunov drift condition (Meyn and Tweedie, 2012). Assumption 2.2 implies that for any $\xi\in{\mathcal{P}}_{V}$ and $k\in{\mathbb{N}}$,
$\left\|{\xi{\mathrm{Q}}^{k}-\mu}\right\|_{V}\leq\varkappa\left({\xi(V)+\mu(V)}\right)\left({1/4}\right)^{\lfloor k/t_{\operatorname{mix}}\rfloor}.$
Hence, Assumption 2.1 is a special case of Assumption 2.2 with $V\equiv e$ and $\varkappa=1$.
Now we are ready to present matrix Rosenthal inequalities for geometrically $V$-ergodic Markov chains. We first present the result parallel to Theorem 2.3 whose leading term matches the Markov chain CLT, with the proof in Section 4.2.
Theorem 2.5 (Matrix Rosenthal Inequality for Geometrically $V$-Ergodic Markov Chains).
Under Assumption 2.2 and $\|\|{\bm{F}}\|\|_{V^{1/p}}<\infty$ for $p\geq 2$, we have that for any $n\geq 2$, and $\xi\in{\mathcal{P}}_{V}$,
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $C_{1}\sqrt{\frac{(p\vee\log d)\left\|{{\bm{\Sigma}}_{\mu}({\bm{F}})}\right\|}{n}}+D_{4}p^{3/2}\left({p\vee\log d}\right)^{3/4}t_{\operatorname{mix}}^{5/4}n^{-3/4}\left[{\varkappa\mu(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
$+D_{5}p^{3/2}\left({p\vee\log d}\right)t_{\operatorname{mix}}^{3/2}n^{-(1-1/p)}\left[{\varkappa\mu(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}},$
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}+\frac{32}{15n}\varkappa^{1/p}\left({\xi(V)+\mu(V)}\right)^{1/p}pt_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}},$
where $C_{1}$ is defined in Theorem 2.1 and $D_{4},D_{5}>0$ are universal constants, which can be traced in the proof.
The proof strategy is identical to that under uniform geometric ergodicity assumption. And the result also has a suboptimal dependence on $t_{\operatorname{mix}}$.
Next, we present a parallel result to Theorem 2.4, whose leading term is given by the variance proxy $t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{1/p}}^{2}$, and the proof can be found in Lemma 4.3 (b).
Theorem 2.6.
Under Assumption 2.2 and $\|\|{\bm{F}}\|\|_{V^{1/p}}<\infty$ for $p\geq 2$, we have that,
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $\frac{8}{3}C_{1}p\sqrt{\frac{(p\vee\log d)t_{\operatorname{mix}}}{n}}\left[{\varkappa\mu(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
$+\left({\frac{16}{3}C_{2}(p\vee\log d)+\frac{11}{3}}\right)\frac{pt_{\operatorname{mix}}}{n^{1-1/p}}\left[{\varkappa\mu(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$
We can also use Markov inequality to obtain high-probability upper bounds similar to Corollary 2.1 and Corollary 2.2. For brevity, we omit them here.
Discussion.
As discussed in Related Work, Garg et al. (2018); Qiu et al. (2020); Neeman et al. (2024); Wu et al. (2025) derived matrix concentration inequalities for Markov chains with a non-zero absolute $L^{2}$-spectral gap; and Van Werde and Sanders (2025) derived universality-based concentration inequalities for $\psi$-mixing Markov chains. Instead of the multi-matrix Golden-Thompson inequality (Sutter et al., 2017; Garg et al., 2018), the León-Perron argument (León and Perron, 2004), or the universality-based approach (Bandeira et al., 2023; Brailovskaya and van Handel, 2024) used in these works, we leverage the Poisson equation approach (Durmus et al., 2023), which allows us to overcome several limitations of the previous results: First, our results can be applied to general non-reversible geometrically ergodic Markov chains without assuming a non-zero absolute spectral gap or $\psi$-mixing condition. Second, we do not need to assume the boundedness of functions under the geometric ergodicity assumption. Third, the leading term of our results matches the Markov chain CLT, rather than relying on a sub-optimal variance proxy as in (Garg et al., 2018; Qiu et al., 2020; Neeman et al., 2024; Wu et al., 2025). Finally, under further assumptions, we obtain dimension-free results, while the approaches employed in previous works failed to achieve them. Applying the dimension-free results, we can also obtain the concentration inequalities for linear operators in infinite-dimensional Hilbert spaces.
Nevertheless, our results have the following limitations: As mentioned before, the residual term in the results where the leading term matches the Markov chain CLT (Theorem 2.3, Theorem 2.2, Theorem 2.5) has a suboptimal dependence on $t_{\operatorname{mix}}$, failing to achieve the sharp dependence on $t_{\operatorname{mix}}$ in the residual term. This issue arises because operator convexity (monotonicity) for a scalar-valued function acted on matrices imposes stronger requirements than convexity (monotonicity), and consequently, arguments from the scalar case (Durmus et al., 2023) fail to hold. We leave the improvement to the sharp dependence on $t_{\operatorname{mix}}$ for future work. Another limitation is that the Poisson equation approach is only applicable to time-homogeneous functions (additive functionals of Markov chains), whereas the results in Neeman et al. (2024) apply to time-inhomogeneous functions. This makes our results inapplicable to problems such as stochastic approximation with Markovian data (Li et al., 2023; Mou et al., 2024; Huo et al., 2024; Samsonov et al., 2025). We also leave the extension to time-inhomogeneous functions as future work. A final limitation is that the universality-based inequalities in Van Werde and Sanders (2025) yield sharper bounds, although they require the stronger $\psi$-mixing condition and a finite dimension $d$. A key direction for future research is to determine if universality-based sharp concentration inequalities can be derived under weaker ergodicity assumptions. One potential approach is to first establish a universality-based moment inequality for matrix martingales and then apply the Poisson equation method. We leave this as future work as well.
3 Applications
In this section, we apply our results to covariance estimation and principal component analysis (PCA) problems on Markovian data, to showcase their effectiveness.
We use the same notation as in Section 2.2. Suppose $(Z_{k})_{k\in{\mathbb{N}}}$ is a uniformly geometrically ergodic stationary Markov chain (Assumption 2.1) and $\bm{f}\colon{\mathcal{Z}}\to{\mathbb{R}}^{d}$ is a bounded vector-valued function. We first consider the problem of covariance estimation on Markovian data, which is also considered in Fan et al. (2021). That is, we are interested in analyzing the spectral norm error in estimating the covariance matrix ${\bm{\Sigma}}=\mu(\bm{f}\bm{f}^{\top})={\mathbb{E}}_{\mu}[\bm{f}(Z_{0})\bm{f}(Z_{0})^{\top}]$ using the sample covariance matrix $\hat{{\bm{\Sigma}}}_{n}=n^{-1}\sum_{k=0}^{n-1}\bm{f}(Z_{k})\bm{f}(Z_{k})^{\top}$. Let ${\bm{F}}(z):=\bm{f}(z)\bm{f}(z)^{\top}-{\bm{\Sigma}}$. Then $\bm{F}$} satisfies $\mu({\bm{F}})=\bm{0}_{d{\times}d}$ and $\|\|{\bm{F}}\|\|_{\infty}\leq\|\|\bm{f}\|\|_{\infty}^{2}$. Using the matrix Bernstein inequality for Markov chains (Corollary 2.2), we obtain that for any $\delta\in(0,1)$, with probability at least $1-\delta$,
[…]
where ${\bm{\Sigma}}_{\mu}({\bm{F}})$ is the long-run variance of $\bm{F}$} defined in Eqn. (3).
Next, we consider the problem of principal component analysis on Markovian data, which is also considered in Neeman et al. (2024); Kumar and Sarkar (2023). We denote by $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{d}\geq 0$ the eigenvalues of $\bm{\Sigma}$}, and ${\bm{v}}_{1},{\bm{v}}_{2},\ldots,{\bm{v}}_{n}$ the corresponding eigenvectors. We assume $\lambda_{1}-\lambda_{2}>0$. In the PCA problem, we are interested in analyzing the $\sin^{2}$ error in estimating the leading eigenvector ${\bm{v}}_{1}$ of $\bm{\Sigma}$} using the leading eigenvector $\hat{{\bm{v}}}_{1,n}$ of $\hat{{\bm{\Sigma}}}_{n}$. By Eqn. (11) and Wedin’s Theorem (Wedin, 1972), we immediately have
$1-\left({{\bm{v}}_{1}^{\top}\hat{{\bm{v}}}_{1,n}}\right)^{2}\leq C\frac{\log(d/\delta)\left\|{{\bm{\Sigma}}_{\mu}({\bm{F}})}\right\|}{(\lambda_{1}-\lambda_{2})^{2}n}+O\left({\frac{1}{n^{3/2}}}\right),$
where $C>0$ is a universal constant.
Under further assumptions on $\bm{f}$, we can obtain dimension-free results. Let ${\bm{\Upsilon}}\in{\mathbb{H}}_{d}^{+}$ be a real PSD matrix with spectral decomposition ${\bm{\Upsilon}}=\sum_{i=1}^{d}\alpha_{i}{\bm{u}}_{i}{\bm{u}}_{i}^{\top}$ such that $\alpha_{1}\geq\alpha_{2}\geq\ldots\geq\alpha_{d}\geq 0$, and $\bm{f}(z)=\sum_{i=1}^{d}\beta_{i}(z){\bm{u}}_{i}$—the representation of $\bm{f}$ on the orthonormal basis $\{{\bm{u}}_{1},\ldots,{\bm{u}}_{d}\}$. Suppose that $\sup_{z\in{\mathcal{Z}}}\sum_{i=1}^{d}\beta_{i}^{2}(z)/\alpha_{i}\leq 1$ ($0/0:=0$ and $1/0:=\infty$). Then by the Schur complement lemma (Zhang, 2006), one has that $\bm{f}(z)\bm{f}(z)^{\top}\preccurlyeq{\bm{\Upsilon}}$ for any $z\in{\mathcal{Z}}$. Furthermore, ${\bm{F}}^{2}(z)=(\bm{f}(z)\bm{f}(z)^{\top}-{\bm{\Sigma}})^{2}\preccurlyeq 2[\bm{f}(z)\bm{f}(z)^{\top}]^{2}+2{\bm{\Sigma}}^{2}\preccurlyeq 2\|\|\bm{f}\|\|_{\infty}^{2}[\bm{f}(z)\bm{f}(z)^{\top}+{\bm{\Sigma}}]\preccurlyeq 4\|\|\bm{f}\|\|_{\infty}^{2}{\bm{\Upsilon}}$. Using the dimension-free results, we can replace the ambient $d$ in the bounds by the effective rank $e\cdot r([(64/9)t_{\operatorname{mix}}^{2}\|\|\bm{f}\|\|_{\infty}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\mu,(1/2)\log(1/\delta)}])$. Similar results hold for the more general geometrically $V$-ergodic Markov chains and unbounded function $\bm{f}$, and we omit them for brevity.
Compared with the results in Kumar and Sarkar (2023); Neeman et al. (2024), our results obtain a leading term that matches the Markov chain CLT and gain dimension-free results. Moreover, our results can be extended to the more general geometrically $V$-ergodicity assumption and unbounded functions. These aspects demonstrate the advantages of our results over the previous ones as we discussed earlier.
4 Proofs
In this section, we outline the proofs of our main results.
4.1 Proof of the Rosenthal-Burkholder Inequality for Matrix Martingales
Based on the Bennett inequality for matrix martingales (Lemma A.1), we can obtain the following good $\lambda$ inequality for conditionally symmetric matrix martingales, which is a matrix version of Pinelis (1994, Lemma 4.2). Before that, we need to introduce the concept of conditionally symmetric martingales.
Definition 4.1 (Conditionally Symmetric Martingales).
The martingale $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ is said to be conditionally symmetric if the conditional distributions of ${\bm{X}}_{k}$ and ${\bm{X}}_{k}$ given ${\mathscr{F}}_{k-1}$ coincide almost surely for any $k\in{\mathbb{N}}$.
With the assumption that the martingale $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ is conditionally symmetric, the analysis can be greatly simplified, because the truncated martingale remains a martingale. Later on, we will introduce how to perform symmetrization to extend the conclusion to general martingales.
Theorem 4.1 (Good $\lambda$ Inequality).
Suppose that $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ is conditionally symmetric. For any $\lambda,\delta_{1},\delta_{2}>0$ and $\beta>1+\delta_{2}$, letting $N=[\beta-(1+\delta_{2})]/{\delta_{2}}$, we have that
${\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|>\beta\lambda,\ \left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\leq\delta_{1}\lambda,\ \sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda}\right)\leq 2d\left({\frac{e}{N}\frac{\delta_{1}^{2}}{\delta_{2}^{2}}}\right)^{N}{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|>\lambda}\right).$
Furthermore, if there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that $\langle{\bm{M}}\rangle_{\infty}\preccurlyeq{\bm{\Upsilon}}$ almost surely, then the above inequality also holds in which $d$ is replaced by $e\cdot r({\bm{\Upsilon}}\wedge(\delta_{1}^{2}\lambda^{2}))$.
The proof is similar to Pinelis (1994, Lemma 4.2) and can be found in Appendix A.1. The key ingredient is the Bennett inequality for matrix martingales (Lemma A.1). Roughly speaking, this good $\lambda$ inequality indicates that if the supremum of a conditionally symmetric martingale is sufficiently large, then with high probability, both the quadratic variation and the supremum of the martingale difference will also be large.
We observe that the dimension-free factor $r\left({{\bm{\Upsilon}}\wedge(\delta_{1}^{2}\lambda^{2})}\right)$ in Theorem 4.1 increases to $d$ as $\lambda$ decreases to 0. Therefore, to obtain a valid dimension-free result, it is necessary to truncate the error level $\lambda$ before applying the good $\lambda$ inequality.
The following lemma, the truncated $\Phi$ inequality, translates a truncated good $\lambda$ inequality into a moment inequality, which is an extension of Burkholder (1973, Lemma 7.1).
Lemma 4.1 (Truncated $\Phi$ Inequality).
Suppose that $\Phi\colon{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$ is non-decreasing continuous satisfying $\Phi(\beta\lambda)\leq\gamma\Phi(\lambda)$ for all $\lambda\geq 0$, where $\beta>1$ and $\gamma>0$ are some constants; and $Y,Z$ are nonnegative random variables satisfying the following good $\lambda$ inequality
${\mathbb{P}}\left({Y>\beta\lambda,Z\leq\lambda}\right)\leq\epsilon{\mathbb{P}}\left({Y>\lambda}\right),\quad\forall\lambda\geq\frac{a}{\beta}$
for some constants $a\geq 0$ and $\epsilon\in(0,1/\gamma)$. Then we have that
${\mathbb{E}}\left[{\Phi(Y)}\right]\leq\frac{1}{1-\gamma\epsilon}\Phi(a)+\frac{\gamma}{1-\gamma\epsilon}{\mathbb{E}}\left[{\Phi(Z)}\right].$
The proof can be found in Appendix A.3.
Now, we are ready to give a full proof of the Rosenthal-Burkholder inequality (Theorem 2.1). As in the proof of the scalar Rosenthal-Burkholder inequality (Hitczenko, 1990; Pinelis, 1994), we first deal with conditionally symmetric martingales, and then generalize the results to general martingales.
Step 1: Conditionally Symmetric Martingales.
Assume that $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ is conditionally symmetric. We start with the dimension-dependent case. Applying the good $\lambda$ inequality (Theorem 4.1) and the truncated $\Phi$ inequality (Lemma 4.1) with $Y=\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|$, $Z=\delta_{1}^{-1}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}+\delta_{2}^{-1}\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|$, $\Phi(x)=x^{p}$, $\beta>1$ to be determined, $\gamma=\beta^{p}$, $a=0$, and
$\epsilon=2d\left({\frac{e}{N}\frac{\delta_{1}^{2}}{\delta_{2}^{2}}}\right)^{N},$
we have the following inequality
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq\left({\frac{1}{1-\gamma\epsilon}}\right)^{1/p}\beta\left({\delta_{1}^{-1}\left\|{\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}}\right\|_{p}+\delta_{2}^{-1}\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|}\right\|_{p}}\right),$
provided that
[…]
We can choose the following parameters such that $\beta>1+\delta_{2}$ and the inequality (12) holds:
$\beta=1+e^{-p/c}+\frac{1}{c\vee\log d},\quad\delta_{1}=\frac{1}{5\sqrt{c\vee\log d}\ e^{p/c}},\quad\delta_{2}=\frac{1}{5(c\vee\log d)},$
where $c\in[1,p]$ is a parameter to be determined. Now, we check that inequality (12) indeed holds. First note that $e^{-p/c}\geq e^{-1}$ and then
$N=\frac{\beta-1-\delta_{2}}{\delta_{2}}=4+5(c\vee\log d)e^{-p/c}\geq 4+5e^{-1}\log d,$
$\beta^{p}=\left[{1+(N+1)\delta_{2}}\right]^{p}\leq e^{2pN\delta_{2}}=e^{2pN/[5(c\vee\log d)]}\leq e^{2pN/(5c)},$
$\frac{e}{N}\frac{\delta_{1}^{2}}{\delta_{2}^{2}}=\frac{e}{4e^{2p/c}(c\vee\log d)^{-1}+5e^{p/c}}\leq\frac{e}{5}e^{-p/c},$
hence
$\gamma\epsilon=2d\left({\frac{e}{N}\frac{\delta_{1}^{2}}{\delta_{2}^{2}}}\right)^{N}\beta^{p}$ $\leq 2d\left({\frac{e}{5}}\right)^{4+5e^{-1}\log d}e^{-21p/(5c)}<2(e/5)^{4}e^{-21/5}<1,$
which is the desired conclusion. Here we used the fact $(e/5)^{5/e}<e^{-1}$ and $2(e/5)^{4}<1/2$. Then we have
$\left({\frac{1}{1-\gamma\epsilon}}\right)^{1/p}<\frac{1}{\sqrt{1-2(e/5)^{4}e^{-21/5}}}<1.002.$
Up to this point, we have shown that for the conditionally symmetric martingale $\{{\bm{M}}_{k}\}_{k\in{\mathbb{N}}}$, the following inequality holds:
[…]
Setting $c=p$, we obtain the inequality in the form given in Theorem 2.1. Alternatively, we can choose $c$ as the unique solution to equation $\sqrt{c}=e^{p/c}$, which yields another common form of the Rosenthal-Burkholder inequality $\|\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|\|_{p}\lesssim[(p/\log p)\vee\log d](\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\|_{p}+\|\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\|_{p})$. We shall not delve into this form in further detail in this paper.
Dimension-Free: To extend the result to a dimension-free version, we only need to replace $a=0$ with
$\tilde{a}=\beta\left({\delta_{1}^{-1}\left\|{\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}}\right\|_{p}+\delta_{2}^{-1}\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|}\right\|_{p}}\right),$
and $d$ with
$\tilde{d}=e\cdot r\left({{\bm{\Upsilon}}\wedge\left[{\left\|{\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|}\right\|_{\frac{p}{2}}}\right]}\right).$
And we need to double the right-hand side of the inequality (13) due to the additional appearance of $\Phi(a)$, that is, if $\langle{\bm{M}}\rangle_{\infty}\preccurlyeq{\bm{\Upsilon}}$ almost surely, then it holds that
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq 10.02\!\left({1{+}e^{-p/c}{+}\frac{1}{c{\vee}\log\tilde{d}}}\right)\left({\sqrt{c{\vee}\log\tilde{d}}\ e^{p{/}c}\left\|{\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1{/}2}}\right\|_{p}{+}(c{\vee}\log\tilde{d})\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|}\right\|_{p}}\right).$
Step 2: Generalize to Arbitrary Martingales
In order to generalize the result under the conditionally symmetry assumption to arbitrary martingales, in the special case of independence increment, it suffices to introduce i.i.d. Rademacher random variables; while in the case of martingales, a more sophisticated technique, named decoupling tangent sequence (Kwapien and Woyczynski, 1989; Hitczenko, 1988; De la Pena and Giné, 2012), is required. The proof in this part is similar to those of the scalar Rosenthal-Burkholder inequality (Hitczenko, 1990), and we defer the presentation to Appendix A.4.
4.2 Proof of the Matrix Rosenthal Inequality for Markov Chains
We consider the Poisson equation with the matrix forcing function $\bm{F}$}:
[…]
In the following lemma, we show that ${\bm{G}}=\sum_{k=0}^{\infty}{\mathrm{Q}}^{k}{\bm{F}}$ is a solution to the Poisson equation under different assumptions, and provide several upper bounds for $\bm{G}$}. The proof can be found in Appendix B.1.
Lemma 4.2 (Solution to the Poisson Equation).
Consider the matrix Poisson equation (14). Then
  1. (a)
    Under Assumption 2.1 and $\|\|{\bm{F}}\|\|_{\infty}<\infty$, we have that ${\bm{G}}=\sum_{k=0}^{\infty}{\mathrm{Q}}^{k}{\bm{F}}$ is a solution to the Poisson equation and satisfies $\|\|{\bm{G}}\|\|_{\infty}\leq(8/3)t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{\infty}$. Furthermore, if there exists an ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$, then ${\bm{G}}^{2}(z)\preccurlyeq(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$.
  2. (b)
    Under Assumption 2.2 and $\|\|{\bm{F}}\|\|_{V^{\alpha}}<\infty$ for some $\alpha\in(0,1]$, we have that ${\bm{G}}=\sum_{k=0}^{\infty}{\mathrm{Q}}^{k}{\bm{F}}$ is a solution to the Poisson equation and satisfies $\|{\bm{G}}(z)\|\leq[\varkappa^{\alpha}2^{3-\alpha}/(3\alpha)](\mu(V)+V(z))^{\alpha}t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{\alpha}}$ for any $z\in{\mathcal{Z}}$.
Using the Poisson equation Eqn. (14), we decompose the summation into a martingale, which is the dominant term, plus a remainder term. That is,
${\bm{S}}_{n}=\sum_{k=0}^{n-1}{{\bm{F}}}(Z_{k})=\sum_{k=0}^{n-1}\left[{{\bm{G}}(Z_{k})-({\mathrm{Q}}{\bm{G}})(Z_{k})}\right]=\sum_{k=1}^{n}{\bm{X}}_{k}+{\bm{G}}(Z_{0})-{\bm{G}}(Z_{n})={\bm{M}}_{k}+{\bm{G}}(Z_{0})-{\bm{G}}(Z_{n}),$
where ${\bm{X}}_{k}={\bm{G}}(Z_{k})-({\mathrm{Q}}{\bm{G}})(Z_{k-1})$ is a $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$-adapted martingale difference sequence, and we denote ${\bm{M}}_{k}=\sum_{j=1}^{k}{\bm{X}}_{j}$ as the related martingale and $\langle{\bm{M}}\rangle_{k}=\sum_{j=1}^{k}{\mathbb{E}}_{j-1}[({\bm{X}}_{j})^{2}]$ as the quadratic variation. We define ${\bm{M}}_{0}={\bm{X}}_{0}=\bm{0}_{d\times d}$. We first deal with stationary Markov chains, and then generalize the results to non-stationary chains.
Step 1: Stationary Markov Chains.
We assume $(Z_{k})_{k\in{\mathbb{N}}}$ is a stationary Markov chain, namely $Z_{0}\sim\mu$. By Minkowski inequality and Lemma 4.2, for all $p\geq 2$, we have
[…]
We start with the dimension-dependent results. For the martingale term, by the Rosenthal-Burkholder inequality (Theorem 2.1), we have
$\left\|{\sup_{k\leq n}\left\|{{\bm{M}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $C_{1}\sqrt{p\vee\log d}\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{n}}\right\|^{1/2}\Big\|_{\mu,p}+C_{2}(p\vee\log d)\left\|{\sup_{k\leq n}\|{\bm{X}}_{k}\|}\right\|_{\mu,p},$ (16)
where the quadratic variation term is dominant, because the second term is independent of $n$.
Now, we aim to give an upper bound for $\|\|\langle{\bm{M}}\rangle_{n}\|^{1/2}\|_{\mu,p}$. Note that $\langle{{\bm{M}}}\rangle_{n}=\sum_{k=1}^{n}{\mathbb{E}}_{k-1}[({\bm{X}}_{k})^{2}]$ and
${\mathbb{E}}_{k-1}\left[{\left({{\bm{X}}_{k}}\right)^{2}}\right]=$ ${\mathbb{E}}_{k-1}\left[{\left({{\bm{G}}(Z_{k})-({\mathrm{Q}}{\bm{G}})(Z_{k-1})}\right)^{2}}\right]$
$=$ $\left({{\mathrm{Q}}{\bm{G}}^{2}}\right)(Z_{k-1})-\left[{\left({{\mathrm{Q}}{\bm{G}}}\right)(Z_{k-1})}\right]^{2}$
$=$ ${\bm{H}}(Z_{k-1}),$
where we defined ${\bm{H}}(z)=({\mathrm{Q}}{\bm{G}}^{2})(z)-[({\mathrm{Q}}{\bm{G}})(z)]^{2}$ for any $z\in{\mathcal{Z}}$. According to the Poisson equation Eqn. (14) and the definition of ${\bm{\Sigma}}_{\mu}({\bm{F}})$ (Eqn. (3)), we have
$\mu\left({{\bm{H}}}\right)=$ $\mu\left({{\mathrm{Q}}{\bm{G}}^{2}-({\mathrm{Q}}{\bm{G}})^{2}}\right)$
$=$ $\mu\left({{\bm{G}}^{2}-\left({{\bm{G}}-{\bm{F}}}\right)^{2}}\right)$
$=$ $\mu\left({{{\bm{F}}}{\bm{G}}+{\bm{G}}{\bm{F}}-{{\bm{F}}}^{2}}\right)$
$=$ $\mu\left({{{\bm{F}}}^{2}}\right)+\sum_{k=1}^{\infty}\mu\left({{{\bm{F}}}{\mathrm{Q}}^{k}{{\bm{F}}}+\left({{\mathrm{Q}}^{k}{{\bm{F}}}}\right){{\bm{F}}}}\right)$
$=$ ${\bm{\Sigma}}_{\mu}({\bm{F}}).$
By Minkowski’s inequality and the inequality $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$, we have that
$\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{n}}\right\|^{1/2}\Big\|_{\mu,p}=$ $\sqrt{\left\|{\left\|{\sum_{k=1}^{n}{\bm{H}}(Z_{k-1})}\right\|}\right\|_{\mu,p/2}}$ (17)
$\leq$ $\sqrt{\left\|{n\left\|{\mu({\bm{H}})}\right\|+\left\|{\sum_{k=1}^{n}\left({{\bm{H}}(Z_{k-1})-\mu({\bm{H}})}\right)}\right\|}\right\|_{\mu,p/2}}$
$\leq$ $\sqrt{n\left\|{{\bm{\Sigma}}_{\mu}({\bm{F}})}\right\|+\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\mu,p/2}}$
$\leq$ $\sqrt{n}\left\|{{\bm{\Sigma}}_{\mu}({\bm{F}})}\right\|^{1/2}+\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\mu,p/2}^{1/2},$
where we denoted $\bar{{\bm{H}}}={{\bm{H}}}-\mu({\bm{H}})$.
Combining the upper bounds provided in Eqn. (15), Eqn. (16), Eqn. (17), we obtain
$\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $C_{1}\sqrt{n(p\vee\log d)}\left\|{{\bm{\Sigma}}_{\mu}({\bm{F}})}\right\|^{1/2}+C_{1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\mu,p/2}^{1/2}$ (18)
$+C_{2}(p\vee\log d)\left\|{\sup_{k\leq n}\|{\bm{X}}_{k}\|}\right\|_{\mu,p}+2\left\|{\left\|{{\bm{G}}(Z_{0})}\right\|}\right\|_{\mu,p},$
where the first term is the leading term of order $\sqrt{n}$, and the second term is of order $n^{1/4}$ because it involves the square root of the sum of $n$ zero-mean functions of the Markov chain. The remaining two terms are independent of $n$. Thus, the sum of these three residual terms is of order $O(n^{1/4})$.
Dimension-Free: To get a dimension-free counterpart of the inequality (18) under Assumption 2.1, the only difference is that we need to apply the dimension-free Rosenthal-Burkholder inequality to get a dimension-free counterpart of the inequality (16). To this end, we need an upper bound for the quadratic variation:
$\langle{{\bm{M}}}\rangle_{n}=\sum_{k=1}^{n}{\bm{H}}(Z_{k-1})=\sum_{k=1}^{n}\left\{{\left({{\mathrm{Q}}{\bm{G}}^{2}}\right)(Z_{k-1})-\left[{\left({{\mathrm{Q}}{\bm{G}}}\right)(Z_{k-1})}\right]^{2}}\right\}\preccurlyeq\sum_{k=1}^{n}\left({{\mathrm{Q}}{\bm{G}}^{2}}\right)(Z_{k-1}).$
If we assume Assumption 2.1 and ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$, then by Lemma 4.2 (a), it holds that $\left({{\mathrm{Q}}{\bm{G}}^{2}}\right)(Z_{k-1})\preccurlyeq(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}$, thus $\langle{{\bm{M}}}\rangle_{n}\preccurlyeq(64/9)nt_{\operatorname{mix}}^{2}{\bm{\Upsilon}}$. By Theorem 2.1 (b), the inequality (16) still holds if we replace $d$ by $d_{n,p}=e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\mu,\frac{p}{2}}])$, and double the right-hand side. As $n\to\infty$, by ergodicity and the fact that $\mu({\bm{H}})={\bm{\Sigma}}_{\mu}({\bm{F}})$, we find that $d_{n,p}\to e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge\|{\bm{\Sigma}}_{\mu}({\bm{F}})\|)$, which is a natural dimension factor.
Residual Terms: Next, we provide the specific forms of the upper bounds for these three residual terms under different assumptions. For the term of order $n^{1/4}$, we need the following lemma, which is a crude version of the Rosenthal inequality for uniformly geometrically ergodic Markov chains.
Lemma 4.3.
  1. (a)
    Under Assumption 2.1 and $\|\|{\bm{F}}\|\|_{\infty}<\infty$, we have that for any $p\geq 2$,
    $\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $\frac{16}{3}C_{1}\sqrt{(p\vee\log d)nt_{\operatorname{mix}}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}+\left({\frac{16}{3}C_{2}(p\vee\log d)+\frac{19}{3}}\right)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty},$
    Furthermore, if there exists an ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$, then the above inequality also holds when we replace $d$ by $e\cdot r([(64/9){\bm{\Upsilon}}]\wedge[(1/\lfloor{n}/{t_{\operatorname{mix}}}\rfloor)\|\|\langle{{\bm{M}}^{(t_{\operatorname{mix}})}}\rangle_{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\|\|_{\mu,\frac{p}{2}}])$, and double the right-hand side, where ${\bm{M}}^{(t_{\operatorname{mix}})}$ is a matrix martingale defined in Appendix B.2.
  2. (b)
    Under Assumption 2.2 and $\|\|{\bm{F}}\|\|_{V^{1/p}}<\infty$ for $p\geq 2$, we have that,
    $\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $\frac{8}{3}C_{1}p\sqrt{(p\vee\log d)nt_{\operatorname{mix}}}\left[{\varkappa\mu(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
    $+\left({\frac{16}{3}C_{2}(p\vee\log d)+\frac{11}{3}}\right)pt_{\operatorname{mix}}n^{1/p}\left[{\varkappa\mu(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$
The proof can be found in Appendix B.2.
Under Assumption 2.1 and $\|\|{\bm{F}}\|\|_{\infty}<\infty$, by Lemma 4.2 (a) and Lemma 4.3 (a), $\|\|\bar{{\bm{H}}}\|\|_{\infty}\leq(64/9)t_{\operatorname{mix}}^{2}\|\|{{\bm{F}}}\|\|_{\infty}$ and the term of order $n^{1/4}$ has the following upper bound:
$C_{1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\mu,p/2}^{1/2}\lesssim\left({p\vee\log d}\right)^{3/4}t_{\operatorname{mix}}^{5/4}n^{1/4}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}+\left({p\vee\log d}\right)t_{\operatorname{mix}}^{3/2}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$ (19)
And it is easy to see that the other two residual terms are dominated by the term. If we further assume ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$, then by Lemma 4.2 (a), the above inequality still holds if we replace $d$ introduced in the l.h.s. by $e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\mu,\frac{p}{2}}])$, and $d$ introduced in the r.h.s. (see Lemma 4.3) by $e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/\lfloor{n}/{t_{\operatorname{mix}}}\rfloor)\|\|\langle{{\bm{M}}^{(t_{\operatorname{mix}})}_{{\bm{H}}}}\rangle_{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\|\|_{\mu,\frac{p}{4}}])$, and double the r.h.s., where the matrix martingale ${\bm{M}}^{(t_{\operatorname{mix}})}_{\bar{\bm{H}}}$ is defined in Appendix B.2 with $\bm{F}$} replaced by $\bar{\bm{H}}$.
Under Assumption 2.2 and $\|\|{\bm{F}}\|\|_{V^{1/p}}<\infty$, by Lemma 4.2 (b), $\|\|\bar{{\bm{H}}}\|\|_{V^{2/p}}^{1/2}$ has the following upper bound:
$\left\|{\left\|{\bar{{\bm{H}}}}\right\|}\right\|_{V^{2/p}}^{1/2}\leq$ $\left\|{\left\|{{{\bm{G}}}}\right\|^{2}}\right\|_{V^{2/p}}^{1/2}$
$=$ $\left\|{\left\|{{{\bm{G}}}}\right\|}\right\|_{V^{1/p}}$
$\leq$ $\sup_{z\in{\mathcal{Z}}}\frac{[p\varkappa^{1/p}2^{3-1/p}/3](\mu(V)+V(z))^{1/p}t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{1/p}}}{V^{1/p}(z)}$
$\lesssim$ $p\left[{\varkappa\mu(V)}\right]^{1/p}t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{1/p}},$
and by Lemma 4.3 (b) the term of order $n^{1/4}$ has the following upper bound:
$C_{1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\mu,p/2}^{1/2}\lesssim$ $p^{3/2}\left({p\vee\log d}\right)^{3/4}t_{\operatorname{mix}}^{5/4}n^{1/4}\left[{\varkappa\mu(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
$+p^{3/2}\left({p\vee\log d}\right)t_{\operatorname{mix}}^{3/2}n^{1/p}\left[{\varkappa\mu(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$
And it is easy to see that the other two residual terms are dominated by the term.
Step 2: Generalize to Non-Stationary Markov Chains
In order to generalize the result to non-stationary Markov chains, we need to leverage the maximal exact coupling argument (Douc et al., 2018, Lemma 19.3.6 and Lemma 19.3.9). Following (Durmus et al., 2023, Section 4), under Assumption 2.1 and $\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}<\infty$, we obtain
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}+\frac{2}{n}\left({1+128(t_{\operatorname{mix}}/\log 4)^{p}(p+1)^{p+1/2}e^{-p}}\right)^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
Under Assumption 2.2 and $\|\|{\bm{F}}\|\|_{V^{1/p}}<\infty$, we obtain
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}+\frac{32}{15n}\varkappa^{1/p}\left({\xi(V)+\mu(V)}\right)^{1/p}pt_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$
5 Conclusions
In this work, we have studied probability bounds for the spectral norm of sums of dependent random matrices. First, we have proved a Rosenthal-Burkholder inequality for matrix martingales (Theorem 2.1) using the extrapolation method (good $\lambda$ inequality) (Hitczenko, 1990, 1994; Pinelis, 1994) and the matrix Bennett inequality (Tropp, 2011; Kroshnin and Suvorikova, 2024). By discretization, we have shown a sharp Burkholder-Davis-Gundy inequality for continuous matrix local martingales (Theorem 2.2) as a corollary of Theorem 2.1. Furthermore, combining this martingale inequality with the Poisson equation approach (Durmus et al., 2023), we have derived matrix Rosenthal inequalities for Markov chains under both the uniform geometric ergodicity assumption and the geometric $V$-ergodicity assumption in Section 2.2. By the Markov inequality, these moment inequalities can be further translated into concentration inequalities including Hoeffding and Bernstein inequalities. Through more careful analysis, we have also established dimension-free versions of the above inequalities.
Compared with previous works on matrix concentration inequalities for Markov chains, our results yield a sharp leading term that matches the Markov chain CLT, rather than relying on a sub-optimal variance proxy. Our results do not require the assumption of a non-zero absolute spectral gap or $\psi$-mixing condition. This expands its applicability to a broader range of problems, such as policy evaluation in reinforcement learning (Duan and Wainwright, 2022) and PCA (Kumar and Sarkar, 2023) with Markovian data. And our results do not need the assumption of the boundedness of functions under the geometric ergodicity assumption. Moreover, our results can be extended to dimension-free versions, enabling applications to infinite-dimensional settings. The advantages of our results over previous works are demonstrated in the covariance estimation and PCA problems on Markovian data. And as mentioned in this paper, our results can also be extended to more general geometrically ergodic Markov chains with respect to the Wasserstein semimetric, as well as to geometrically ergodic continuous Markov processes.
One direction for future work is to explore whether we can obtain residual terms with sharp dependence on $t_{\operatorname{mix}}$. This might require more sophisticated tools from matrix analysis, beyond the properties of self-adjoint matrices related to the Löwner order. Another promising direction is to extend our results to time-inhomogeneous functions while retaining the advantages mentioned above. For time-inhomogeneous functions, the standard Poisson equation approach no longer works, and more advanced tools from Markov chain theory would be necessary. Finally, an interesting question is: do certain form of universality-based probabilistic inequalities hold for martingales and ergodic Markov chains? The universality-based inequalities may yield tighter error bounds in some applications, and we would leave the question for future work.
Appendix A Omitted Proof for the Rosenthal-Burkholder Inequality for Matrix Martingales
In this section, we complete the proofs omitted in Section 4.1.
A.1 Proof of Theorem 4.1
Proof.
We first recall the Bennett inequality for matrix martingales with bounded martingale difference, which is provided in Tropp [2011, Theorem 3.1] Kroshnin and Suvorikova [2024, Theorem 2.3].
Lemma A.1 (Bennett Inequality for Matrix Martingale).
Suppose that there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that $\langle{\bm{M}}\rangle_{\infty}\preccurlyeq{\bm{\Upsilon}}$ almost surely, and $\|{\bm{X}}_{k}\|\leq U$ for any $k\in{\mathbb{N}}$ almost surely for some $U>0$. Then for any $\varepsilon>0$, it holds that
${\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\left\|{{\bm{M}}_{k}}\right\|\geq\varepsilon}\right)\leq$ $2(d\wedge(e\cdot r({\bm{\Upsilon}})))\exp\left\{{-\frac{\left\|{{\bm{\Upsilon}}}\right\|}{U^{2}}h\left({\frac{U\varepsilon}{\left\|{{\bm{\Upsilon}}}\right\|}}\right)}\right\}\leq 2(d\wedge(e\cdot r({\bm{\Upsilon}})))\left({\frac{e\left\|{{\bm{\Upsilon}}}\right\|}{\varepsilon U}}\right)^{{\varepsilon}/{U}},$
where $h(x)=(1+x)\log(1+x)-x$ for any $x>0$.
For any $k\in{\mathbb{N}}$, let $\bar{{\bm{X}}}_{k}={\bm{X}}_{k}\mathds{1}\{\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda\}$ be the truncated version of ${\bm{X}}_{k}$, $\bar{{\bm{M}}}_{k}=\sum_{j=0}^{k}\bar{{\bm{X}}}_{j}$ be the truncated martingale, and $\langle\bar{{\bm{M}}}\rangle_{k}=\sum_{j=0}^{k}{\mathbb{E}}_{j-1}\bar{{\bm{X}}}_{j}^{2}\preccurlyeq\langle{\bm{M}}\rangle_{k}$ be the corresponding quadratic variation sequence. $\bar{{\bm{M}}}_{k}$ is indeed a martingale because
${\mathbb{E}}_{k-1}\left[{\bar{{\bm{X}}}_{k}}\right]=$ ${\mathbb{E}}_{k-1}\left[{{\bm{X}}_{k}\mathds{1}\{\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda\}}\right]$
$=$ ${\mathbb{E}}_{k-1}\left[{{{\bm{X}}}_{k}}\right]-{\mathbb{E}}_{k-1}\left[{{\bm{X}}_{k}\mathds{1}\{\|{\bm{X}}_{k}\|>\delta_{2}\lambda\}}\right]$
$=$ ${\mathbb{E}}_{k-1}\left[{{\bm{X}}_{k}\mathds{1}\{\|{\bm{X}}_{k}\|>\delta_{2}\lambda\}}\right]$
$=$ ${\mathbb{E}}_{k-1}\left[{{{\bm{X}}}_{k}}\right]-{\mathbb{E}}_{k-1}\left[{{\bm{X}}_{k}\mathds{1}\{\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda\}}\right]$
$=$ $-{\mathbb{E}}_{k-1}\left[{\bar{{\bm{X}}}_{k}}\right]$
$=$ $\bm{0}_{d\times d},$
where we used the fact that $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ is a martingale and it is conditionally symmetric. By definition of the truncated martingale, on the event $\{\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda\}$, we have ${\bm{M}}_{k}=\bar{{\bm{M}}}_{k}$ for any $k\in{\mathbb{N}}$.
Next, we define three stopping times: $\tau_{1}=\inf\{k\in{\mathbb{N}}\colon\|{{\bm{M}}_{k}}\|>\lambda\}$, $\tau_{2}=\inf\{k\in{\mathbb{N}}\colon\|{{\bm{M}}_{k}}\|>\beta\lambda\}$ ($\tau_{1}\leq\tau_{2}$), $\tau_{3}=\inf\{k\in{\mathbb{N}}\colon\|\langle\bar{{\bm{M}}}\rangle_{k+1}\|^{1/2}>\delta_{1}\lambda\}$. By definition of the stopping times, we have
[…]
We also consider the stopped martingale $\hat{{\bm{M}}}_{k}=\bar{{\bm{M}}}_{(k\wedge\tau_{2}\wedge\tau_{3})\vee\tau_{1}}-\bar{{\bm{M}}}_{\tau_{1}}$ conditionally on the sigma-algebra of the $\tau_{1}$-past
${\mathscr{F}}_{\tau_{1}}=\left\{{A\in{\mathscr{F}}\colon A\cap\{\tau_{1}\leq k\}\in{\mathscr{F}}_{k},\quad\forall k\in{\mathbb{N}}}\right\}.$
That is $(\hat{{\bm{M}}}_{k})_{k\in{\mathbb{N}}}$ is a $({\mathscr{F}}_{k\vee\tau_{1}})_{k\in{\mathbb{N}}}$-adapted martingale with the martingale difference $\hat{{\bm{X}}}_{k}=\hat{{\bm{M}}}_{k}-\hat{{\bm{M}}}_{k-1}=\bar{{\bm{X}}}_{k}\mathds{1}\{\tau_{1}<k<\tau_{2}\wedge\tau_{3}\}$ upper-bounded by $\delta_{2}\lambda$ and square root of the quadratic variation upper-bounded by $\delta_{1}\lambda$.
Now, we are ready to derive the good $\lambda$ inequality.
${\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|>\beta\lambda,\ \left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\leq\delta_{1}\lambda,\ \sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda}\right)$ (21)
$\qquad\leq{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\|{{\bm{M}}}_{k}\|>\beta\lambda,\ \left\|{\left\langle{\bar{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\leq\delta_{1}\lambda,\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda}\right)$
$\qquad\leq{\mathbb{P}}\left({\tau_{2}<\infty,\tau_{3}=\infty,\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda}\right)$
$\qquad\leq{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\|\hat{{\bm{M}}}_{k}\|>(\beta-1-\delta_{2})\lambda}\right)$
$\qquad={\mathbb{E}}\left[{{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\|\hat{{\bm{M}}}_{k}\|>(\beta-1-\delta_{2})\lambda\Big|{\mathscr{F}}_{\tau_{1}}}\right)\mathds{1}\left\{{\tau_{1}<\infty}\right\}}\right]$
$\qquad\leq 2d\left({\frac{e}{N}\frac{\delta_{1}^{2}}{\delta_{2}^{2}}}\right)^{N}{\mathbb{E}}\left[{\mathds{1}\left\{{\tau_{1}<\infty}\right\}}\right]$
$\qquad=2d\left({\frac{e}{N}\frac{\delta_{1}^{2}}{\delta_{2}^{2}}}\right)^{N}{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|>\lambda}\right),$
which is the desired conclusion. Here, we explain some details in the derivation of Eqn (21). In the first inequality, we used the fact that $\langle\bar{{\bm{M}}}\rangle_{\infty}\preccurlyeq\langle{\bm{M}}\rangle_{\infty}$ almost surely. In the second inequality, we used Eqn. (20). In the third inequality, we used the fact that on the event $\{\tau_{2}<\infty,\tau_{3}=\infty,\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda\}$, it holds that $\sup_{k<\tau_{1}}\left\|{{{\bm{M}}}_{k}}\right\|=\sup_{k<\tau_{1}}\left\|{\bar{{\bm{M}}}_{k}}\right\|<\lambda$ and $\left\|{\bar{{\bm{M}}}_{(\tau_{2}\wedge\tau_{3})-1}}\right\|>\left\|{{{\bm{M}}}_{\tau_{2}}}\right\|-\left\|{{\bm{X}}_{\tau_{2}}}\right\|>(\beta-\delta_{2})\lambda$. And in the last inequality, we used the matrix Bennett inequality (Lemma A.1) with error term $\varepsilon=(\beta-1-\delta_{2})\lambda$, parameter $U=\delta_{2}\lambda$, and $\|\langle\hat{{\bm{M}}}_{k}\rangle_{\infty}\|\leq\delta_{1}^{2}\lambda^{2}$ given ${\mathcal{F}}_{\tau_{1}}$.
If we further assume that there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that $\langle{\bm{M}}\rangle_{\infty}\preccurlyeq{\bm{\Upsilon}}$ almost surely, then the fourth inequality also holds if we replace $d$ with $e\cdot r({\bm{\Upsilon}}\wedge(\delta_{1}^{2}\lambda^{2}))$ by the dimension-free matrix Bennett inequality (Lemma A.1). ∎
A.2 Proof of Theorem 2.2
Proof.
We first consider $\tau=\infty$ and the bounded case: $({\bm{M}}_{t})_{t\geq 0}$ is a continuous martingale satisfying $\sup_{t\geq 0}\left\|{{\bm{M}}_{t}}\right\|\leq K,\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|\leq K$ for some constant $K>0$. For a fixed discretization level $n\in{\mathbb{N}}$, we define ${\bm{M}}_{k}^{(n)}={\bm{M}}_{\frac{k}{2^{n}}}$ for any $k\in{\mathbb{N}}$ (${\bm{M}}_{-1}^{(n)}:=\bm{0}_{d{\times}d}$). Apply Theorem 2.1 to $({\bm{M}}_{k}^{(n)})_{k\in{\mathbb{N}}}$, we have
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{\frac{k}{2^{n}}}\|}\right\|_{p}$ (22)
$\leq$ $C_{R,1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right\|^{1/2}}\right\|_{p}+C_{R,2}(p\vee\log d)\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}}\|}\right\|_{p}.$
Now we take the limit of each part of the above inequality as $n\to\infty$. By Beppo Levi’s Theorem, the left hand side converges to $\left\|{\sup_{t\geq 0}\|{\bm{M}}_{t}\|}\right\|_{p}$. By dominated convergence theorem, the second term of the right hand side converges to 0. It suffices to prove $\lim_{n\to\infty}\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]=\left\langle{{\bm{M}}}\right\rangle_{\infty}$ in $L^{\frac{p}{2}}$. Here the convergence of matrices is defined entry-wise.
Lemma A.2.
[…]
Proof.
${\mathbb{E}}\left\|{\sum_{k=0}^{\infty}({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}}\right\|^{2}$ $\leq{\mathbb{E}}\left({\operatorname{Tr}\left({{\sum_{k=0}^{\infty}({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}}}\right)}\right)^{2}$ (24)
$={\mathbb{E}}\left({\sum_{i,j=1}^{d}{\sum_{k=0}^{\infty}\left({[{\bm{M}}_{\frac{k}{2^{n}}}]_{ij}-[{\bm{M}}_{\frac{k-1}{2^{n}}}]_{ij}}\right)^{2}}}\right)^{2}$
$\leq d^{2}\sum_{i,j=1}^{d}{\mathbb{E}}\left({{\sum_{k=0}^{\infty}\left({[{\bm{M}}_{\frac{k}{2^{n}}}]_{ij}-[{\bm{M}}_{\frac{k-1}{2^{n}}}]_{i,j}}\right)^{2}}}\right)^{2}.$
By [Karatzas and Shreve, 1988, Lemma 1.5.9], each term in the last summation is upper bounded by $48K^{4}$. ∎
Lemma A.3.
$\lim_{n\to\infty}{\mathbb{E}}\left\|{\sum_{k=0}^{\infty}({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{4}}\right\|=0$ (25)
Proof.
${\mathbb{E}}\left\|{\sum_{k=0}^{\infty}({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{4}}\right\|$ $\leq{\mathbb{E}}\left\|{\sum_{k=0}^{\infty}\left({\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}}\|^{2}}\right)({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}}\right\|$ (26)
$={\mathbb{E}}\left({\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}}\|^{2}}\right)\left\|{\sum_{k=0}^{\infty}({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}}\right\|$
$\leq\left({{\mathbb{E}}\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}}\|^{4}}\right)^{{1}/{2}}\left({{\mathbb{E}}\left\|{\sum_{k=0}^{\infty}({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}}\right\|^{2}}\right)^{{1}/{2}}$
$\leq\sqrt{48}d^{2}K^{2}\left({{\mathbb{E}}\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}}\|^{4}}\right)^{{1}/{2}}.$
In the last step, we use the upper bound in Lemma A.2. The expectation in the last step converges to 0 by dominated convergence theorem. ∎
Lemma A.4.
$\lim_{n\to\infty}\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]=\left\langle{{\bm{M}}}\right\rangle_{\infty}\ {\rm in}\ L^{2}.$ (27)
Proof.
Note that $(({\mathbb{E}}[({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}]-(\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}})))_{k\in{\mathbb{N}}}$ is a martingale difference sequence. Thus,
${\mathbb{E}}\left\|{\sum_{k=0}^{\infty}\left({{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]-(\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}})}\right)}\right\|_{F}^{2}$ (28)
$=$ $\sum_{k=0}^{\infty}{\mathbb{E}}\left\|{\left({{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]-(\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}})}\right)}\right\|_{F}^{2}$
$\leq$ $2\sum_{k=0}^{\infty}{\mathbb{E}}\left\|{{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right\|_{F}^{2}+2\sum_{k=0}^{\infty}{\mathbb{E}}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}}}\right\|_{F}^{2}$
$\leq$ $2\sum_{k=0}^{\infty}{\mathbb{E}}\left\|{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}}\right\|_{F}^{2}+2\sum_{k=0}^{\infty}{\mathbb{E}}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}}}\right\|\operatorname{Tr}\left({\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}}}\right)$
$\leq$ $2\sum_{k=0}^{\infty}{\mathbb{E}}\operatorname{Tr}\left({{\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}}}\right)^{4}+2\sum_{k=0}^{\infty}{\mathbb{E}}\left({\sup_{k\in{\mathbb{N}}}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}}}\right\|}\right)\operatorname{Tr}\left({\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}}}\right)$
$=$ $2{\mathbb{E}}\operatorname{Tr}\sum_{k=0}^{\infty}({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{4}+2{\mathbb{E}}\left({\sup_{k\in{\mathbb{N}}}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}}}\right\|}\right)\operatorname{Tr}(\left\langle{{\bm{M}}}\right\rangle_{\infty}),$
where the first term converges to 0 by Lemma A.3, and the second term converges to 0 by the dominated convergence theorem. ∎
Lemma A.5.
$\left({\left\|{\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right\|^{\frac{p}{2}}}\right)_{n\in{\mathbb{N}}}$ (29)
is uniformly integrable.
Proof.
Take $q>\frac{p}{2}$. It suffices to prove
$\left({{\mathbb{E}}\left\|{\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right\|^{q}}\right)_{n\in{\mathbb{N}}}$ (30)
is uniformly bounded.
${\mathbb{E}}\left\|{\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right\|^{q}$ $\leq{\mathbb{E}}\left({\operatorname{Tr}\left({\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right)}\right)^{q}$ (31)
$={\mathbb{E}}\left({\sum_{i,j=1}^{d}\sum_{k=0}^{\infty}{\mathbb{E}}\left[{([{\bm{M}}_{\frac{k}{2^{n}}}]_{ij}-[{\bm{M}}_{\frac{k-1}{2^{n}}}]_{ij})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right)^{q}$
$\leq d^{2(q-1)}\sum_{i,j=1}^{d}{\mathbb{E}}\left({\sum_{k=0}^{\infty}{\mathbb{E}}\left[{\left({[{\bm{M}}_{\frac{k}{2^{n}}}]_{ij}-[{\bm{M}}_{\frac{k-1}{2^{n}}}]_{ij}}\right)^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right)^{q}.$
Apply scalar Rosenthal-Burkholder inequality [Burkholder, 1973] to $([{\bm{M}}_{k}^{(n)}]_{ij})_{k\in{\mathbb{N}}}$:
${\mathbb{E}}\left({\sum_{k=0}^{\infty}{\mathbb{E}}\left[{\left({[{\bm{M}}_{\frac{k}{2^{n}}}]_{ij}-[{\bm{M}}_{\frac{k-1}{2^{n}}}]_{ij}}\right)^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right)^{q}\leq C_{q}K^{2q}.$ (32)
Thus
$\left({{\mathbb{E}}\left\|{\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right\|^{q}}\right)_{n\in{\mathbb{N}}}$ (33)
is uniformly bounded by $d^{2q}C_{q}K^{2q}$. ∎
By Lemma A.3 and Lemma A.4, $\lim_{n\to\infty}\sum_{k=1}^{2^{n}}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}T}-{\bm{M}}_{\frac{k-1}{2^{n}}T})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}T}}}\right]=\left\langle{{\bm{M}}}\right\rangle_{T}$ in $L^{\frac{p}{2}}$.
Now suppose that $({\bm{M}}_{t})$ is a continuous local martingale that is not necessarily bounded. We use the standard localization technique. Define a sequence of stopping times $\tau_{m}={\rm inf}\{t\geq 0:\left\|{{\bm{M}}_{t}}\right\|\geq m\text{ or }\left\|{\left\langle{{\bm{M}}}\right\rangle_{t}}\right\|\geq m\}$. Then ${\bm{M}}_{t}^{(m)}:={\bm{M}}_{t\wedge\tau\wedge\tau_{m}}$ is a proper martingale. By the previous results, we have
$\left\|{\sup_{t\geq 0}\|{\bm{M}}_{t\wedge\tau\wedge\tau_{m}}\|}\right\|_{p}\leq C_{R,1}\sqrt{p\vee\log d}\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau\wedge\tau_{m}}}\right\|^{1/2}\Big\|_{p}\leq C_{R,1}\sqrt{p\vee\log d}\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|^{1/2}\Big\|_{p}.$ (34)
Let $m\to\infty$ and we finish the proof. ∎
A.3 Proof of Lemma 4.1
Proof.
Consider the Lebesgue-Stieltjes integral, we have
${\mathbb{E}}\left[{\Phi(Y)}\right]=$ ${\mathbb{E}}\left[{\Phi(Y)\mathds{1}\left\{{Y\leq a}\right\}}\right]+{\mathbb{E}}\left[{\Phi(Y)\mathds{1}\left\{{Y>a}\right\}}\right]$
$\leq$ $\Phi(a)+{\mathbb{E}}\left[{\Phi(Y)\mathds{1}\left\{{Y>a}\right\}}\right],$
and by the good $\lambda$ inequality, we further have
${\mathbb{E}}\left[{\Phi(Y)\mathds{1}\left\{{Y>a}\right\}}\right]=$ ${\mathbb{E}}\left[{\Phi(\beta\beta^{-1}Y)\mathds{1}\left\{{Y>a}\right\}}\right]$
$\leq$ $\gamma{\mathbb{E}}\left[{\Phi(\beta^{-1}Y)\mathds{1}\left\{{Y>a}\right\}}\right]$
$=$ $\gamma\int_{{a}/{\beta}}^{\infty}{\mathbb{P}}\left({Y>\beta\lambda}\right)\Phi(d\lambda)$
$=$ $\gamma\int_{{a}/{\beta}}^{\infty}\left[{{\mathbb{P}}\left({Y>\beta\lambda,Z\leq\lambda}\right)+{\mathbb{P}}\left({Y>\beta\lambda,Z>\lambda}\right)}\right]\Phi(d\lambda)$
$\leq$ $\gamma\int_{{a}/{\beta}}^{\infty}\left[{\epsilon{\mathbb{P}}\left({Y>\lambda}\right)+{\mathbb{P}}\left({Z>\lambda}\right)}\right]\Phi(d\lambda)$
$\leq$ $\gamma\epsilon{\mathbb{E}}\left[{\Phi(Y)\mathds{1}\left\{{Y>{a}/{\beta}}\right\}}\right]+\gamma{\mathbb{E}}\left[{\Phi(Z)}\right]$
$\leq$ $\gamma\epsilon{\mathbb{E}}\left[{\Phi(Y)\mathds{1}\left\{{Y>a}\right\}}\right]+\gamma\epsilon\Phi(a)+\gamma{\mathbb{E}}\left[{\Phi(Z)}\right]$
$\leq$ $\frac{\gamma\epsilon}{1-\gamma\epsilon}\Phi(a)+\frac{\gamma}{1-\gamma\epsilon}{\mathbb{E}}\left[{\Phi(Z)}\right].$
hence
${\mathbb{E}}\left[{\Phi(Y)}\right]\leq\frac{1}{1-\gamma\epsilon}\Phi(a)+\frac{\gamma}{1-\gamma\epsilon}{\mathbb{E}}\left[{\Phi(Z)}\right],$
which is the desired conclusion. ∎
A.4 Remaining Proof of Theorem 2.1
We continue from the Step 2 in Section 4.1. In this section, we only prove the dimension-dependent version of the inequality. For the dimension-free version, it can be obtained by repeating the proof line by line as presented in this section, and we omit it for brevity.
First, let us recall some important concepts. The first one is the tangent sequence, which is used to construct conditionally symmetric martingales. The second is the conditional independence (CI) condition; the additional random components introduced in constructing conditionally symmetric martingales will satisfy this condition as shown in Theorem A.6, and sequences satisfying the CI condition can be treated as independent sequences.
Definition A.1.
[Hitczenko, 1990, Tangent Sequences, Definition 2.1] Two $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$-adapted sequences $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$ and $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ are said to be tangent if the conditional distributions of ${\bm{X}}_{k}$ and ${\bm{Y}}_{k}$ given ${\mathscr{F}}_{k-1}$ coincide almost surely for any $k\in{\mathbb{N}}$.
From this definition, it follows that $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$ is conditionally symmetric if and only if $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$ and $(-{\bm{X}}_{k})_{k\in{\mathbb{N}}}$ are tangent. Furthermore, suppose that there exists a sequence $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ tangent to $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$, and ${\bm{X}}_{k},{\bm{Y}}_{k}$ are ${\mathscr{F}}_{k-1}$-conditionally independent for any $k\in{\mathbb{N}}$, it holds that $(\sum_{j=0}^{k}({\bm{X}}_{j}-{\bm{Y}}_{j}))_{k\in{\mathbb{N}}}$ is a conditionally symmetric martingale.
Definition A.2.
[Hitczenko, 1990, Conditional Independence (CI) Condition, Definition 2.2] An $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$-adapted sequences $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ satisfies (CI) condition if there exists a $\sigma$-algebra $\mathscr{G}$$\subset{\mathscr{F}}$ such that the conditional distributions of ${\bm{Y}}_{k}$ given ${\mathscr{F}}_{k-1}$ and $\mathscr{G}$} coincide almost surely for any $k\in{\mathbb{N}}$, and the sequence $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ is $\mathscr{G}$}-conditionally independent.
The following lemma guarantees the existence of the desired sequence $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$.
Lemma A.6.
[Hitczenko, 1990, Lemma 2.3] In some enlarged probability space of $(\Omega,{\mathscr{F}},{\mathbb{P}})$, there exists a sequence $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ which is tangent to $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$, satisfies the (CI) condition with $\sigma$-algebra $\mathscr{G}$}, and satisfies ${\bm{X}}_{k},{\bm{Y}}_{k}$ are ${\mathscr{F}}_{k-1}$-conditionally independent for any $k\in{\mathbb{N}}$.
The statement and proof of this lemma require rigorous measure-theoretic language (such as regular conditional probabilities). To avoid introducing excessive notations, we omit the details and merely explain the idea of constructing $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ in words. Interested readers may refer to Hitczenko [1990, Lemma 2.3]. For simplicity, from now on, we shall not change the notations even though we will work in the enlarged probability space. At time 0, we generate an independent copy ${\bm{Y}}_{0}$ of ${\bm{X}}_{0}$ (in an enlarged probability space to accommodate the sequence $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$). At time $k\geq 1$, given ${\mathscr{F}}_{k-1}$, we generate a copy ${\bm{Y}}_{k}$ that is conditionally independent of ${\bm{X}}_{k}$. It is easy to see that $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ is tangent to $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$ and satisfies the (CI) condition with $\mathscr{G}$$\mathscr{F}$}.
Based on the sequence $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ given in Theorem A.6, we introduce some notations. For any $k\in{\mathbb{N}}$, we denote ${\bm{N}}_{k}=\sum_{j=0}^{k}{\bm{Y}}_{j}$, $\tilde{{\bm{X}}}_{k}={\bm{X}}_{k}-{\bm{Y}}_{k}$, $\tilde{{\bm{M}}}_{k}=\sum_{j=0}^{k}\tilde{{\bm{X}}}_{k}={\bm{M}}_{k}-{\bm{N}}_{k}$. Then $({\bm{N}}_{k})_{k\in{\mathbb{N}}}$ is a martingale with (CI) increments, and we denote by $(\langle{\bm{N}}\rangle_{k})_{k\in{\mathbb{N}}}$ the related quadratic variation process. In addition, $(\tilde{{\bm{M}}}_{k})_{k\in{\mathbb{N}}}$ is a conditionally symmetric martingale, and we denote by $(\langle\tilde{\bm{M}}\rangle_{k})_{k\in{\mathbb{N}}}$ the related quadratic variation process.
Because ${\bm{X}}_{k},{\bm{Y}}_{k}$ are ${\mathscr{F}}_{k-1}$-conditionally independent, we have $\langle\tilde{\bm{M}}\rangle_{k}=\langle{\bm{M}}\rangle_{k}+\langle{\bm{N}}\rangle_{k}$ almost surely. In addition, because $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ is tangent to $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$, we have $\langle{\bm{M}}\rangle_{k}=\langle{\bm{N}}\rangle_{k}$ almost surely, thus $\langle\tilde{\bm{M}}\rangle_{k}=2\langle{\bm{M}}\rangle_{k}$ almost surely.
By Minkowski’s inequality, we have
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq\left\|{\sup_{k\in{\mathbb{N}}}\|\tilde{\bm{M}}_{k}\|}\right\|_{p}+\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{N}}_{k}\|}\right\|_{p}.$
According to Rosenthal-Burkholder inequality for conditionally symmetric martingale (Eqn. (13)), we have
[…]
where
$\left\|{\left\|{\left\langle{\tilde{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}}\right\|_{p}=\sqrt{2}\left\|{\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}}\right\|_{p},$
and
$\left\|{\sup_{k\in{\mathbb{N}}}\left\|{\tilde{\bm{X}}_{k}}\right\|}\right\|_{p}\leq\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|}\right\|_{p}+\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{Y}}_{k}}\right\|}\right\|_{p}.$
To give an upper bound for $\|\sup_{k\in{\mathbb{N}}}\|{\bm{Y}}_{k}\|\|_{p}$, we need the following lemma.
Lemma A.7.
[Hitczenko, 1988, Lemma 1] Suppose that $A_{k},B_{k}\in{\mathscr{F}}_{k}$ satisfy ${\mathbb{P}}(A_{k}\mid{\mathscr{F}}_{k-1})={\mathbb{P}}(B_{k}\mid{\mathscr{F}}_{k-1})$ for any $k\in{\mathbb{N}}$. Then it holds that
${\mathbb{P}}\left({\bigcup_{k\in{\mathbb{N}}}A_{k}}\right)\leq 2{\mathbb{P}}\left({\bigcup_{k\in{\mathbb{N}}}B_{k}}\right).$
For any $t>0$, we define the events $A_{k}=\{\|{\bm{Y}}_{k}\|\geq t\}$, $B_{k}=\{\|{\bm{X}}_{k}\|\geq t\}$ which satisfy ${\mathbb{P}}(A_{k}\mid{\mathscr{F}}_{k-1})={\mathbb{P}}(B_{k}\mid{\mathscr{F}}_{k-1})$ because $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ is tangent to $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$. By Lemma A.7, we obtain
${\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\left\|{{\bm{Y}}_{k}}\right\|\geq t}\right)=$ ${\mathbb{P}}\left({\bigcup_{k\in{\mathbb{N}}}A_{k}}\right)\leq 2{\mathbb{P}}\left({\bigcup_{k\in{\mathbb{N}}}B_{k}}\right)=2{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|\geq t}\right),$
hence
$\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{Y}}_{k}}\right\|}\right\|_{p}^{p}=$ ${\mathbb{E}}\left[{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{Y}}_{k}}\right\|^{p}}\right]$
$=$ $\int_{0}^{\infty}{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\left\|{{\bm{Y}}_{k}}\right\|^{p}\geq u}\right)du$
$=$ $p\int_{0}^{\infty}{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\left\|{{\bm{Y}}_{k}}\right\|\geq t}\right)t^{p-1}dt$
$\leq$ $2p\int_{0}^{\infty}{\mathbb{P}}\left({\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|\geq t}\right)t^{p-1}dt$
$=$ $2\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|}\right\|_{p}^{p},$
which implies $\|\sup_{k\in{\mathbb{N}}}\|{\bm{Y}}_{k}\|\|_{p}\leq 2^{1/p}\|\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\|_{p}$.
Therefore, we have
[…]
We only need to give an upper bound for $\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{N}}_{k}\|}\right\|_{p}$. Since $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ satisfies (CI) condition, the proof strategy is almost identical to that of an independent sum. The only difference is that we need to replace the expectation with the mathematical expectation conditional on $\mathscr{G}$}. The key tool is the following maximal version of the symmetrization inequality [Ledoux and Talagrand, 2013, Lemma 6.3].
Lemma A.8 (Maximal Symmetrization Inequality).
Let $F\colon{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$ be a non-decreasing convex function, and $({\mathcal{X}},\|\cdot\|)$ be a separable Banach space. Suppose that $(X_{k})_{k\in{\mathbb{N}}}$ is a sequence of independent $\mathcal{X}$}-valued mean zero random elements such that ${\mathbb{E}}[F(\sup_{k\in{\mathbb{N}}}\|\sum_{j=0}^{k}X_{j}\|)]<\infty$ for any $k\in{\mathbb{N}}$, and $(\epsilon_{k})_{k\in{\mathbb{N}}}$ is a sequence of i.i.d. Rademacher random variables (${\mathbb{P}}(\epsilon_{1}=1)={\mathbb{P}}(\epsilon_{1}=-1)=1/2$) independent of $(X_{k})_{k\in{\mathbb{N}}}$. Then it holds that
${\mathbb{E}}\left[{F\left({\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}X_{j}}\right\|}\right)}\right]\leq{\mathbb{E}}\left[{F\left({2\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}X_{j}}\right\|}\right)}\right].$
The proof can be found in Appendix A.5.
By applying Lemma A.8 with $F(x)=x^{p}$, we have
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{N}}_{k}\|}\right\|_{p}^{p}=$ ${\mathbb{E}}\left[{\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}{\bm{Y}}_{j}}\right\|^{p}}\right]$
$=$ ${\mathbb{E}}\left[{{\mathbb{E}}\left[{\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}{\bm{Y}}_{j}}\right\|^{p}\Bigg|{\mathscr{G}}}\right]}\right]$
$\leq$ $2^{p}{\mathbb{E}}\left[{{\mathbb{E}}\left[{\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}{\bm{Y}}_{j}}\right\|^{p}\Bigg|{\mathscr{G}}}\right]}\right]$
$=$ $2^{p}{\mathbb{E}}\left[{\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}{\bm{Y}}_{j}}\right\|^{p}}\right]$
$=$ $2^{p}\left\|{\sup_{k\in{\mathbb{N}}}\|\tilde{\bm{N}}_{k}\|}\right\|_{p}^{p},$
where we used the fact that $({\bm{Y}}_{k})_{k\in{\mathbb{N}}}$ satisfies (CI) condition and we denoted $\tilde{\bm{N}}_{k}=\sum_{j=0}^{k}\epsilon_{j}{\bm{Y}}_{j}$ which is also a conditionally symmetric martingale with $\langle\tilde{\bm{N}}\rangle_{k}=\langle{\bm{N}}\rangle_{k}=\langle{\bm{M}}\rangle_{k}$ and $\sup_{k\in{\mathbb{N}}}\left\|{\epsilon_{k}{\bm{Y}}_{k}}\right\|=\sup_{k\in{\mathbb{N}}}\left\|{{\bm{Y}}_{k}}\right\|$ which implies $\|\sup_{k\in{\mathbb{N}}}\|\epsilon_{k}{\bm{Y}}_{k}\|\|_{p}=\|\sup_{k\in{\mathbb{N}}}\|{\bm{Y}}_{k}\|\|_{p}\leq 2^{1/p}\|\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\|_{p}$ for any $k\in{\mathbb{N}}$.
Applying the Rosenthal-Burkholder inequality for conditionally symmetric martingale (Eqn. (13)) again, we have
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{N}}_{k}\|}\right\|_{p}\leq$ $2\left\|{\sup_{k\in{\mathbb{N}}}\|\tilde{\bm{N}}_{k}\|}\right\|_{p}$ (36)
[…]
Combining Eqn. (35) with Eqn. (36), we arrive at the conclusion.
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq$ $\left\|{\sup_{k\in{\mathbb{N}}}\|\tilde{\bm{M}}_{k}\|}\right\|_{p}+\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{N}}_{k}\|}\right\|_{p}$
[…]
When we take $c=p$, we determine the constants in the inequality: The coefficient of the quadratic variation term is
$5.01(1+e^{-1}+\frac{1}{p\vee\log d})(2+\sqrt{2})e\leq 87,$
when $p\vee\log d\geq 117$, the constant can be taken as $64$. The coefficient of the martingale difference term is
$5.01(1+e^{-1}+\frac{1}{p\vee\log d})(1+3\cdot 2^{1/p})\leq 50,$
when $p\geq 117$, the constant can be taken as $28$.
A.5 Proof of Lemma A.8
Proof.
We denote by $(X_{k}^{\prime})_{k\in{\mathbb{N}}}$ an independent copy of $(X_{k})_{k\in{\mathbb{N}}}$, which is also independent of $(\epsilon_{k})_{k\in{\mathbb{N}}}$. And we denote $\tilde{X}_{k}=X_{k}-X_{k}^{\prime}$. Then $\epsilon_{k}\tilde{X}_{k}$ and $\tilde{X}_{k}$ have the same distribution. Hence
${\mathbb{E}}\left[{F\left({\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}X_{j}}\right\|}\right)}\right]=$ ${\mathbb{E}}\left[{F\left({\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\left({X_{j}-{\mathbb{E}}[X_{j}^{\prime}]}\right)}\right\|}\right)}\right]$
$\leq$ ${\mathbb{E}}\left[{F\left({{\mathbb{E}}\left[{{\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\tilde{X}_{j}}\right\|\Bigg|(X_{k}^{\prime})_{k\in{\mathbb{N}}}}}\right]}\right)}\right]$
$\leq$ ${\mathbb{E}}\left[{F\left({\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\tilde{X}_{j}}\right\|}\right)}\right]$
$=$ ${\mathbb{E}}\left[{F\left({\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}\tilde{X}_{j}}\right\|}\right)}\right]$
$\leq$ ${\mathbb{E}}\left[{F\left({\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}X_{j}}\right\|+\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}X_{j}^{\prime}}\right\|}\right)}\right]$
$\leq$ $\frac{1}{2}\left\{{{\mathbb{E}}\left[{F\left({2\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}X_{j}}\right\|}\right)}\right]+{\mathbb{E}}\left[{F\left({2\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}X_{j}^{\prime}}\right\|}\right)}\right]}\right\}$
$=$ ${\mathbb{E}}\left[{F\left({2\sup_{k\in{\mathbb{N}}}\left\|{\sum_{j=0}^{k}\epsilon_{j}X_{j}}\right\|}\right)}\right],$
where we used the fact that the function $(x_{k})_{k\in{\mathbb{N}}}\mapsto\sup_{k\in{\mathbb{N}}}\left\|{x_{k}}\right\|$ from $\mathcal{X}$$^{\mathbb{N}}$ to $\mathbb{R}$$_{+}$ is a convex function and Jensen’s inequality. ∎
Appendix B Omitted Proof for the Matrix Rosenthal Inequality for Markov Chains
In this section, we complete the proofs omitted in Section 4.2.
B.1 Proof of Lemma 4.2
Proof.
To derive the desired conclusion, we need the following useful lemma, whose proof is in Appendix B.3.
Lemma B.1.
For any $\xi_{1},\xi_{2}\in{\mathcal{P}}$, vector-valued function ${\bm{h}}\colon{\mathcal{Z}}\to{\mathbb{C}}^{d}$, and matrix-valued function ${\bm{H}}\colon{\mathcal{Z}}\to{\mathbb{C}}^{d_{1}{\times}d_{2}}$, it holds that
  1. 1.
    Assume $\left\|{\left\|{{\bm{h}}}\right\|}\right\|_{\infty},\left\|{\left\|{{\bm{H}}}\right\|}\right\|_{\infty}<\infty$. Then it holds that
    $\left\|{\xi_{1}({\bm{h}})-\xi_{2}({\bm{h}})}\right\|\leq\left\|{\left\|{{\bm{h}}}\right\|}\right\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}},$
    $\left\|{\xi_{1}({\bm{H}})-\xi_{2}({\bm{H}})}\right\|\leq\left\|{\left\|{{\bm{H}}}\right\|}\right\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}}.$
  2. 2.
    Assume $\left\|{\left\|{{\bm{h}}}\right\|}\right\|_{V^{\alpha}},\left\|{\left\|{{\bm{H}}}\right\|}\right\|_{V^{\alpha}}<\infty$ for some $\alpha\in(0,1]$. Then it holds that
    $\left\|{\xi_{1}({\bm{h}})-\xi_{2}({\bm{h}})}\right\|\leq\left\|{\left\|{{\bm{h}}}\right\|}\right\|_{V^{\alpha}}\left\|{\xi_{1}-\xi_{2}}\right\|_{V^{\alpha}},$
    $\left\|{\xi_{1}({\bm{H}})-\xi_{2}({\bm{H}})}\right\|\leq\left\|{\left\|{{\bm{H}}}\right\|}\right\|_{V^{\alpha}}\left\|{\xi_{1}-\xi_{2}}\right\|_{V^{\alpha}}.$
Equipped with Lemma B.1, we are ready to prove this lemma.
  1. (a)
    By Douc et al. [2018, Proposition 21.2.3], we only need to show that $\|{\bm{G}}(z)\|<\infty$ for any $z\in{\mathcal{Z}}$.
    $\left\|{\left\|{{\bm{G}}}\right\|}\right\|_{\infty}=$ $\sup_{z\in{\mathcal{Z}}}\left\|{\sum_{k=0}^{\infty}\left({{\mathrm{Q}}^{k}{\bm{F}}}\right)(z)}\right\|$
    $\leq$ $\sup_{z\in{\mathcal{Z}}}\sum_{k=0}^{\infty}\left\|{\left({{\mathrm{Q}}^{k}{\bm{F}}}\right)(z)}\right\|$
    $=$ $\sup_{z\in{\mathcal{Z}}}\sum_{k=0}^{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)({\bm{F}})-\mu({\bm{F}})}\right\|$
    $\leq$ $\sup_{z\in{\mathcal{Z}}}\sum_{k=0}^{\infty}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\mu}\right\|_{\textup{TV}}$
    $\leq$ $\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}\sum_{k=0}^{\infty}\sup_{z\in{\mathcal{Z}}}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\mu}\right\|_{\textup{TV}}$
    $\leq$ $2\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}\sum_{k=0}^{\infty}\left({\frac{1}{4}}\right)^{\lfloor k/t_{\operatorname{mix}}\rfloor}$
    $\leq$ $2t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}\sum_{k=0}^{\infty}\left({\frac{1}{4}}\right)^{k}$
    $=$ $\frac{8}{3}t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
    where we used Lemma B.1 and Assumption 2.1.
    If we further assume ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$, for some ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$, then we have the following result: for any ${\bm{u}}\in{\mathbb{C}}^{d}$ and $z\in{\mathcal{Z}}$, it holds that
    $\left\|{{\bm{G}}(z){\bm{u}}}\right\|=$ $\left\|{\sum_{k=0}^{\infty}\left({{\mathrm{Q}}^{k}{\bm{F}}}\right)(z){\bm{u}}}\right\|$
    $\leq$ $\sum_{k=0}^{\infty}\left\|{\left({{\mathrm{Q}}^{k}{\bm{F}}}\right)(z){\bm{u}}}\right\|$
    $=$ $\sum_{k=0}^{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)({\bm{F}}{\bm{u}})-\mu({\bm{F}}{\bm{u}})}\right\|$
    $\leq$ $\sum_{k=0}^{\infty}\left\|{\left\|{{\bm{F}}{\bm{u}}}\right\|}\right\|_{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\mu}\right\|_{\textup{TV}}$
    $\leq$ $\frac{8}{3}t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}{\bm{u}}}\right\|}\right\|_{\infty},$
    where we used Lemma B.1. Hence
    ${\bm{u}}^{*}{\bm{G}}^{2}(z){\bm{u}}=$ $\left\|{{\bm{G}}(z){\bm{u}}}\right\|^{2}$
    $\leq$ $\frac{64}{9}t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}{\bm{u}}}\right\|}\right\|_{\infty}^{2}$
    $=$ $\frac{64}{9}t_{\operatorname{mix}}\sup_{z^{\prime}\in{\mathcal{Z}}}{\bm{u}}^{*}{\bm{F}}^{2}(z^{\prime}){\bm{u}}$
    $\leq$ $\frac{64}{9}t_{\operatorname{mix}}{\bm{u}}^{*}{\bm{\Upsilon}}{\bm{u}},$
    which is the desired conclusion.
  2. (b)
    We first state a lemma, which is a immediate corollary of Jensen’s inequality.
    Lemma B.2.
    [Durmus et al., 2023, Lemma 12] Assume Assumption $2.$ Then for any $\alpha\in(0,1]$, $k\in{\mathbb{N}}$, and $\xi\in{\mathcal{P}}_{V}$, it holds that
    $\left\|{\xi{\mathrm{Q}}^{k}-\mu}\right\|_{V^{\alpha}}\leq 2^{1-\alpha}\varkappa^{\alpha}\left({\mu(V)+\xi(V)}\right)^{\alpha}(1/4)^{{\lfloor k/t_{\operatorname{mix}}\rfloor}\alpha}.$
    The proof is almost identical:
    $\left\|{{\bm{G}}(z)}\right\|=$ $\left\|{\sum_{k=0}^{\infty}\left({{\mathrm{Q}}^{k}{\bm{F}}}\right)(z)}\right\|$
    $\leq$ $\sum_{k=0}^{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)({\bm{F}})-\mu({\bm{F}})}\right\|$
    $\leq$ $\sum_{k=0}^{\infty}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{\alpha}}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\mu}\right\|_{V^{\alpha}}$
    $\leq$ $2^{1-\alpha}\varkappa^{\alpha}\left({\mu(V)+V(z)}\right)^{\alpha}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{\alpha}}\sum_{k=0}^{\infty}(1/4)^{{\lfloor k/t_{\operatorname{mix}}\rfloor}\alpha}$
    $\leq$ $\left[{\varkappa^{\alpha}2^{3-\alpha}/(3\alpha)}\right](\mu(V)+V(z))^{\alpha}t_{\operatorname{mix}}\left\|{\|{\bm{F}}\|}\right\|_{V^{\alpha}},$
    where we used Lemma B.1, Lemma B.2, and the estimate in Durmus et al. [2023, Eqn (146)] in the last inequality.
B.2 Proof of Lemma 4.3
Proof.
  1. (a)
    By Minkowski inequality, we have
    $\left\|{\left\|{{\bm{S}}_{n}}\right\|}\right\|_{\mu,p}\leq\left\|{\left\|{\sum_{k=0}^{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor t_{\operatorname{mix}}-1}{{\bm{F}}}(Z_{k})}\right\|}\right\|_{\mu,p}+t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
    We define the function ${\bm{G}}_{t_{\operatorname{mix}}}\colon{\mathcal{Z}}\to{\mathbb{H}}^{d}$ as ${\bm{G}}_{t_{\operatorname{mix}}}=\sum_{k=0}^{\infty}{\mathrm{Q}}^{kt_{\operatorname{mix}}}{\bm{F}}$, which satisfies the Poisson equation ${\bm{G}}_{t_{\operatorname{mix}}}-{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}={\bm{F}}$ . By Lemma 4.2 (a), ${\bm{G}}_{t_{\operatorname{mix}}}$ satisfies the Poisson equation ${\bm{G}}_{t_{\operatorname{mix}}}-{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}={\bm{F}}$ and $\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}\leq({8}/{3})\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}$. By Minkowski inequality, we have
    $\left\|{\left\|{\sum_{k=0}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor t_{\operatorname{mix}}-1}{{\bm{F}}}(Z_{k})}\right\|}\right\|_{\mu,p}$ $=\left\|{\left\|{\sum_{r=0}^{t_{\operatorname{mix}}-1}\sum_{k=0}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor-1}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{kt_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\mu,p}$
    $\leq\sum_{r=0}^{t_{\operatorname{mix}}-1}\left\|{\left\|{\sum_{k=0}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor-1}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{kt_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\mu,p}$
    $\leq\sum_{r=0}^{t_{\operatorname{mix}}-1}\left\|{\left\|{\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\mu,p}+2t_{\operatorname{mix}}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}.$
    The summand of the first term is the spectral norm of a matrix martingale ${\bm{M}}^{(t_{\operatorname{mix}})}_{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}$, by Rosenthal-Burkholder inequality for matrix martingales (Theorem 2.1), we have
    $\left\|{\left\|{\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\mu,p}$
    $\leq$ $C_{1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}{\mathbb{E}}\left[{\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})}\right)^{2}\big|{{\mathscr{F}}_{r,k-1}}}\right]}\right\|^{1/2}}\right\|_{\mu,p}$
    $+C_{2}(p\vee\log d)\left\|{\sup_{k\leq\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\|{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})\|}\right\|_{\mu,p}$
    $\leq$ $C_{1}\sqrt{p\vee\log d}\Big\|\left({\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left({2\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}}\right)^{2}}\right)^{1/2}\Big\|_{\mu,p}+C_{2}(p\vee\log d)\left\|{2\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}}\right\|_{\mu,p}$
    $=$ $2C_{1}\sqrt{(p\vee\log d)\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}+2C_{2}(p\vee\log d)\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}.$
    where ${\mathscr{F}}_{r,k-1}=\sigma(Z_{jt_{\operatorname{mix}}+r},0\leq j\leq k-1)$. Combining the upper bounds above, we have
    $\left\|{\left\|{{\bm{S}}_{n}}\right\|}\right\|_{\mu,p}$ $\leq t_{\operatorname{mix}}(2C_{1}\sqrt{(p{\vee}\log d)\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}{+}2C_{2}(p{\vee}\log d)\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}){+}2t_{\operatorname{mix}}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}{+}t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}$
    $\leq\frac{16}{3}C_{1}\sqrt{(p\vee\log d)nt_{\operatorname{mix}}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}+\left({\frac{16}{3}C_{2}(p\vee\log d)+\frac{19}{3}}\right)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
    Furthermore, if there exists an ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$, then by Lemma 4.2 (a), it holds that ${\bm{G}}^{2}_{t_{\operatorname{mix}}}(z)\preccurlyeq(64/9){\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$. By Theorem 2.1 (b), the above inequality still holds if we replace $d$ with $e\cdot r([(64/9){\bm{\Upsilon}}]\wedge[(1/\lfloor{n}/{t_{\operatorname{mix}}}\rfloor)\|\|\langle{{\bm{M}}^{(t_{\operatorname{mix}})}}\rangle_{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\|\|_{\mu,\frac{p}{2}}])$, and double the right-hand side.
  2. (b)
    The proof is identical to that of part (a), except that we need to use Lemma 4.2 (b) instead.
B.3 Proof of Lemma B.1
Proof.
  1. 1.
    We first show the inequality for the vector-valued function. Recall the dual representation of the total variation [Douc et al., 2018, Proposition D.2.4]:
    $\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}}=\sup_{f\colon\|f\|_{\infty}\leq 1}(\xi_{1}(f)-\xi_{2}(f)),$
    then by the dual representation of the Euclidean norm, we have
    $\left\|{\xi_{1}({\bm{h}})-\xi_{2}({\bm{h}})}\right\|=$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}{\bm{u}}^{*}\left({\xi_{1}({\bm{h}})-\xi_{2}({\bm{h}})}\right)$
    $=$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}\left({\xi_{1}({\bm{u}}^{*}{\bm{h}})-\xi_{2}({\bm{u}}^{*}{\bm{h}})}\right)$
    $\leq$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}\|{\bm{u}}^{*}{\bm{h}}\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}}$
    $\leq$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}\|{\bm{u}}\|\left\|{\|{\bm{h}}\|}\right\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}}$
    $=$ $\left\|{\|{\bm{h}}\|}\right\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}},$
    which is desired.
    Now, we deal with the matrix case:
    $\left\|{\xi_{1}({\bm{H}})-\xi_{2}({\bm{H}})}\right\|=$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}\left\|{\left({\xi_{1}({\bm{H}})-\xi_{2}({\bm{H}})}\right){\bm{u}}}\right\|$
    $=$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}\left\|{\xi_{1}({\bm{H}}{\bm{u}})-\xi_{2}({\bm{H}}{\bm{u}})}\right\|$
    $\leq$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}\left\|{\|{\bm{H}}{\bm{u}}\|}\right\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}}$
    $\leq$ $\sup_{{\bm{u}}\colon\left\|{{\bm{u}}}\right\|\leq 1}\left\|{{\bm{u}}}\right\|\left\|{\|{\bm{H}}\|}\right\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}}$
    $=$ $\left\|{\|{\bm{H}}\|}\right\|_{\infty}\left\|{\xi_{1}-\xi_{2}}\right\|_{\textup{TV}},$
    where we used the inequality for the vector-valued function.
  2. 2.
    The proof is identical to (a) if we note that
    $\xi_{1}(f)-\xi_{2}(f)=$ $\int_{\mathcal{Z}}f(z)(\xi_{1}-\xi_{2})(dz)$
    $=$ $\int_{\mathcal{Z}}\frac{f(z)}{V^{\alpha}(z)}V^{\alpha}(z)(\xi_{1}-\xi_{2})(dz)$
    $\leq$ $\int_{\mathcal{Z}}\frac{\left|{f(z)}\right|}{V^{\alpha}(z)}V^{\alpha}(z)\left|{\xi_{1}-\xi_{2}}\right|(dz)$
    $\leq$ $\sup_{z\in{\mathcal{Z}}}\frac{\left|{f(z)}\right|}{V^{\alpha}(z)}\int_{\mathcal{Z}}V^{\alpha}(z)\left|{\xi_{1}-\xi_{2}}\right|(dz)$
    $=$ $\left\|{f}\right\|_{V^{\alpha}}\left\|{\xi_{1}-\xi_{2}}\right\|_{V^{\alpha}}.$
    Then we only need to replace $\|\cdot\|_{\infty}$ by $\left\|{\cdot}\right\|_{V^{\alpha}}$, and $\|\cdot\|_{\textup{TV}}$ by $\left\|{\cdot}\right\|_{V^{\alpha}}$ in (a).